A Bibliography of Publications by, and about, Olga Taussky [Olga Taussky-Todd]

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

16 August 2024 Version 1.11

Abstract

This bibliography records publications of Olga Taussky [Olga Taussky-Todd].

Title word cross-reference

#128 [Tau33e]. #19 [Tau78e].

(1, 2, 4, 8) [Tau71a]. 1 [TZ70]. $1.50 [Cas54]. 2 × 2 [Tau74a, Tau75a, Tau75b, Tau76a, Tau76b, Tau77c, Tau77b, Tau81a, Tau82d, Tau86b, Tau89a]. 4 [Tau77b]. A [KT56, Tau58c, Tau70b, Tau71d]. A∗ [KT56, Tau70b]. A − A′ [Tau79d]. A^{−1}A′ [Tau79d]. A^{∗} [Tau58c]. AX + X′A′ [TW62b].

C [Tau64b]. C^{n} → 0 [Tau64b]. G [Tau31d]. GL(n, F) [Tau68f]. I [TZ69]. L [MT52a, MT52c, MT53b, MT53a, MT55, RS71, Tau59, Tau62a, Tau72c, Tau72e, Tau74f, TZ69]. n [TW48a, TW48b]. n = 2 [Tau73d]. P [Tau72e, Tau74f, Tau60b, Tau60d]. p × p [Tau85b]. Q(\sqrt{−m}) [Tau77b]. Q(\sqrt{m}) [Tau77b]. ad [Tau71d]. SL(n, F) [Tau68f].

-cohomology [TZ69, TZ70]. -Dimensional [TW48a, TW48b]. -property [Tau72e, Tau74f]. -Quadrupel [Tau31d]. -Quadruples [Tau31d]. -rank [Tau77b]. -sums [Tau71a]. -th [Tau60b, Tau60d].

1948 [ABB^{+}48]. 1954 [Mey56, Tau55d]. 1969 [AB71]. 1973 [Tau74c]. 1975 [TWHM81]. 1976 [TWHM81]. 1978/79 [TT78]. 1981 [BW83]. 1983 [WS87]. 1st [van76].

20 [Tau74c]. 208 [ST35]. 223 [Sha97]. 23 [PT53]. 25 [Tau78a]. 2nd [van76].

3rd [BW83].

4723-4727 [LET^{+}57]. 4845-4850 [KTB^{+}59]. 4893-4898 [LTK^{+}60]. 4911-4916 [UEM^{+}60].

6 [TT78]. 60th [ABB^{+}48].

7.10.1995 [Ano15].

9 [GdG54].

A. [ST35, Tau58d]. AAAS [Whi86]. Abbildungen [Tau31c, Tau33f, Tau34a]. abelian [NT56]. Abgrenzung [Ger31]. Abhandlungen [Tau32a, Tau33a]. absolute [HNST54, HNST56, HNST03]. Abstract [Tau55a, Tau74b, Tau35a]. Abstrakte [Tau35a]. Abt [Tau33e]. Acknowledgements [Ano98a]. Activity [Tau55b, Tau55d]. addition [Tau36c, Tau36d]. Additive [Tau74a, Tau75a, TW62a]. adjugate [Tau68c]. Administrators [Ano17]. Advanced [Abh71, BF60, BKM^{+}67, BvL68, BRM70, BEM76, CGJ60, CTT74, DKB^{+}69, DW70, ET46, Erd54, EY61, ES61, GLF61, GG74, GC68, HH74, HC63, JS69, KO70, Kee61, KP69, KH69, KRO80, KF60, KS68, KTB^{+}59, Koe60, KMCS63, KMS64, LET^{+}57, Lam58, LTK^{+}60, LC61, Mac58, Mau73, MB74, MTD68, MD69, MHS82, MT54, MC61, MTP^{+}78, New61, PPdB80, PJR^{+}68, RWA70, RH61, RM60, SST^{+}68, SR69, STS^{+}52, Sha52, Sha54, SN61, SD63, SRvR69, SK63, SyW54, SAG68, SG68, Tau52a, TE54, TS54, Tau57a, TSE58, TP60, Tau60a, TM61, Tau61a, TWS^{+}62, TM63, TC69, TS70, TG80, TL58, UEM^{+}60, UvL61, VM68, WB63, War64, WW74, WF68, WF69, YCJ63, Yff64]. aerofoil [Tau45, Tau48a, Tau48b, Tod45]. ähnliche [Tau31c]. Alan [Mic03, Tau56a]. Alexander [SDGT81]. Algebra [HHOT94, Ost83, Sto11, TWHM81, Tau33a, Tau58a, Tau64a, Tau67a, Zas77, vdW31, Laf98, Tau36e, Tau37c, TT47, Tau81b, Tau84, Ten08, TT78, TH31, TH32, Tau33b, Tau36e, HT32, Tau60f]. Algebraic [Coh78, Coh88, Ost83, Tau55a, DTZ61, DTZ62, DT64, FTT55a, NT56, Tau32d, Tau36b, Tau39, TT40a, Tau54d, Tau56b, Tau62f, Tau68b, Tau78c, Tau80b, Tau80e, Tau81e, Tau82b, Tau83b, Tau83c, Tau87b, Tau88c, Tau89b, ST56]. algebraische [Tau32d]. Algebras [CR62, Tau78b]. Algorithms [Tau70d]. Alternative [ST80, ST80]. American [Cur56, Car00, TWHM81, Tau88b]. AMS [Tau78a, Ano98c, Tau74c]. Amsterdam [GdG54]. analogs [FTT55b]. Analysis [PT53, Tau50b, Tod62, Var62, Var00, Var09, Bha01, Cur56, Mad12, Mei96, Tau55d, Tau55b]. Analytical [Tau36a]. André [Bre96, Bre06a, Bre06b, Bre11, Bre14, Ben17]. André-Louis [Bre14, Ben17]. Angeles [PT53]. Answer [Tau78a]. Appendix [TT44]. Application [PT57]. Applications [TTT57, TT57a, TT12a, Du00, Tau47b, Tra92, HHOT94]. Applied [Cur56, TWHM81]. approach [Tau74b]. April [BW83]. arising [Tau48a, Tau48b, Tau66d, Tau79a]. Arlington [TWHM81]. Army [Ben17, Bre14]. Arnold [LR16b, Tau52b, Zas77]. Array [BL77]. Art [Tau70d]. Articles [CV76, MT75, Tau77d]. Artin [Tau78d, Tau32b, Tau88a]. aspects [Laf98, TT47, Tau82a]. Association [ST35]. Associative [CR62]. Atlas [AB71]. Aufgabe [Tau33e, ST35, Sch33]. August [AB71, PT53]. Autobiographical [TT85a]. Autobiography [Tau80a, Ter80]. Automorphs [Tau68a, Tau72a]. Axiomatik [Tau33g, Tau34b]. axioms [Tau33g, Tau34b].

B [TT78, Tau33b, Zas77]. B. [TH31, TH32, van76]. Band [Tau32a, Tau33a, vdW31]. Bases [Tau47a]. Basic [Tod78]. basis [NT56]. Baumert [Tau78a]. Bd [TH31, TH32, Tau33b]. be [Tau80d]. Became [Tau88e]. Beitrag [Tau87d]. Berlin [Tau33b, Tau32a, Tau33a]. Bestimmung [ST34, TS34]. Between [TZ59, Tau46, TW62a, Tau66c, Tau74a, Tau77b, Tau78c, Tau85b, Tau88c]. Beweis [Bre56]. beyond [Han98]. Bibliography [OTT49, Tau51a]. bilinear [TT56a]. binary [PT57, Tau81a]. binomial [Tau25]. Biobibliographic [GC87]. Biography [Car00]. bird [Tod97]. Birthday [ABB^{+}48]. bis [Sta88]. bodies [Tau35a]. body [ST34, Tau32e, TS34, Tau35a]. Book [Ben17, Cas54, Djo68, Fle55, Tau46, Tau47a, Tau55b, Tau64a]. Books [Ano54a, Ano54b, Ano54c, Ano55, Ano78]. boundary [Tau45, Tau48a, Tau48b, Tod45]. bounding [Ger31]. Bounds [Tau48c, Tau48d, Tau51a, Tau51b, Tau47b, Tau62b]. Brezinski [Ben17]. Briefwechsel [LR16a]. Brouwer [Tau91b]. Bryn [SS83, QMLT83]. Bureau [Cas54].

CA [Tau96]. Calculations [BL77, Tau47b]. California [PT53]. can [Tau80d]. caractéristiques [Tau53b]. Careers [Cas02]. Carlo [Mey56]. carrés [Tau53b]. case [Tau73d]. Cayley [PT70]. Celebration [Cas02]. celestial [Tau88d]. Centennial [Tau88b]. central [Tau78b]. Century [Lid01]. changes [Tau64c]. characterisation [TT40a]. Characteristic [LT60, Tau48c, Tau49c, Tau51a, Tau54a, Gai75, Tau47b, Tau48d, Tau51b, TT51, Tau64c, Tau66c, Tau68a, Tau68e, Tau70a, Tau83a, Tau86a]. characterization [HT53, HT54, Tau62a]. Cholesky [Ben17, Bre96, Bre05, BGC05, Bre06a, Bre06b, Bre11, Bre14, TT06]. Chronicle [Lid01]. Circles [Var04, Tau85c]. circulant [Kai12]. circulants [NT57, Tau55e, Tau85b, Tra92]. Class [Coh78, Coh88, Tau32b, Tau37a, Tau37b, Tau78d, BT42, Kis97, Kis98, ST34, Tau32e, TS34, Tau35b, Tau77b, Tau88a]. class-body-tower [ST34]. class-objects [ST34]. Classes [NT57, Tau50a, Tau51c, Tau52c, DTZ61, DTZ62, Tau57d, Tau71c, Tau74d, Tau79a, Tau81a]. Classical [Coh78, Coh88, Tau64a]. Claude [Ben17]. cohomology [TZ69, TZ70]. Collected [KS88, Ost83, Zas77]. Collection [CV76, MT75]. Commentaries [Tau96]. commentary [Mic03]. comments [TT78]. Common [Tau78e]. Communications [TZ69]. commutative [Tau36c, Tau36d, Tau83b]. Commutativity [Gai66, Tau57b, Sha05, Tau69a, Tau78e]. commutator [ET81, Gai68, Gai75, MK60, Tau58c]. Commutators [KT56, Tau61b, Tau53a, Tau54c, TW62a, Tau74a, Tau75a, Tau80b]. commute [Tau61b]. Commuting [TT56a, Tau76a, Tau77c]. compact [JT35]. complex [Tau80d, Tau85c, Tau86a]. Complexities [CL05]. Composition [Tau81a, Tau80, TH80]. computational [Tau56b, Tau61f, Tau62f, Tau70e, Tau74b, Tau60f]. Computations [Dwy53, TT60]. Computer [Tau70d]. Computers [AB71, Tau61e]. Concerning [LT60, Tau49c, Tau74f, Tau82d, ET81, GT79, TT78, Tau62b, Tau68f, Tau69c, Tau70b, Tau71c, Tau72e, Tau73h, Tau74d, Tau76b, Tau80, Tau79d, TH80, Tau80e, Tau81e, Tau85b]. Condition [Tau50b]. Conference [BW83, Mei96, TWHM81]. Congress [GdG54]. congruence [Tau83a]. connected [DT64, DT65]. Connections [Tau88c, Tau78c, Tau85b]. contribution [Tau87d]. Contributions [Tau54b]. converse [Tho83, Yin94]. Converses [Alz91]. correlation [HNST54, HNST56, HNST03]. Correspondence [LR16a, LR16b, LR16c]. correspondences [EG84]. corresponding [Tau57d]. Council [AB71, Whi86]. Couples [Tau47a]. Courant [ABB^{+}48]. course [Sha97]. Covering [TT48a, TT48b, TT49]. Craig [Tau58d]. criterion [Ste84]. cubic [ST34, TS34, Tau87c]. Curriculum [Ano98b]. curve [Tau86a]. curves [Gai75, Tau87c]. cyclic [Tau78b].

D [Sch33, Tau78a]. D.C [Cas54]. Dade [Fal71]. dame [McL96]. d’André [BGC05]. David [Tau43, Tau32a, Tau33a]. DC [Tau96]. Dedekind [Tau33d]. Dedicated [CV76, MT75, Zas77]. Definite [OT83, NT57, OT51, Tau67d, Tau68e, Tau70a]. Department [Cas02]. Descriptions [Ano55]. Determinant [OT83, OT51, TW63]. determinantal [Tau57c, Tau66a, Tau87c, Tho83]. Determinants [Ost83, Pri78, TTT57, TT78, TT40b, Tau49b, TT52a, TT52b, TT57a, TT57b, Tau74e, TT12a, TT12b]. Determination [Ano54a, Cas54, Fle55, PT53, Tau54b, TS34]. Deutsche [ST35]. Development [Sto11, Tau46, Mad12, Ste84]. diagonal [Pri78, TT78, Tau73h]. did [Hof98]. different [DT64, Tau35b, Tau74a]. Differential [O’L01, Tau39, Tau48a, Tau48b, Tau58f, Tau67c]. Digital [Tau61e]. Dimensional [Gai66, TW48a, TW48b]. Diophantine [Tau79a]. direct [Tra92]. Discrete [FTT55b, TT60]. discriminant [Tau68b]. Discursive [Sch77]. Discuss [Tau60e]. Distinguished [Ano17]. Divisors [DRTW64, Tau84]. doctorate [Hla97b]. Documents [Cas54]. DOE [Cas02]. d’Olga [Ser70]. dominant [Pri78, TT78]. Dominique [Ben17]. Donald [Tau70d]. Dover [Tau60f]. Dr [Ano15].

E. [Tau32b, Tau78d]. E771

[CBR^{+}47, CF48]. E772 [BMK48]. E773 [RM48]. E774 [AH50]. E775 [BT48, CBR^{+}47]. Economic [Tau55b, Tau55d]. ed [Tau55d]. edited [Ano54a, Koh84, Tau61e]. Editor [Tau32b, Tau32a, Tau33a, Tau91b]. Eds [Cas54]. Eigenvalue [O’L01]. Eigenvalues [PT53, Tau54b, Tau62b, Ger31, Tau60e, Tau85c, Ano54a, Cas54, Fle55]. Eigenwerte [Ger31]. eight [Tau66a]. Einfluß [ST34, TS34]. Einführung [Tau33c, DB31]. Einzeldarstellungen [Tau33b]. Electron [BL77]. Elementare [Tau64a]. Elementary [AH50, BT48, BMK48, CBR^{+}47, CF48, RM48, Tau64a]. elements [TT51, Tau64c]. Elisabeth [LR16c]. Emigration [Sta88]. Emmy [BS81, Dic70a, Dic70b, Dic81, QMLT83, SS83, Tau81c, Ten08]. Endliche [Tau35a]. Energy [Cas02]. Engendrent [Tau47a]. Engineer [Bec13a, Bec13b]. Engineers [Ano17]. entries [Tau86a]. equation [Tau39, Tau48a, Tau48b]. Equations [Ano54a, Cas54, Fle55, Ost83, PT53, TTT57, Tau54b, ST56, Tau50c, TT52a, TT52b, TT57a, TT57b, Tau58f, Tau67c, TT12a, TT12b]. Erratum [JZ95]. Errors [Dwy53]. Erster [Tau32a]. Essay [TT85a]. Essays [ABB^{+}48]. Excellence [Lid01]. exercise [Tau33e, ST35]. Exil [Sta88]. extensions [EG84]. Eyewitness [Tau88h].

factor [Tau61b]. Factorization

[PT70, Tau68c, Tau72b, Tau73b, Tau73c, Tau73a, Tau73d, Tau86b, TT06]. factors [Tau54c]. Facts [Tau82d, Tau76b, Tau80e, Tau81e]. Faddeeva [Tau60f]. family [Bre11]. Fan [Alz91, Lun94, Yin94, ZX05]. Fermat [Tau85b]. Field [Tau37a, Tau60b, Tau78d, DTZ61, DTZ62, DT64, Kis97, Kis98, Tau32b, TS34, Tau60d, Tau68b, Tau80e, Tau81e, Tau88a, Tauxx]. Fields [Coh78, Coh88, Tau35b, Tau37b, NT56, ST34, Tau32d, Tau50a, Tau51c, Tau52c, Tau74a, Tau75b, Tau76a, Tau77a, Tau77c, Tau77b, Tau78b, Tau80b, Tra92]. Final [Cas02]. Finite [CR62, Gai66, LT60, Tau49c, Tau51a, Tau57b, Tau74f, MT52b, TT40c, TT41b, Tau62b, Tau64c, Tau68c, Tau70b, Tau72e, Tau35a]. First [Bec13a]. Florida [Mey56]. flutter [Tau47b]. For.Mem.R.S. [Tau43]. Foreword [Tau91a]. form [PT57, Tau79d, van76]. Forms [Koh84, Tau82d, Tau82c, Tau68a, Tau72a, Tau76b, Tau80d, Tau81a]. Foster [TF85]. four [PT57]. fourth [TWHM81]. Fox [Sch33]. French [Bre05, Ser70]. function [TW62b]. functions [Tau85c].

G. [TT78]. Gallery [Ano17]. Gatlinburg [Sto11]. Gauss [Tau74b]. Gauthier [Tau53b]. geb [Tau33b]. Gebiete [Tauxx]. Gedächtnis [Tau52b]. Geiringer [Bre56]. genera [DTZ62]. General [BW83, Tau72d, Tau68e, TZ69, TZ70, Tau70a, Tau73d, BW83]. Generalization [Fal71, Tau61c, Du00, Fri81, Lun94, NT56, RS71, Tau69a]. Generalizations [Gai66]. Generalized [Tau53a, Tau54c, Tau68a]. Generated [Tau60b, Tau60d]. Generation [TT54, TT56b]. genus [Tau77a, Tau77c]. Geometrie [Tau33a]. Geometry [TW48a, TW48b]. German [ST35, Bre56, DB31, FT36, Ger31, LR16a, LR16b, LR16c, LR16d, ST34, ST35, Tau25, Tau30b, Tau30a, Tau31c, Tau31e, Tau31d, Tau31a, Tau32e, Tau32d, Tau33g, Tau33f, Tau33e, Tau34b, TS34, Tau34a, Tau35a, Tau36e, Tau52b, Tau64a, Tau87d, Tau88h, Tauxx, Tod97]. Gersgorin [Var04]. Gesammelte [Tau32a, Tau33a, Tau32c, Tau33d]. gestellt [ST35, Sch33]. glue [Tod97]. Gödel [WS87, Tau87e, Tau88g]. Göttingen [Tau78d, Tau88a]. GPO [Cas54]. Grande [McL96]. grant [Cas02]. group [Tau55c, Tau58c, TZ69]. Groupe [Tau47a]. Groups [CR62, BT42, GT79, MT52b, Tau30b, Tau31c, Tau31e, Tau31b, Tau33g, Tau33f, Tau34b, Tau34a, TT41a, TT48a, TT48b, TT49, TZ70, Tau77b]. Grundlehren [Tau33b]. Gruppen [Tau30b, Tau31c, Tau31e, Tau33g, Tau33f, Tau34b, Tau34a].

H [Tau61e]. H. [Tau57c, Tau58d]. h.c [Ano15]. Hadamard [Tau71a]. Hahn [Tau97]. Handbook [CO58, CO67]. Hans [KS88, Tau88f, Tau97]. Hasse [LR16a, LR16a]. Hauptideale [ST34, TS34]. Hauptidealsatzes [Tau32d, Tau30a, Tau31a]. having [Tau72c]. Heilbronn [KS88, Tau88f]. held [AB71, Mey56, PT53, TWHM81]. Henry [Zas77]. Her [BS81, TT88, TT05]. Heritage [TWHM81]. Hermitian [Tau60e]. higher [TW62a]. Hilbert [Tau32a, Tau33a, Tau43, Tau49c, Tau54f, Tau69b, Tau69c, Tau71b, Tau73e, Tau91b]. historical [Pri78, TT78, Tau87c]. History [Tau81b]. Hoffman [Mic03]. Honoring [SDGT81]. honour [Mei96]. Hotelling [Tau58d]. hyperbolic [Tau48a, Tau48b]. hypercomplex [Tau36a].

Ideal [Tau30a, Tau31a, Tau32d, Tau62c, Tau63, Tau85a, Tau87a, DTZ61, DTZ62, Tau57d, Tau77b]. ideals [ST34, TS34, Tau80e, Tau81e, Tra92]. identical [Tau74a, Tau83a]. identities [ET81]. identity [Tau66a]. ihr [ST34, TS34]. ihre [ST34, TS34]. II [Tau33b, MT53b, MT55, Ost83, TH32, Tau50b, Tau51b, Tau52c, TT52b, Tau63, Tau69b, Tau72b, Tau73d, Tau74f, Tau77c, vdW31]. III [Ost83, Tau77b, Tau85a]. Illus [Cas54]. im [ST35]. imaginär [TS34]. imaginär-quadratischer [TS34]. imaginärquadratischer [ST34]. Imaginary [LT60, ST34, TS34]. imaginary-square [ST34]. indecomposable [Ste84]. index [HHOT94]. Inequalities [BW83, Alz91, FTT55b, Lun94, ZX05]. inequality [FTT55a, Joh90, JZ95, Tau57c, Tho83, Yin94]. Infinite [TT42, Tau84]. Infinity [Tau82e]. Influence [Sch77, Sto11, ST34, TS34]. Institute [PT53]. Instruction [ET46]. integers [DT65, Tau60c, Tau80c]. Integral [BL77, O’L01, Tau58b, Tau66b, Tau82d, Tau87b, TT51, Tau55e, Tau61f, Tau66c, Tau72b, Tau73b, Tau73c, Tau73a, Tau73d, Tau74a, Tau76b, Tau78c, Tau79a, Tau80e, Tau81a, Tau81e, Tau85b, Tau86b, Tau88c, Tau89a]. International [BW83, GdG54]. Internationale [Sta88]. interview [Tro73]. Interviews [AA85]. Introduction [Sha98a, Tau69d, Tau78c, DB31]. invariant [Tau78e]. Invariantentheorie [Tau33a]. inverse [Tau57d]. Inversion [TT41a]. Invitation [Coh78, Coh88]. involving [Tau61f]. irreducible [Tau66c]. isomorphe [Tau33f, Tau34a]. isomorphic [Tau33f, Tau34a]. isoperimetric [FTT55a]. issue [ABR97b, Sha98b]. italics [ST35]. Iteration [O’L01]. Iterative [Var62, Var00, Var09]. IV [Tau87a].

J [Ano54a, Cas54, Fle55, Sch33, Tau32a, Tau33a, Tau33b]. J.T. [Tro73]. Jahresbericht [Sch33]. January [ABB^{+}48]. Jber [ST35]. John [Alb07, Alz91, Bau07, Mad12, Tau87c]. Jordan [Tau46, Tau60e, Tau65b]. July [WS87].

Kac [TT91]. Kato [Gai68, Gai75]. Klassenkörper [ST34, TS34]. Klassenkörpers [Tau32e].

Klassenkörperturm [ST34, TS34]. klassische [Tau64a]. kleinere [Tauxx]. Knuth [Tau70d]. Körper [Tau35a]. Körperpotenzen [Tau35a]. Korrespondenz [LR16b, LR16c, NT45]. Korrodi [LR16c]. Krull [Tau64a, Tau64a]. kubischen [ST34, TS34]. kursiv [ST35, Sch33]. Kurt [Tau87e, Tau88g]. Kvadrata [Tau69d]. Ky [Alz91].

L [Ano54a, Cas54, Fle55, TH31, TH32, Tau33b, Tau78a, van76]. L. [TF85]. Laplace [Tau39]. last [Tau85b]. Latimer [EG84, Han98, Tau48e, Tau49a]. Leading [Tau82d, Tau76b]. Leben [LR16d]. Lectures [Tau78d, Tau32b, Tau88a]. left [Tau84]. Leim [Tod97]. Letter [Tau91b]. letters [van76]. Life [BS81, Car00, LR16d, Bre06a, Bre06b, Bre11]. Linear [Ano54a, Cas54, Fle55, Frö53, Gai66, O’L01, Ost83, PT53, Sto11, Tau54b, Tau60f, TW62a, Tau50c, TZ69, TZ70, TT78, HHOT94]. link [Tau77b]. Literaturberichte [Tau32c, HT32, Tau33c, Tau33b, Tau33d]. Lives [Car00]. Locally [JT35]. location [Tau60e]. loci [Tau87c]. Logic [Tau46]. Logical [Tau46]. look [TT91]. Lösung [Tau33e]. Lösungen [ST35, Sch33]. Louis [Bre96, Bre11, Bre14, Ben17]. Lyapunov [Tau61c, Tau61d].

M [SDGT81, Sch33, Tau53b]. MacDuffee [EG84, Tau48e, Tau49a, Han98]. Machine [Frö53]. main [ST34, TS34]. Mann [Zas77]. many [Tau82a]. mappings [Tau33f, Tau34a]. March [Mey56]. Marcus [ST80, ST80]. Mark [TT91]. Math [ST35, Sha97]. Math-Verein [ST35]. Mathematical [AA85, Cur56, Ost83, TWHM81, Tau46, TW48a, Tau60b, Tau62d, ST34, TS34, Tau61e]. Mathematician [Ben17, Bre14, Oli95, Zun00]. mathematicians [GdG54, Uhl07].

Mathematics [Alb07, Bec13a, Bec13b, Cas02, CL05, Cur56, ET46, GC87, JB57, JB12, TWHM81, Tod78, McL96, Tau88b, Tauxx, Uhl07]. Mathematik [Tauxx]. Mathématiques [Tau53b]. mathematische [Tau32c, Tau33d]. mathematischen [Tau33b]. Mathison [Tau56a]. Matrices [MT52a, MT53b, TTT57, TT40c, TT41b, Tau48c, Tau50b, Tau51a, Tau57b, Tau58b, Tau60c, Tau62d, Tau62e, Tau64b, Tau67b, Tau68d, Tau72d, Tau74f, Tau82d, DT65, HT53, HT54, MT52c, MT53a, MT55, Sha05, TT42, Tau44, Tau47b, Tau48d, Tau50a, Tau51b, Tau51c, TT51, Tau52c, TT52a, TT52b, Tau53a, Tau53b, Tau54a, Tau54c, Tau54d, Tau55c, TT56a, TT57a, TT57b, Tau59, Tau60e, Tau61b, Tau61f, Tau62b, Tau62c, Tau63, TW63, Tau66b, Tau67d, Tau68b, Tau68e, Tau70a, Tau70c, Tau71c, Tau71a, Tau72b, Tau72c, Tau72e, Tau73b, Tau73c, Tau73a, Tau73d, Tau74d, Tau75a, Tau75b, Tau76a, Tau76b, Tau77c, Tau77b, Tau78c, Tau79a, Tau79d, Tau80b, Tau80d, Tau81a, Tau83a, Tau84, Tau85a, Tau85d, Tau86a, Tau87b, Tau87a, Tau88c, Tau89a, TT06, TT12a, TT12b, TT44]. matrices [Tod44, Djo68]. Matrix [Dwy53, LT60, OT83, Sch77, Tau49c, TZ59, Tau65a, Tau88e, Var62, Var00, Var09, Ger31, GT79, JF98, OT51, PT65, ST80, ST80, Sha97, Sha05, Ste84, Tau54f, Tau57d, Tau58d, TW62b, Tau64c, Tau65b, Tau66d, Tau66c, Tau68c, Tau69a, Tau70b, Tau71d, Tau72b, Tau73b, Tau73c, Tau73a, Tau73d, Tau73e, Tau73h, Tau74a, Tau74b, Tau77a, Tau80b, Tau81a, Tau82b, Tau86b, Tau87d, TT91, Ger31, Tau55a]. Matrizen [Tau87d]. matters [Tau84]. Mawr [SS83, QMLT83]. May [BW83]. me [Hof98]. Measurements [Lid01]. mechanics [Tau88d]. Meeting [Whi86]. Meinguet [Mei96]. Mémorial [Tau53b]. Memoriam [ABR97a, ABR98, Bau07, LM96, Tau52b]. memory [ABR97b, Sha98b]. Mengen [Tau35a]. Method [O’L01, Bre05, Tau47b, Tau86b]. méthode [Bre05]. Methods [Mey56, Tau55a, Tau60f, Tau61e, Tau36a, Tau83b, Tau83c, Tau89b]. metric [Tau30b, Tau31e, Tau35a]. Metrik [Tau30b, Tau31e, Tau35a]. Mladih [Tau69d]. modal [Bha01]. Modern [Bec13a, Bec13b, Tau36b, Mad12, TT47, Tau64a, Tau74e, Ten08]. Moderne [HT32, TH31, TH32, Tau33b, vdW31]. modernen [Tau64a]. Molecular [BL77]. Mon [ST35]. monotonic [Tau54e]. Monte [Mey56]. Monthly [Ano54b]. Morgenstern [Tau55d]. mother [Ten08]. motion [Tau45, Tau48a, Tau48b, Tod45]. Motzkin [Fri81, RS71]. Moyls [ST80, ST80]. Multilinear [TT78]. Multiply [Tau54e]. My [Tau79b, Tau81c, Tau88f].

N [Tau60f]. Nation [Car00]. National [Car00, Cas54]. natural [ST56]. Naucni [Tau69d]. NBS [Lid01]. NBS/NIST [Lid01]. Neuer [Bre56]. NIST [Ano17, Lid01]. no [AB71, TT78, Tau53b]. Noether [SS83, BS81, Dic70a, Dic70b, Dic81, QMLT83, Tau81c, Ten08]. non [Tau36c, Tau36d, Tau45, Tau48a, Tau48b, Tau76a, Tau77c, Tau83b, Tau84, Tod45]. non-commutative [Tau36c, Tau36d, Tau83b]. non-left [Tau84]. non-uniform [Tau45, Tau48a, Tau48b, Tod45]. Noncommutative [Fal71, Tau83c]. noncommutativity [Tau89b]. noncommuting [Tau75b, Tau77b]. Nonsingular [Tau87c]. nonsingularity [Ste84]. Normal [Tau54d, HT53, HT54, NT56, Tau85d]. Norms [Tau75b, Tau76a, Tau77c, Tau77b, Tau82c, Tra92, Tau80b, Koh84]. Note [TW48a, TW48b, Pri78, TT78, Tau44, Tau55c, Tau58c, TT44, Tod44, Tau50b]. Noted [Oli95]. Notes [Sha97, TW48a, Tau50b, Tau60b, Tau62d]. Notices [Tau78a, Tau74c]. November [TWHM81]. Number [DB31, Tau55a, Tau62f, Tau70e, Tau79b, Zas77, AB71, DTZ61, DTZ62, DT64, HNST54, HNST56, HNST03, NT56, ST34, Tau32d, TS34, Tau36b, Tau54d, Tau56b, Tau68b, Tau74a, Tau74b, Tau78c, Tau80b, Tau80e, Tau81e, Tau82b, Tau83b, Tau83c, Tau87b, Tau88c, Tau89b, Tra92, Tau79c]. number-fields [ST34]. Numbers [Coh78, Coh88, Sha05, TT54, PT70, ST56, TT40a, TT56b]. Numerical [Alb07, PT53, Sto11, Tau50b, Tod62, Tod78, Mad12, Mei96, ST80, ST80, Cur56, Mei96].

O [Tau55d]. O.T. [Tro73]. Oberwolfach [BW83, Tod83]. Obituary [Tau43, Wei96]. objects [ST34]. obtaining [Tau47b]. October [TWHM81, Tau74c]. off [Tau73h]. Offered [Ano98c]. officer [Bre14, Ben17]. Oktober [Sta88]. Olga [Ano54a, Cas54, Djo68, Fle55, Koh84, LR16b, LR16c, Sch33, Alz91, Ano98b, Ano98d, Ano98e, Ano98f, Ano98g, Ano15, ABR97a, ABR97b, ABR98, Bau98, CV76, Cas02, Dav97, Dav05, Goo20, Hla97a, Hla97b, Hof98, JF98, Kai12, Kis97, Kis98, Laf98, LR16d, LV66, Luc87, LM96, MT75, McL96, New80, Oli95, Rom91, Sch77, Sch98a, Sch98b, ST35, Ser70, Sha97, Sha98b, Sha05, Ste84, TT77, TF85, Ter80, Tra92, Uhl07, Wei96, Zas61, Zas77, Zun00, van76]. one [Tau61b]. Online [Car00]. operator [Tau71d]. Operators [O’L01, Tau58e, Tau70c, Tau72a, Tau78e]. Order [Cas54, DTZ61, DTZ62]. Orders [Fal71, DTZ62, DT64, Tau74a]. Ordinary [Tau58f, Tau67c]. österreichischer [Sta88]. Ostrowski [BH54, SDGT81, Tho83]. ovals [Tau85c]. Own [TT88, TT05]. Oxford [AB71].

P [Tau57c]. P106 [Tau72f]. P107 [Tau72f]. P91 [Tau72f]. Paige [Ano54a, Cas54, Fle55].

Pairs [MT52c, MT53b, MT53a, MT55, Tau72c, Tau74f, Tau80c, Tau59, Tau72e, Tau89a]. paper [Red65, ST56, Zas61]. Papers [KS88, Ost83, Zas77, Mic03, TWHM81, Tau96]. Paris [Tau53b]. Parodi [Tau53b]. part [Tau73h]. particular [DTZ62]. Parts [LT60]. Pasadena [Tau96]. Paso [TWHM81]. pencil [Tau83a, Tau86a]. People [AA85, Car00]. period [TT40c, TT41b]. periods [Tau85c]. permutations [Tau54c]. Personal [Sch77, Tau88f, Pri78, Sch98a, Sch98b, TT78, Tau81c]. Physics [CO58, CO67]. Piccard [Tau47a]. pictures [Tau31c]. Pioneer [Alb07]. Poem [Tau79c]. poems [Tau82e]. points [HNST54, HNST56, HNST03]. Poland [Tau46]. polygons [FTT55a]. polynomial [Tau25, Tau66c, Tau83a, Tau86a]. polynomials [Tau78b]. Positive [OT83, Tau67d, Tau68e, Tau70a, NT57, OT51, Tau72a]. Positive-definite [Tau67d, Tau68e, Tau70a]. Positivism [Tau46]. possible [Tau85b]. potencies [Tau35a]. powers [TT42, Tau84]. pp [Cas54]. Practical [ET46]. preceding [ST56]. Preface [ABR97b]. Preis [Tau33b]. Preliminary [Fal71]. Presented [ABB^{+}48]. Price [TT78]. principal [Pri78, TT78, Tau77a, Tau77c, Tau30a, Tau31a, Tau32d]. Problem [O’L01, Rom91, Tau78e, JF98, Kai12, LV66, New80, Tau48a, Tau48b, Tau54f, Tau54e, Tau60e, Tau65b, Tau70c, Tau73f, Tau74b, Tau79a, TG81, Tau85c, TF85]. Problems [Abh71, AH50, BF60, BKM^{+}67, BvL68, BT48, BRM70, BMK48, BEM76, CBR^{+}47, CF48, CGJ60, CTT74, DKB^{+}69, DW70, Erd54, EY61, ES61, GLF61, GG74, GC68, HH74, HC63, JS69, KO70, Kee61, KP69, KH69, KRO80, KF60, KS68, Koe60, KMCS63, KMS64, Lam58, LC61, Mac58, Mau73, MB74, MTD68, MD69, MHS82, MT54, MC61, MTP^{+}78, New61, PPdB80, PJR^{+}68, RWA70, RH61, RM60, RM48, SST^{+}68, SR69, STS^{+}52, Sha52, Sha54, SN61, SD63, SRvR69, SK63, SyW54, SAG68, SG68, Tau52a, TE54, TS54, Tau57a, TSE58, TP60, Tau60a, TM61, Tau61a, TWS^{+}62, TM63, Tau68f, TC69, TS70, TG80, TL58, UvL61, VM68, WB63, War64, WW74, WF68, WF69, YCJ63, Yff64, Ano98d, TT85b, Tau36b, Tau37c, Tau45, Tau54d, Tau56b, Tau58g, Tau61f]. problems [Tau62f, Tau65b, Tau70c, Tau70e, Tau72f, Tau73g, Tau73i, Tau81d, Tau85d, Tod45, Tra92, KTB^{+}59, LET^{+}57, LTK^{+}60, UEM^{+}60]. Proceedings [Cur56, SS83, AB71, PT53, GdG54]. product [Tau54c, Tau60e, Tau73b, Tau73c, Tau73a, Tau73d, Tau80c]. Products [Tau65b, TT51]. Prof [Ano15]. Prof. [Tau43]. Profiles [AA85]. Programming [Tau70d]. Progress [Kai12]. proof [Bre56, FTT55a, ZX05]. proofs [Lun94, ST80, ST80]. Property [MT53b, MT52a, MT52c, MT53a, MT55, RS71, Tau39, Tau59, Tau62a, Tau72c, Tau72e, Tau74f, Tau80c]. Property- [MT53b, MT52a]. propriétés [Tau53b]. provided [ST35]. Pseudo [Tau31d, TT54, TT56b]. Pseudo- [Tau31d, Tau31d]. Pseudo-Random [TT54, TT56b]. Publications [Ano98c, Ano98e, Ano98f, Lid01, Tauxx]. purpose [ST34]. Pythagoras [Tau88d]. Pythagorean [Tau82a].

Quadratic [Koh84, Tau60b, Tau82c, PT57, PT65, TS34, Tau50a, Tau51c, Tau52c, Tau60d, Tau68a, Tau72a, Tau74a, Tau75b, Tau76a, Tau77a, Tau77c, Tau77b, Tau78b, Tau81a, van76]. quadratischer [TS34]. Quadrupel [Tau31d]. Quadruples [Tau31d]. quantities [Tau35a]. quaternion [ET81, Tau54a]. quaternions [PT57, TT40b, Tau66a]. quelques [Tau53b].

Query [Tau74c, Tau78a]. question [Ser70]. qui [Tau47a].

R [Tau33d]. R. [ABB^{+}48]. Rad [Tau69d]. Radon [Tau87d]. Radons [Tau87d]. Ralston [Tau61e]. Random [TT54, TT56b]. range [ST80, ST80]. rank [Tau77b]. Rational [Tau82d, DT65, TT51, Tau60c, Tau75a, Tau75b, Tau76b, Tau80b, Tau86b]. Real [LT60, Tau83a]. Received [Ano54a, Ano78]. rechnerische [ST34, TS34]. recollection [Tau81c]. Recollections [Tau88f, Tau97]. recurrent [DRTW64]. Recurring [Tau49b]. reference [Tau35b]. reflections [Tau88b]. Régulier [Tau47a]. related [Tau84]. relation [Tau75b, Tau77b]. relations [Gai68, Gai75, TW62a, Tau68a, Tau76a, Tau77c]. Relay [Frö53]. Remark [SN07, ST56, Tau37a, Tau37b, Tau49c, GT79, Tau61d, Tau69c, Tau70b, Zas61]. Remarks [ET81, LT60, Red65, Tau66d, Tau85b, Tau71d, Tau73h, Tau79d, Tau87c]. remembered [WS87]. Remembering [Dav97, Dav05]. Remembrances [Tau87e, Tau88g]. reminiscences [Sch98a, Sch98b]. Renewal [Hla97b]. Report [Fal71, Cas02]. Representation [CR62]. Representations [EG84, MT52b, PT57, Tau74a, Tau80e, Tau81e]. Representing [Tau82d, Tau76b]. Republication [Tau77d]. Research [AB71, TT85b, Tau54f, Tau54e, Tau58g, Tau60e, Tau65b, Tau70c, Tau73f, Tau73g, Tau81d, TG81, Tau85c, TF85, Tau73i, Tau85d]. residues [PT65]. Response [Tau74c]. Result [Tau71c, MK60, Tau74d]. Results [Tau74f, Tau80, TH80, DT65, Tau72e]. Review [Ben17, Cas54, Djo68, Fle55, Koh84, TH31, TH32, Tau46, Tau47a, Tau53b, Tau55b, Tau55d, Tau60f, Tau61e, Tau64a, Sha05]. Reviewed [Ano54b]. Reviews [Ano55, Tau70d]. revisited [Bha01, Joh90, JZ95]. ring [EG84]. Rings [Tau36c, Tau36d, FT36, JT35, Tau73e]. river [Tau82e]. Robertson [JZ95, Joh90, Tau57c]. Role [Tau72d, TW63, Tau68e, Tau70a]. Root [Tau60b, Tau60d]. Roots [LT60, Tau48c, Tau49c, Tau51a, ST56, Tau47b, Tau48d, Tau51b, TT51, Tau54a, Tau64c, Tau68e, Tau70a]. Ross [Zas77].

S [Tau61e]. Salzburg [WS87]. same [Tau80c]. satisfied [ST56]. Satzes [Bre56]. Scalar [PT65]. Schiefringe [FT36]. Schiffer [Sch33]. Scholz [LR16a, LR16b, Tau52b, ST35, Tau52b]. Schulz [Sch33]. Schur [Tau33e]. Science [AB71]. Sciences [Tau53b]. Scientific [Ano54b, Tau69d]. Scientists [Ano17]. Search [Tau79b]. Second [Bec13b]. Segments [Tau49c]. Selected [Mic03, Lid01]. semigroup [DTZ61, DTZ62, Tau84]. semigroups [TW63]. Seminumerical [Tau70d]. September [GdG54]. sequences [DRTW64, Tau54e]. Serbian [Tau69d]. Series [Bec13a, Bec13b]. Sets [Tau80d]. sharpening [Tau31a]. sheaf [FT36]. Short [TZ69]. similar [Tau31c]. similarities [Tau89a]. Similarity [TZ59, Tau31b, Tau66c, Tau70b, Tau79a]. simple [ZX05]. Simplifications [BL77]. Simultaneous [PT53, Tau89a, Ano54a, Cas54, Fle55]. Sixth [Cur56]. skew [Tau44, TT44, Tod44]. skew-symmetric [Tau44, TT44, Tod44]. Small [Tau74c]. smaller [Tauxx]. Society [Cur56]. Solution [CBR^{+}47, O’L01, STS^{+}52, Sha52, Tau52a, Tau54b, Tau57a, Tau60a, TWS^{+}62, KTB^{+}59, LET^{+}57, LTK^{+}60, Tau33e, UEM^{+}60]. Solutions [Abh71, AH50, BF60, BKM^{+}67, BvL68, BT48, BRM70, BMK48, BEM76, CBR^{+}47, CF48, CGJ60, CTT74, DKB^{+}69, DW70, Erd54, EY61, ES61, Frö53, GLF61, GG74, GC68, HH74, HC63, JS69, KO70, Kee61, KP69, KH69, KRO80, KF60, KS68, Koe60, KMCS63, KMS64, Lam58, LC61, Mac58, Mau73, MB74, MTD68, MD69, MHS82, MT54, MC61, MTP^{+}78, New61, PPdB80, PJR^{+}68, RWA70, RH61, RM60, RM48, SST^{+}68, SR69, STS^{+}52, Sha52, Sha54, SN61, SD63, SRvR69, SK63, SyW54, SAG68, SG68, Tau52a, TE54, TS54, Tau57a, TSE58, TP60, Tau60a, TM61, Tau61a, TWS^{+}62, TM63, TC69, TS70, TG80, TL58, UvL61, VM68, WB63, War64, WW74, WF68, WF69, YCJ63, Yff64, Ano98d, KTB^{+}59, LET^{+}57, LTK^{+}60, ST35, UEM^{+}60]. Some [DT65, Gai66, Laf98, LT60, Sch98a, Sch98b, TT78, Tau37c, TT47, Tau56b, TT60, Tau61f, Tau62b, Tau62f, Tau70e, Tau71d, Tau72e, Tau73h, Tau74e, Tau74f, Tau79d, Tau80e, Tau81e, Tau82b, Tau83b, Tau83c, Tau89b, Sha05, Tau45, Tau54d, Tod45]. Song [Tau82e]. Sophie [Tau47a]. Sourcebook [GC87]. Space [Gai66]. Special [ABR97b, Sha98b, Tau68f, DT64, TZ69, TZ70]. Springer [Tau32a, Tau33a, Tau33b]. square [ST34, ST80, ST80, Tau66a]. Squares [Tau70f, ET81, PT57, Tau71a, Tau80, Tau80c, TH80, Tau81b, Tau88d, Tau69d]. stable [Tau68d]. Standards [Cas54, Lid01]. Standpunkt [Tau64a]. statistics [Tau66d]. Students [Ano98g]. Studies [ABB^{+}48]. Study [Tau72d, Tau68e, Tau70a]. Subfield [Tau60b, Tau60d]. subspaces [Tau78e]. Substitutions [Tau47a]. suggested [New80]. sum [PT57]. Sums [Tau70f, ET81, Tau71a, Tau80, Tau80c, TH80, Tau81b, Tau88d, Tra92]. supersonic [Tau45, Tau48a, Tau48b, Tod45]. Supt [Cas54]. Survey [Tod62]. Symmetric [Tau72d, Tau44, Tau72b, Tau73b, Tau73c, Tau73a, Tau73d, Tau83a, Tau89a, TT06, TT44, Tod44]. Symmétrique [Tau47a]. symmetrizability [Bha01]. Symmetrization [Tau50c]. Symposia [Sto11]. Symposion [Sta88]. Symposium [AB71, Cur56, Mey56, PT53, SS83]. Systems [Frö53, TTT57, TT52a, TT52b, Tau54b, TT57b, TT12b, Tau36a, Tau50c, TT57a, TT12a].

T [Tau58d]. Tables [Ano55]. Taussky [Ano54a, Ano98b, Ano98d, Ano98e, Ano98f, Ano98g, Ano15, Bre56, CV76, Cas02, Cas54, Dav97, Dav05, Djo68, Fal71, Gai68, Gai75, Hla97b, JZ95, LR16d, Sha97, Sha98b, Sha05, ZX05, Alz91, ABR97a, ABR97b, ABR98, Bau98, BH54, Bha01, Bre56, Du00, EG84, Fle55, Fri81, Goo20, Hla97a, Joh90, JF98, Kai12, Kis97, Kis98, Koh84, Laf98, LR16a, LR16b, LR16c, LR16d, LV66, Luc87, LM96, Lun94, MK60, MT75, McL96, New80, Oli95, Red65, RS71, Rom91, SN07, Sch77, Sch98a, Sch98b, ST35, Sch33, Ser70, Ste84, TT77, TF85, Ter80, Tho83, Tra92, Uhl07, Wei96, Yin94, Zas61, Zas77, Zun00, van76]. Taussky-Todd [Ano15, ABR97a, ABR97b, ABR98, Bau98, Goo20, Hla97a, Kai12, Kis97, Kis98, Luc87, LM96, MT75, McL96, New80, Oli95, SN07, Sch77, Sch98a, Sch98b, TT77, Ter80, Tra92, Wei96, Zas77, Zun00, van76]. teacher [Uhl07]. Technology [Lid01]. Ternary [Tau82d, Tau82c, Tau76b, Koh84]. Testing [TT54, TT56b]. Texas [TWHM81]. th [Tau60b, Tau60d]. Their [Sto11, ST34, TS34, Tau68e, Tau70a, Tau75b, Tau76a, Tau77c, Tau77b]. Theorem [TW48a, Tau49b, Tau61c, Tau69b, Tau71b, BH54, Bha01, Bre56, Du00, Fri81, Han98, SN07, ST80, ST80, Tau25, TW48b, Tau48e, Tau49a, Tau58d, Tau61d, Tau66d, Tau69c, Tau73e, Tau85b, Tau88d, Tau25, Tau30a, Tau31a, Tau32d]. Theorems [Fal71, TT48a, TT48b, TT49]. Theorie [Tau32e, Tau87d]. Theorists [Sch77]. Theory [CR62, DB31, Sch77, Tau65a, Tau70e, Tau78d, Tau79b, Tau79c, Tau88e, Zas77, AB71, DTZ62, JF98, Kis97, Kis98, Sha97, Sha05, Tau32b, Tau32e, Tau36b, Tau45, Tau48a, Tau48b, Tau54d, Tau56b, Tau62f, Tau65b, Tau74b, Tau78c, Tau82b, Tau83b, Tau83c, Tau87b, Tau87d, Tau88a, Tau88c, Tau89b, Tod45, Tau55a]. third [TWHM81]. Three [Tau72f, ET81, Tau54c, Tau65b, Tau80, Tau80c, TH80]. tightening [Tau30a, Tau32d]. Todd [Ano15, Bau07, Alb07, Alz91, Ano98b, Ano98d, Ano98e, Ano98f, Ano98g, ABR97a, ABR97b, ABR98, Bau98, CV76, Cas02, Dav97, Dav05, Goo20, Hla97a, Hla97b, Kai12, Kis97, Kis98, Luc87, LM96, Lun94, Mad12, MT75, McL96, New80, Oli95, SN07, Sch77, Sch98a, Sch98b, Sha97, Sha98b, Sha05, TT77, Tau87c, Ter80, Tra92, Wei96, Yin94, Zas77, ZX05, Zun00, van76]. Toeplitz [TT06]. Told [Car00]. topics [Tau62b]. Topographer [Ben17, Bre14]. topological [Tau36e, Tau37c]. topologischen [Tau36e]. Torchbearer [Tau88e, Uhl07]. Tournès [Ben17]. Tower [Tau37a, ST34, TS34]. Trace [Tau62d, Tau62e]. Transcript [Tro73]. transform [Tau80b]. Transformation [TZ59, Tau66c, Tau83a]. Transformations [Gai66, TT56a]. transformed [Tau80d]. Transition [Tau74f, Tau72e]. Transpose [TZ59, Tau66c, Tau80b]. travaux [BGC05]. treated [Tau68a]. Tree [JB57, JB12]. triangles [Tau82a]. triangular [Tau80d, TT06]. Tribute [BS81, Sch77]. Turing [Tau56a]. Two [BL77, Djo68, Tau46, Tau67b, Tau73i, Tau76b, Tau82d, Tau82e, Tau85d, Alz91, Tau60e, Tau72b, Tau73b, Tau73c, Tau73a, Tau73d, van76]. Two-Electron [BL77].

u [Tau69d]. U.S. [Cas02]. unification [JF98]. uniform [Tau45, Tau48a, Tau48b, Tod45]. Unimodular [Tau55e, NT57]. unipotent [GT79]. unitary [Tau61b]. units [DT64]. Unity [Tau60b, Tau60d]. University [Mey56, TWHM81]. unknown [Tau76c]. Unramified [Tau37b]. Unsolved [Rom91]. uses [Tau82b]. Uvodenje [Tau69d].

V [Sch33, Tau60f]. valeurs [Tau53b]. value [Tau45, Tau48a, Tau48b, Tau68a, Tod45]. variable [TT60]. Variation [OT83, OT51, Tau64c]. various [Tau64c, Tauxx]. Vector [Gai66]. Verein [ST35]. Vernunft [Sta88]. Veröffentlichungen [Tauxx]. Verschärfung [Tau30a, Tau31a, Tau32d]. Verschiedene [Tauxx]. version [Tau77a]. Vertriebene [Sta88]. VI [Tau69d]. via [Tau81a, Tau86b, Tau88d]. vie [BGC05]. viewpoint [Tau64a]. Villars [Tau53b]. vitae [Ano98b]. Vogel [Tod97]. Vol [Ost83]. Volume [Tau70d, HHOT94]. vom [Tau64a, Tau25].

W [Sch33]. Waerden [TH31, TH32, Tau33b]. Wars [Tau46]. Washington [Cas54, Tau96]. weak [Tau59]. Werk [LR16d]. Werke [Tau32c, Tau33d]. Whereabouts [Ano50]. which [Tau61b, Tau80b, Tau80d]. whose [Tau80c]. Wielandt [Gai68, Gai75, Tau96]. Wien [Sta88]. Wiley [Tau55d]. Wilf [Tau61e]. Wirtinger [FTT55b]. Wissenschaft [Sta88, Tau33b]. Wolfgang [Tau64a]. Women [Cas02, CL05, GC87, Tau88b]. Words [TT88, TT05]. Work [BS81, LR16d, Tau69d, Bre06a, Bre06b, Kis97, Kis98, Laf98, Sha05, Tau74e]. works [Bre11].

years [HHOT94]. York [Tau55d]. Young [Tau69d].

Z [Tau46]. Zahlentheorie [DB31, Tau33c, Tau32a]. Zahlkörper [ST34, TS34, Tau32d]. Zassenhaus [Fal71, MK60]. Zbirovi [Tau69d, Tau69d]. Zeitzeugin [Tau88h]. Zero [Tau62d, Tau82d, Tau62e, Tau76b, Tau84]. zum [Tau25, Tau52b]. zur [Tau87d, Tau30b, Tau31e, Tau32e, Tau33g, Tau34b, Tau36e]. Zweiter [Tau33a].

References

Albers:1985:MPP

[AA85] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical People: Profiles and Interviews. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1985. ISBN 0-8176-3191-7, 3-7643-3191-7. xvi + 2 + 372 pp. LCCN QA28 .M37 1985.

Atkin:1971:CNT

[AB71] Arthur Oliver Lonsdale Atkin and Bryan John Birch, editors. Computers in number theory: proceedings of the Science Research Council Atlas Symposium no. 2 held at Oxford, from 18–23 August, 1969. Academic Press, New York, USA, 1971. ISBN 0-12-065750-3. LCCN QA241 .S35 1969.

Artin:1948:SEP

[ABB^{+}48] Emil Artin, Reinhold Baer, Paul Bernays, Herbert Busemann, Ernest D. Courant, William Feller, W. Fenchel, Jens Nielsen, James Franck, Kurt O. Friedrichs, Fritz John, Hans Lewy, Otto Neugebauer, W. Prager, Hans Rademacher, Hans J. Reissner, Hans Schwerdtfeger, Carl Ludwig Siegel, Olga Taussky, and Alexander Weinstein. Studies and Essays Presented to R. Courant on his 60th Birthday, January 8, 1948. Interscience Publishers, Inc., New York, NY, USA, 1948. 470 pp.

Abhyankar:1971:PSS

[Abh71] S. Abhyankar. Problems and solutions: Solutions of advanced problems: 5693. American Mathematical Monthly, 78(1):86–88, January 1971. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DKB^{+}69].

Aschbacher:1997:OTT

[ABR97a] Michael Aschbacher, Don Blasius, and Dinakar Ramakrishnan, editors. Olga Taussky-Todd: in Memoriam. Pacific Journal of Mathematics, Berkeley, CA, USA, 1997. ISSN 0030-8730 (print), 1945-5844 (electronic). i– vi + 1–357 pp. Pacific J. Math. 1997, Special Issue.

Aschbacher:1997:SIM

[ABR97b] Michael Aschbacher, Don Blasius, and Dinakar Ramakrishnan. Special issue in memory of Olga Taussky-Todd — preface. Pacific Journal of Mathematics, 181 (3):1–357, December 1997. CO-DEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (elec-

tronic). URL https://msp.org/
pjm/1997/181-3/.

Aschbacher:1998:OTT

[ABR98] Michael Aschbacher, Don Blasius, and Dinakar Ramakrishnan, editors. Olga Taussky-Todd: in Memoriam. International Press, Cambridge, MA, USA and Hong Kong, 1998. ISBN 1-57146-051-9 (hardcover). ISSN 0030-8730 (print), 1945-5844 (electronic). iv + 357 pp. LCCN QA7 .O43 1997. URL https://

www.intlpress.com/site/pub/
pages/books/items/00000092/

index.html. Special issue of the Pacific Journal of Mathematics, December 1997.

Anning:1950:EPS

[AH50] Norman Anning and Vern Hoggatt. Elementary problems and solutions: Solutions: E774. American Mathematical Monthly, 57(7):484–486, August/September 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CBR^{+}47].

Albers:2007:JTN

[Alb07] Don Albers. John Todd — numerical mathematics pioneer. College Mathematics Journal, 38 (1):2–23, January 2007. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (elec-

tronic).  URL http://www.
tandfonline.com/doi/abs/10.
1080/07468342.2007.11922213.

Alzer:1991:CTI

[Alz91] Horst Alzer. Converses of two inequalities of Ky Fan, Olga Taussky, and John Todd. Journal of Mathematical Analysis and Applications, 161(1):142–147, October 1991. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Anonymous:1950:W

[Ano50]  Anonymous.  Whereabouts.

Physics Today, 3(3):37, March 1950. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). From the article: “Two British mathematicians, Mr. and Mrs. John Todd, have joined the staff of the National Bureau of Standards’ applied mathematics laboratories. Mr. Todd, formerly of Kings College, London University, and Queens University in Belfast, has been appointed chief of the Bureau’s computation laboratory, where he will be in charge of mathematical research as well as of the computing services offered by the laboratory to various government agencies. Mrs. Todd (Dr. Olga Taussky), as a technical consultant, will do fundamental research in problems of mathematical analysis, algebra, and number theory in their relation to the use of high-speed computing machines”.

Anonymous:1954:BRSa

[Ano54a] Anonymous. Books received: Simultaneous Linear Equations and the Determination of Eigenvalues, edited by L. J. Paige and Olga Taussky; . . .. Physics Today, 7(4):27, April 1954. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Anonymous:1954:BRSb

[Ano54b] Anonymous. Books reviewed in The Scientific Monthly. Science, 119(3101):804–805, June 1954.

CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1954:NB

[Ano54c] Anonymous. New books. Science, 119(3101):805, June 1954. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1955:RDT

[Ano55] Anonymous. Reviews and descriptions of tables and books. Mathematical Tables and Other Aids to Computation, 9(51): 126–136, July 1955. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

                                        URL http://www.jstor.org/
                                        stable/2002069.

Anonymous:1978:BR

[Ano78] Anonymous. Books received. Science, 200(4345):1043, June 1978. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Anonymous:1998:A

[Ano98a] Anonymous. Acknowledgements. Linear Algebra and its Applications, 280(1):64, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:CVO

[Ano98b] Anonymous. [Curriculum vitae of] Olga Taussky Todd. Linear Algebra and its Applications, 280 (1):62–63, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:NPO

[Ano98c] Anonymous. New publications offered by the AMS. Notices of the American Mathematical Society, 45(10):1379–1391, November 1998. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

www.ams.org/notices/199810/
newpubs-nov98.pdf.

Anonymous:1998:OTT

[Ano98d] Anonymous. Olga Taussky Todd problems and solutions. Linear Algebra and its Applications, 280 (1):84–85, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:PAO

[Ano98e] Anonymous. Publications about Olga Taussky Todd. Linear Algebra and its Applications, 280 (1):59–61, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:1998:POT

[Ano98f] Anonymous. Publications of Olga Taussky Todd. Linear Algebra and its Applications, 280(1): 67–83, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Special issue in memory of Olga Taussky Todd.

Anonymous:1998:SOT

[Ano98g] Anonymous. Students of Olga Taussky Todd. Linear Algebra and its Applications, 280(1): 65–66, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Anonymous:2015:OTT

[Ano15] Anonymous. Olga Taussky-Todd, Prof. Dr. Dr. h.c. 30.8.1906– 7.10.1995. Web site, 2015. URL

                                        http://geschichte.univie.ac.
                                        at/en/persons/olga-taussky-
                                        todd-prof-dr-dr-hc.

Anonymous:2017:NGD

[Ano17] Anonymous. NIST gallery of distinguished scientists, engineers, and administrators — 2017. Web document, 2017. URL https://

                                        www.nist.gov/sites/default/
                                        files/documents/2017/11/03/
                                        portrait_gallery_honorees_
                                        through_2016.pdf. Olga Taussky

Toddy (at NBS 1947–1957) was inducted into the Gallery in 2002.

Bauer:1998:OTT

[Bau98] Friedrich L. Bauer. Olga Taussky-Todd 30.8.1906–7.10.1995. Linear Algebra and its Applications, 280 (1):9–12, August 30, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Bauer:2007:MJT

[Bau07] Friedrich L. Bauer. In memoriam John Todd (1911–2007). Numerische Mathematik, 108(1):1– 6, November 2007. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Beckenbach:2013:MMEa

[Bec13a] Edwin Ford Beckenbach, editor.

Modern Mathematics for the Engineer: First Series. Dover books on engineering. Dover, New York, NY, USA, 2013. ISBN 0-486-49746-1 (paperback), 0-486-31611-4 (e-book). xx + 514 pp. LCCN QA401 .B4 2013. With an introduction by Royal Weller. Reprint of [?].

Beckenbach:2013:MMEb

[Bec13b] Edwin Ford Beckenbach, editor.

Modern Mathematics for the Engineer: Second Series. Dover, New York, NY, USA, 2013. ISBN 0-486-31612-2 (e-book), 0-486-49747-X (paperback), 1-306-32668-0 (e-book). xviii + 456 pp. LCCN QA401 .B39 2013. With an introduction by Magnus R. Hestenes. Reprint of [?].

Bumby:1976:PSS

[BEM76] Richard Bumby, Erik Ellentuck, and Dan Marcus. Problems and solutions: Solutions of advanced problems: 5589. American Mathematical Monthly, 83(2):141–142, February 1976. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

Benzi:2017:BRA

[Ben17] Michele Benzi. Book review: André-Louis Cholesky: Mathematician, Topographer and Army Officer by Claude Brezinski and Dominique Tournès. The Mathematical Intelligencer, 39 (2):99–101, April 2017. CO-DEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Bateman:1960:APS

[BF60] P. T. Bateman and N. J. Fine. Advanced problems and solutions: Solutions: 4847. American Mathematical Monthly, 67(3): 299, March 1960. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KTB^{+}59].

Brezinski:2005:VTA

[BGC05] C. Brezinski and M. Gross-Cholesky. La vie et les travaux d’André Cholesky. Bulletin de la Société des Amis de la Bibliothèque de l’École Polytechnique, 39(??):7–32, ???? 2005. ISSN 0989-3059 (print), 2114-2130 (electronic). URL http://

                                        www.sabix.org/bulletin/sabixb39.
                                        htm.

Bellman:1954:TOT

[BH54] Richard Bellman and Alan Hoffman. On a theorem of Ostrowski and Taussky. Archiv der Mathematik, 5:123–127, 1954. CO-DEN ACVMAL. ISSN 0003-889X (print), 1420-8938 (electronic).

Bhaskar:2001:TTS

[Bha01] Atul Bhaskar. Taussky’s theorem, symmetrizability and modal analysis revisited. Proceedings of the Royal Society A: Mathematical, Physical, and Engineering Sciences, 457(2014):2455– 2480, 2001. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Behrmann:1967:PSA

[BKM^{+}67] Adriano Behrmann, M. S.

Klamkin, W. A. McWorter, W. G. Spohn, Jr., C. M. Petty, W. E. Johnson, M. S. Klamkin, Benjamin Volk, M. V. Subbarao, S. W. Williams, and Charles Green. Problems and solutions: Advanced problems: 5460–5469. American Mathematical Monthly, 74(2):206–208, February 1967. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BvL68, GC68, KS68, MTD68, PJR^{+}68, SAG68, SG68, VM68, WF68].

Beebe:1977:STE

[BL77] Nelson H. F. Beebe and Jan Linderberg. Simplifications in the two-electron integral array in molecular calculations. International Journal of Quantum Chemistry, 12(4):683–705, October 1977. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Browne:1948:EPS

[BMK48] D. H. Browne, Norman Miller, and J. B. Kelly. Elementary problems and solutions: E772. American Mathematical Monthly, 55 (2):98–99, February 1948. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CBR^{+}47].

Brenner:1956:NBS

[Bre56] J. L. Brenner. Neuer Beweis eines Satzes von Taussky und Geiringer. (German) [New proof of a theorem by Taussky and Geiringer]. Archiv der Mathematik, 7:274–275, 1956. CO-DEN ACVMAL. ISSN 0003-889X (print), 1420-8938 (electronic).

Brezinski:1996:ALC

[Bre96] Claude Brezinski. André Louis Cholesky. In Meinguet [Mei96], pages 45–50. LCCN QA297 .N825 1996.

Brezinski:2005:MC

[Bre05] Claude Brezinski. La méthode de Cholesky. (French) [The method of Cholesky]. Revue d’Histoire des Math’ematiques = Journal for the History of Mathematics, 11(2):205–238 (2006), 2005. ISSN 1262-022X (print), 1777-568X (electronic). URL http://

                                        smf4.emath.fr/Publications/
                                        RevueHistoireMath/11/html/
                                        smf_rhm_11_205-238.html.

Brezinski:2006:LWAa

[Bre06a] Claude Brezinski. The life and work of André Cholesky. Report, Laboratoire Paul Painlevé, UMR CNRS 8524, UFR de Mathématiques Pures et Appliquées, Université des Sciences et Technologies de Lille, 59655-Villeneuve d Ascq cedex, France, December 18, 2006. 10 pp. URL

http://math.univ-lille1.fr/
~brezinsk/cholNUMA.pdf.

Brezinski:2006:LWAb

[Bre06b] Claude Brezinski. The life and work of André Cholesky. Numerical Algorithms, 43(3):279– 288, November 2006. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic).  URL http://www.
springerlink.com/openurl.asp?
genre=article&issn=1017-1398&
volume=43&issue=3&spage=279.

Brezinski:2011:ALC

[Bre11] Claude Brezinski. André Louis Cholesky: family, life and works. Lecture slides., September 21,

2011. URL http://bugs.unica.
it/SC2011/slides/slides/brezinski.
pdf.

Brezinski:2014:ALC

[Bre14] Claude Brezinski. André-Louis Cholesky: mathematician, topographer and Army officer. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-08134-9 (hardcover). xiii + 311 + 134 pp. LCCN ????

Bond:1970:PSS

[BRM70] John Bond, Harold Reiter, and J. C. Morgan, II. Problems and solutions: Solutions of advanced problems: 5691. American Mathematical Monthly, 77(8):

894, October 1970. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DKB^{+}69].

Brewer:1981:ENT

[BS81] James W. Brewer and Martha Kate Smith, editors. Emmy Noether: a Tribute to Her Life and Work, volume 69 of Monographs and Textbooks in Pure and Applied Mathematics. Marcel Dekker, Inc., New York, NY, USA, 1981. ISBN 0-8247-1550-0. x + 180 pp. LCCN QA29.N6 E47.

Best:1942:CG

[BT42] Ernest Best and Olga Taussky. A class of groups. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 47:55–62, 1942. CODEN PRIAAK. ISSN 0035-8975.

Boas:1948:PSE

[BT48] R. P. Boas, Jr. and Olga Taussky (Mrs. John Todd). Problems and solutions: Elementary problems: Solutions: E775. American Mathematical Monthly, 55(2): 99, February 1948. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CBR^{+}47].

Behrmann:1968:PSS

[BvL68] Adriano Behrmann and J. H. van Lint. Problems and solutions: Solutions of advanced problems: 5460. American Mathematical Monthly, 75(2):202, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Beckenbach:1983:GII

[BW83] Edwin F. Beckenbach and Wolfgang Walter, editors. General Inequalities 3: 3rd International Conference on General Inequalities Oberwolfach, April 26– May 2, 1981, volume 64 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1983. ISBN 3-7643-1539-3. LCCN QA295 .I57 1981. URL https:/

/link.springer.com/book/10.
1007/978-3-0348-6290-5.

Carnes:2000:ANB

[Car00] Mark C. (Mark Christopher)

Carnes, editor. American National Biography Online: the Life of a Nation is Told by the Lives of its People. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2000. ISBN 0-19-520635-5 (24-volume set), 0-19-860669-9 (online). [various] pp. LCCN CT213. URL http:

//www.anb.org.

Cassity:1954:BRS

[Cas54] C. R. Cassity. Book review: Simultaneous Linear Equations and the Determination of Eigenvalues. L. J. Paige and Olga Taussky, Eds. National Bureau of Standards, Washington, 1953. (Order from Supt. of Documents, GPO, Washington, D.C.) 126 pp. Illus. $1.50. Science, 120(3112):302, August 20, 1954. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Case:2002:FGR

[Cas02] Bettye Anne Case. Final grant report to the U.S. Department of Energy (DOE): Olga Taussky Todd celebration of careers in mathematics for women. Report, Office of Scientific and Technical Information (OSTI), June 3, 2002.

Carter:1947:EPS

[CBR^{+}47] C. C. Carter, D. H. Browne, A. L. Rubinoff, Norman Anning, and R. P. Boas. Elementary problems and solutions: Problems for solution: E771–E775. American Mathematical Monthly, 54(5):280–281, May 1947. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AH50, BT48, BMK48, CF48, RM48].

Carter:1948:EPS

[CF48] C. C. Carter and N. J. Fine. Elementary problems and solutions: E771. American Mathematical Monthly, 55(2):98, February 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CBR^{+}47].

Cheney:1960:APS

[CGJ60] Ward Cheney, Allen Goldstein, and Robert C. James. Advanced problems and solutions: Solutions: 4850. American Mathematical Monthly, 67(4):386–387, April 1960. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KTB^{+}59].

Case:2005:CWM

[CL05] Bettye Anne Case and Anne M.

Leggett, editors. Complexities: Women in Mathematics. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (e-book). xix + 412 + 22 pp. LCCN QA27.5 .C66 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0668/2004048843-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0668/
2004048843-d.html;   http:/
/www.loc.gov/catdir/toc/fy0612/
2004048843.html. Based on the

conference “1999 Olga Taussky Todd Celebration of Careers in Mathematics for Women”.

Condon:1958:HP

[CO58] Eward Uhler Condon and Hugh Odishaw, editors. Handbook of Physics. McGraw-Hill, New York, NY, USA, 1958. [various] pp. LCCN QC21 .C74 1958.

Condon:1967:HP

[CO67] Eward Uhler Condon and Hugh Odishaw, editors. Handbook of Physics. McGraw-Hill, New York, NY, USA, second edition, 1967. ISBN 0-07-012403-5. [various] pp. LCCN QC21 .C7 1967.

Cohn:1978:CIA

[Coh78] Harvey Cohn. A Classical Invitation of Algebraic Numbers and Class Fields. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. ISBN 1-4612-9950-0 (e-book), 0-387-90345-3 (print), 3-540-90345-3 (print). xiii + 328 pp. LCCN QA247 .C63.

URL http://catdir.loc.gov/
catdir/enhancements/fy0902/
78013785-d.html;   http://
catdir.loc.gov/catdir/enhancements/
fy0902/78013785-t.html;   http:
//www.gbv.de/dms/hbz/toc/ht000071512.

pdf. With two appendices by Olga Taussky: ”Artin’s 1932 Göttingen lectures on class field theory” and ”Connections between algebraic number theory and integral matrices.”.

Cohn:1988:CIA

[Coh88] Harvey Cohn. A Classical Invitation to Algebraic Numbers and Class Fields. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-90345-3. xiii + 328 pp. LCCN QA247 .C63. With two appendices by Olga Taussky: ‘Artin’s 1932 Göttingen lectures on class field theory’ and ‘Connections between algebraic number theory and integral matrices’.

Curtis:1962:RTF

[CR62] Charles W. Curtis and Irving Reiner. Representation Theory of Finite Groups and Associative Algebras, volume 11 of Pure and applied mathematics. Interscience Publishers, New York, NY, USA, 1962. ISBN 0-470-18975-4. 685 pp. LCCN QA171 .C86.

Cohn:1974:PSS

[CTT74] Paul Cohn and Olga Taussky-Todd. Problems and solutions: Solutions of advanced problems: 5921. American Mathematical Monthly, 81(8): 913, October 1974. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://www.jstor.org/
stable/2319465.  See also

[Mau73].

Curtiss:1956:NAP

[Cur56] J. H. Curtiss, editor. Numerical analysis. [Proceedings of the Sixth Symposium in Applied Mathematics of the American Mathematical Society], volume 6 of Proceedings of Symposia in Applied Mathematics. McGraw-Hill, New York, NY, USA, 1956. LCCN QA297 .S9 1953.

Carlson:1976:CAD

[CV76] David H. Carlson and Richard S.

Varga, editors. Collection of Articles Dedicated to Olga Taussky Todd. American Elsevier Publishing Co., New York, NY, USA, 1976. iv + 172

pp.  URL https://ac.els-
cdn.com/0024379576900501/1-
s2.0-0024379576900501-main.

pdf. Linear Algebra and Appl. 13 (1976), no. 1–2.

Davis:1997:ROT

[Dav97] Chandler Davis. Remembering Olga Taussky Todd. The Mathematical Intelligencer, 19(1):15– 17, 1997. CODEN MAINDC.

ISSN 0343-6993 (print), 1866-7414 (electronic).

Davis:2005:ROT

[Dav05] Chandler Davis. Remembering Olga Taussky Todd. In Case and Leggett [CL05], pages 6– 7. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (e-book). LCCN QA27.5 .C66 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0668/2004048843-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0668/
2004048843-d.html;   http:/
/www.loc.gov/catdir/toc/fy0612/
2004048843.html. Based on the

conference “1999 Olga Taussky Todd Celebration of Careers in Mathematics for Women”.

Dickson:1931:EZG

[DB31] L. E. Dickson and E. Bodewig. Einführung in die Zahlentheorie. (German) [Introduction to Number Theory]. B. G. Teubner, Leipzig, Germany, 1931. ???? pp.

Dick:1970:ENa

[Dic70a] Auguste Dick. Emmy Noether, 1882–1935, volume 13 of Elemente der Mathematik. Beih.; Kurze Mathematiker-Biographien. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1970. ISBN 3-7643-0519-3. 72 (est.) pp. LCCN QA29 .N6 .D513.

Dick:1970:ENb

[Dic70b] Auguste Dick. Emmy Noether, 1882–1935. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1970.

ISBN 3-7643-3019-8. 72 (est.) pp. LCCN QA29 .N6 .D513. Translation by H. I. Blocher from German original.

Dick:1981:EN

[Dic81] Auguste Dick. Emmy Noether, 1882–1935. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1981. ISBN 3-7643-3019-8. xiv + 193 + 4 pp. LCCN QA29.N6

D513.  URL http://link.
springer.com/10.1007/978-1-

4684-0535-4. Translated by H. I. Blocher.

Djokovic:1968:BRT

[Djo68]  D. Z. Djoković.  Book re-

view: On Two Matrices (Olga Taussky). SIAM Review, 10 (3):379, ???? 1968. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

DeMeyer:1969:PSA

[DKB^{+}69] Frank DeMeyer, Irving Kaplansky, John Bond, Harold Reiter, Olga Taussky, S. Abhyankar, Jurg Ratz, and Albert Wilansky. Problems and solutions: Advanced problems: 5689–5694. American Mathematical Monthly, 76(8):947–948, October 1969. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Abh71, BRM70, DW70, KO70, RWA70, TS70].

Dade:1964:DRS

[DRTW64] E. C. Dade, D. W. Robinson, Olga Taussky, and M. Ward. Divisors of recurrent sequences.

Journal für die reine und angewandte Mathematik, 214/215: 180–183, 1964. CODEN JR-MAA8. ISSN 0075-4102 (print), 1435-5345 (electronic).

Dade:1964:DOA

[DT64] E. C. Dade and Olga Taussky. On the different in orders in an algebraic number field and special units connected with it. Acta Arithmetica, 9:47–51, 1964. CO-DEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Dade:1965:SNR

[DT65] E. C. Dade and Olga Taussky. Some new results connected with matrices of rational integers. Proceedings of Symposia in Pure Mathematics, 8:78–88, 1965. ISSN 0082-0717. URL http:

                                        //www.ams.org/books/pspum/
                                        008/0184924.

Dade:1961:SIC

[DTZ61] E. C. Dade, Olga Taussky, and Hans Zassenhaus. On the semigroup of ideal classes in an order of an algebraic number field. Bulletin of the American Mathematical Society, 67 (3):305–308, May 1961. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

                                        tronic). URL http://www.ams.
                                        org/journals/bull/1961-67-
                                        03/S0002-9904-1961-10594-6/
                                        S0002-9904-1961-10594-6.pdf.

Dade:1962:TOP

[DTZ62] E. C. Dade, Olga Taussky, and Hans Zassenhaus. On the theory of orders, in particular on the semigroup of ideal classes and genera of an order in an algebraic number field. Mathematische Annalen, 148:31–64, 1962. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Du:2000:GTT

[Du00] Yuhui Du. A generalization of Taussky’s theorem and its applications. Acta Math. Appl. Sin., 23 (1):76–81, 2000. ISSN 0254-3079.

DeMeyer:1970:PSS

[DW70] Frank DeMeyer and C. R. B.

Wright. Problems and solutions: Solutions of advanced problems: 5689. American Mathematical Monthly, 77(9):1016, November 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DKB^{+}69].

Dwyer:1953:EMC

[Dwy53] Paul S. Dwyer. Errors of matrix computations. In L. J. (Lowell J.) Paige and Olga Taussky, editors, Simultaneous Linear Equations and the Determination of Eigenvalues, number 29 in Applied Mathematics Series, pages 49–58. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1953.

Estes:1984:RUR

[EG84] Dennis R. Estes and Robert M.

Guralnick. Representations under ring extensions: Latimer– MacDuffee and Taussky correspondences. Advances in Mathematics, 54:302–313, 1984. CO-DEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic).

Erdos:1954:APS

[Erd54] Paul Erdős. Advanced problems and solutions: Solutions: 4506. American Mathematical Monthly, 61(1):56–57, January 1954. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Sha52].

Erdos:1961:APSb

[ES61] Paul Erdős and Jack Silver. Advanced problems and solutions: Solutions: 4913. American Mathematical Monthly, 68(5):515–516, May 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UEM^{+}60].

Erdelyi:1946:AIP

[ET46] A. Erdélyi and John Todd. Advanced instruction in practical mathematics. Nature, 158(4020): 690–692, November 16, 1946. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Estes:1981:RCS

[ET81] Dennis R. Estes and Olga Taussky. Remarks concerning sums of three squares and quaternion commutator identities. Linear Algebra and its Applications, 35(??):279–285, February 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0024379581902792.

Erdos:1961:APSa

[EY61] Paul Erdős and Koichi Yamamoto. Advanced problems and solutions: Solutions: 4912. American Mathematical Monthly, 68(5):512–515, May 1961. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UEM^{+}60].

Falk:1971:GZT

[Fal71] D. A. Falk. Generalization of Zassenhaus–Taussky–Dade theorems to noncommutative orders — preliminary report. Notices of the American Mathematical Society, 18(1):120–??, 1971. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Fletcher:1955:BRS

[Fle55] A. Fletcher. Book review: Simultaneous Linear Equations and the Determination of Eigenvalues, by L. J. Paige and Olga Taussky. Mathematical Gazette, 39(330):335, December 1955. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL http://www.jstor.org/
stable/3608611.

Friedland:1981:GMT

[Fri81] Shmuel Friedland. A generalization of the Motzkin–Taussky theorem. Linear Algebra and its Applications, 36(??):103–109, March 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0024379581902238.

Froberg:1953:SLS

[Frö53] Carl-Erik Fröberg. Solutions of linear systems on a relay machine. In L. J. (Lowell J.) Paige and Olga Taussky, editors, Simultaneous Linear Equations and the Determination of Eigenvalues, number 29 in Applied Mathematics Series, pages 39–42. National Bureau of Standards, United States Department of Commerce, Washington, DC, 1953.

Furtwangler:1936:SGS

[FT36] Ph. Furtwängler and Olga Taussky. Über Schiefringe. (German) [On sheaf rings]. Sitzungsberichte. Akadamie der Wissenschaften in Wien, Mathematisch-Naturwissenschaftliche Klasse, Abt. 2a (Mathematik, Astronomie, Physik, Meteorologie and Technik), 145(7–8):525, 1936.

Fan:1955:API

[FTT55a] Ky Fan, Olga Taussky, and John Todd. An algebraic proof of the isoperimetric inequality for polygons. Journal of the Washington Academy of Sciences, 45(11): 339–342, November 1955. CO-DEN JWASA3. ISSN 0043-

0439. URL http://www.jstor.
org/stable/24534441.

Fan:1955:DAI

[FTT55b] Ky Fan, Olga Taussky, and John Todd. Discrete analogs of inequalities of Wirtinger. Monatshefte für Mathematik, 59:73– 90, 1955. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Gaines:1966:SGC

[Gai66] Fergus Gaines. Some Generalizations of Commutativity for Linear Transformations on a Finite Dimensional Vector Space. Ph.D. dissertation, California Institute of Technology, Pasadena, CA, USA, 1966.

Gaines:1968:KTW

[Gai68] Fergus Gaines. Kato–Taussky– Wielandt commutator relations. Linear Algebra and its Applications, 1(1):127–138, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0024379568900530.

Gaines:1975:KTW

[Gai75] Fergus Gaines. Kato–Taussky– Wielandt commutator relations and characteristic curves. Pacific Journal of Mathematics, 61: 121–128, 1975. CODEN PJ-MAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

Green:1968:PSS

[GC68] Charles Green and L. D. Crowson. Problems and solutions: Solutions of advanced problems: 5469. American Mathematical Monthly, 75(3):306, March 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Grinstein:1987:WMB

[GC87] Louise S. Grinstein and Paul J. Campbell, editors. Women of Mathematics: a Biobibliographic Sourcebook. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1987. ISBN 0-313-24849-4 (hardcover). xx + 292 pp. LCCN QA28 .W66 1987. With a foreword by Alice Schafer.

Gerretsen:1954:PIC

[GdG54] Johan C. H. Gerretsen and Johannes de Groot, editors. Proceedings of the International Congress of mathematicians 1954: Amsterdam, September 2– 9, volume 1, 2, 3. Erven P. Noordhoff, Groningen, The Netherlands, 1954. LCCN ????

Gersgorin:1931:AEM

[Ger31] S. Geršgorin. Über die Abgrenzung der Eigenwerte einer Matrix. (German) [On the bounding of eigenvalues of a matrix]. Bull. Acad. Sci. Leningrad, 1931 (??):740–754, ???? 1931.

Golomb:1974:APS

[GG74] S. W. Golomb and Robert Gilmer. Advanced problems and solutions: Solutions: 5918. American Mathematical Monthly, 81 (7):789, August/September 1974. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mau73].

Golomb:1961:APS

[GLF61] S. W. Golomb, Joe Lipman, and C. H. Franke. Advanced problems and solutions: Solutions: 4915. American Mathematical Monthly, 68(6):579, June/July 1961. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UEM^{+}60].

Goodstein:2020:OTT

[Goo20]  Judith R. Goodstein.  Olga

Taussky-Todd. Notices of the American Mathematical Society, 67(3):345–353, March 2020. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL https://

www.ams.org/notices/202003/
rnoti-p345.pdf.

Guralnick:1979:RCU

[GT79] R. Guralnick and Olga Taussky.

A remark concerning unipotent matrix groups. Linear and Multilinear Algebra, 7(1):87–89, 1979. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Hanlon:1998:LMT

[Han98] Phil Hanlon. To the Latimer– Macduffee theorem and beyond! Linear Algebra and its Applications, 280(1):21–37, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Hornyak:1963:APS

[HC63] F. M. Hornyak and L. Carlitz. Advanced problems and solutions: Solutions: 5057. American Mathematical Monthly, 70(9):

1018–1020, November 1963. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TWS^{+}62].

Harris:1974:PSS

[HH74] L. A. Harris and Finban Holland. Problems and solutions: Solutions of advanced problems: 5919. American Mathematical Monthly, 81(10):1122–1123, December 1974. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mau73].

Hoffman:1994:YIV

[HHOT94] A. J. Hoffman, A. S. Householder, A. M. Ostrowski, and O. Taussky Todd. 20 years index volume “Linear Algebra and its Applications”. Linear Algebra and its Applications, 200(??): 1–83, ???? 1994. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Hlawka:1997:OTT

[Hla97a] Edmund Hlawka. Olga Taussky-Todd, 1906–1995. Monatshefte für Mathematik, 123(3):189–201, 1997. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Hlawka:1997:RDO

[Hla97b] Edmund Hlawka. Renewal of the doctorate of Olga Taussky Todd. The Mathematical Intelligencer, 19(1):18–20, 1997. CO-DEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Hoffman:1954:NAP

[HNST54] A. J. (Alan Jerome) Hoffman, M. Newman, E. G. Straus, and Olga Taussky. The number of absolute points of a correlation. Bulletin of the American Mathematical Society, 60(2):169, March 1954. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Hoffman:1956:NAP

[HNST56] A. J. (Alan Jerome) Hoffman, M. Newman, E. G. Straus, and Olga Taussky. On the number of absolute points of a correlation. Pacific Journal of Mathematics, 6:83–96, 1956. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:

//projecteuclid.org/euclid.
pjm/1103044243.

Hoffman:2003:NAP

[HNST03] A. J. (Alan Jerome) Hoffman, M. Newman, E. G. Straus, and Olga Taussky. On the number of absolute points of a correlation. In Micchelli [Mic03], pages 36–49. ISBN 981-02-4198-4, 981-279-693-2 (e-book). LCCN QA164 .H64 2003eb. URL http:/

/ebookcentral.proquest.com/
lib/warw/detail.action?docID=
1681706;   http://public.ebookcentral.
proquest.com/choice/publicfullrecord.
aspx?p=1681706;   http://site.
ebrary.com/id/10255838; http:
//www.myilibrary.com?id=194808.

Hoffman:1998:WOD

[Hof98] Alan Hoffman. What Olga did for me. Linear Algebra and its Applications, 280(1):13, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Hahn:1932:LMA

[HT32] H. Hahn and Olga Taussky. Literaturberichte: Moderne Algebra. Monatshefte für Mathematik und Physik, 39(1):A11– A12, 1932. CODEN MMPHA8. ISSN 1812-8076. B. L. van der Waerden, (Unter Benützung von Vorlesungen von E. Artin und E. Noether.) I. Teil. (Die Grundlehren der mathematischen Wissenschaften, Band XXXIII.) J. Springer, Berlin 1930. VIII und 243 S. Preis geh. RM 15,60, geb. RM 17,20.

Hoffman:1953:CNM

[HT53] A. J. (Alan Jerome) Hoffman and Olga Taussky. A characterization of normal matrices. Bulletin of the American Mathematical Society, 59(1):54, January 1953. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Hoffman:1954:CNM

[HT54] Alan J. Hoffman and Olga Taussky. A characterization of normal matrices. Journal of Research of the National Bureau of Standards (1934), 52:17–19, 1954. ISSN 0160-1741.

James:1957:TM

[JB57] Glenn James and E. F. Beckenbach, editors. The Tree of Mathematics. Digest Press, Pacoima, CA, USA, 1957. xvii + 403 pp. LCCN QA37 .J26.

James:2012:TM

[JB12] Glenn James and E. F. Beckenbach, editors. The Tree of Mathematics. Digest Press, Literary Licensing, LLC, 2012. ISBN 1-258-28562-2. xvii + 403 pp. LCCN QA37 .J26.

Johnson:1998:OMT

[JF98] Charles R. Johnson and Shaun Fallat. Olga, matrix theory and the Taussky unification problem. Linear Algebra and its Applications, 280(1):39–49, August 30, 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). With the assistance of Shaun Fallat. Special issue in memory of Olga Taussky Todd.

Johnson:1990:RTI

[Joh90]  Charles R. Johnson.  The

Robertson–Taussky inequality revisited. Linear and Multilinear Algebra, 26(1–2):5–8, 1990. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See erratum [JZ95].

Just:1969:PSS

[JS69] Erwin Just and D. P. Sumner.

Problems and solutions: Solutions of advanced problems: 5584. American Mathematical Monthly, 76(3):315, March 1969. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

Jacobson:1935:LCR

[JT35] Nathan Jacobson and Olga Taussky. Locally compact rings. Proceedings of the National Academy of Sciences of the United States of America, 21:106– 108, 1935. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Johnson:1995:ERT

[JZ95] Charles R. Johnson and Fuzhen Zhang. Erratum: “The Robertson–Taussky inequality revisited”. Linear and Multilinear Algebra, 38(3):281–282, 1995. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Joh90].

Kaiblinger:2012:POT

[Kai12] Norbert Kaiblinger. Progress on Olga Taussky-Todd’s circulant problem. The Ramanujan Journal, 28(1):45–60, May 2012. ISSN 1382-4090 (print), 1572-9303 (electronic).

Keeping:1961:APS

[Kee61] E. S. Keeping. Advanced problems and solutions: Solutions: 4895. American Mathematical Monthly, 68(2):186, February 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LTK^{+}60].

Klamkin:1960:APS

[KF60] M. S. Klamkin and A. B. Farnell. Advanced problems and solutions: Solutions: 4848. American Mathematical Monthly, 67(3):

300, March 1960. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KTB^{+}59].

Kessing:1969:PSSb

[KH69] H. D. Kessing and D. A. Herrero.

Problems and solutions: Solutions of advanced problems: 5586. American Mathematical Monthly, 76(3):315–316, March 1969. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

Kisilevsky:1997:OTT

[Kis97] Hershy Kisilevsky. Olga Taussky-Todd’s work in class field theory. Pacific Journal of Mathematics, 181(3):219–224, ???? 1997. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (elec-

tronic). URL https://msp.org/
pjm/1997/181-3/pjm-v181-n3-

p10-p.pdf. Dedicated “To the memory of Olga Taussky-Todd, a friend, a colleague and an inspiration”.

Kisilevsky:1998:OTT

[Kis98] Hershy Kisilevsky. Olga Taussky-Todd’s work in class field theory. In Aschbacher et al. [ABR98], pages 219–224. CODEN PJ-MAAI. ISBN 1-57146-051-9 (hardcover). ISSN 0030-8730 (print), 1945-5844 (electronic). LCCN QA7 .O43 1997. URL

https://www.intlpress.com/
site/pub/pages/books/items/
00000092/index.html. Special

issue of the Pacific Journal of Mathematics, December 1997.

Kohls:1963:APS

[KMCS63] C. W. Kohls, M. E. Mahowald, C. H. Cunkle’s Class in Topology, and E. L. Spitznagel, Jr. Advanced problems and solutions: Solutions: 5056. American Mathematical Monthly, 70(9): 1017–1018, November 1963. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KMS64, TWS^{+}62].

Kohls:1964:PSS

[KMS64] C. W. Kohls, M. E. Mahowald, and P. S. Schnare. Problems and solutions: Solutions of advanced problems: 5056. American Mathematical Monthly, 71(5): 562, May 1964. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KMCS63, TWS^{+}62].

Kaplansky:1970:PSS

[KO70] Irving Kaplansky and Barbara Osofsky. Problems and solutions: Solutions of advanced problems: 5690. American Mathematical Monthly, 77(8):893–894, October 1970. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DKB^{+}69].

Koehler:1960:APS

[Koe60] Fulton Koehler. Advanced problems and solutions: Solutions: 4845. American Mathematical Monthly, 67(2):190–192, February 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KTB^{+}59].

Koh:1984:RTQ

[Koh84] Kwangil Koh. Review of Ternary Quadratic Forms and Norms, edited by Olga Taussky. Linear Algebra and its Applications, 57(??):291–293, February 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379584901952.

Kessing:1969:PSSa

[KP69] H. D. Kessing and Nicholas Passell. Problems and solutions: Solutions of advanced problems: 5585. American Mathematical Monthly, 76(3):315, March 1969. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

Kim:1980:PSS

[KRO80] Ki Hang Kim, Fred Roush, and R. W. K. Odoni. Problems and solutions: Solutions of advanced problems: 6215. American Mathematical Monthly, 87(4): 309, April 1980. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [MTP^{+}78].

Klamkin:1968:PSS

[KS68] M. S. Klamkin and Seymour Schuster. Problems and solutions: Solutions of advanced problems: 5461. American Mathematical Monthly, 75(2):202–203, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Kani:1988:CPH

[KS88] Ernst J. Kani and Robert A. (Robert Arnold) Smith, editors. The Collected Papers of Hans Arnold Heilbronn. Canadian Mathematical Society series of monographs and advanced texts = Monographies et études de la Société mathématique du Canada. Wiley, New York, NY, USA, 1988. ISBN 0-471-81655-8. xvi + 608 + 1 pp. LCCN QA241 .H372513 1988.

Kato:1956:C

[KT56] T. Kato and Olga Taussky. Commutators of A and A∗. Journal of the Washington Academy of Sciences, 46(2):38–40, February 1956. CODEN JWASA3. ISSN 0043-0439. URL http://www.

                                        jstor.org/stable/24533630.

Koehler:1959:APS

[KTB^{+}59] Fulton Koehler, Olga Taussky, P. T. Bateman, M. S. Klamkin, L. A. Rubel, Ward Cheney, and Allen Goldstein. Advanced Problems and Solutions: Problems for Solution: 4845-4850. American Mathematical Monthly, 66 (5):426–427, May 1959. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BF60, CGJ60, KF60, Koe60, RM60, TP60].

Laffey:1998:SAO

[Laf98] Thomas J. Laffey. Some aspects of Olga Taussky’s work in algebra. Linear Algebra and its Applications, 280(1):51–57, August 30, 1998. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Lambek:1958:APS

[Lam58] Joachim Lambek. Advanced problems and solutions: 4723. American Mathematical Monthly, 65(1):48–49, January 1958. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LET^{+}57].

Lipman:1961:APS

[LC61] Joe Lipman and Leonard Carlitz. Advanced problems and solutions: Solutions: 4893. American Mathematical Monthly, 68(2): 183–185, February 1961. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LTK^{+}60].

Lambek:1957:APS

[LET^{+}57] Joachim Lambek, Paul Erdős, Olga Taussky, Victor Thébault, and G. R. MacLane. Advanced Problems and Solutions: Problems for Solution: 4723-4727. American Mathematical Monthly, 64(2):116–117, 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Lide:2001:CEM

[Lid01] D. R. Lide, editor. A Century of Excellence in Measurements, Standards, and Technology: A Chronicle of Selected NBS/NIST Publications, 1901– 2000, volume 958. National Technical Information Service, Washington, DC, USA, 2001. ix + 386 pp. URL https:/

/nvlpubs.nist.gov/nistpubs/
sp958-lide/cntsp958old.htm.

NIST Special Publication.

Luchins:1996:MOT

[LM96] Edith H. Luchins and Mary Ann McLoughlin. In memoriam: Olga Taussky-Todd [1906–1995]. Notices of the American Mathematical Society, 43(8):838– 847, August 1996. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://www.ams.
org/journals/notices/199608/
taussky.pdf.

Lemmermeyer:2016:BHS

[LR16a] Franz Lemmermeyer and Peter Roquette, editors. Der Briefwechsel Hasse–Scholz–Taussky. (German) [Correspondence: Hasse– Scholz–Taussky]. Universitätsverlag Göttingen, Göttingen, Germany, 2016. ISBN 3-86395-253-7. xxii

+ 574 pp. URL http://d-nb.
info/1090794738/04.

Lemmermeyer:2016:KAS

[LR16b] Franz Lemmermeyer and Peter Roquette. Die Korrespondenz Arnold Scholz–Olga Taussky. (German) [Correspondence: Arnold Scholz–Olga Taussky]. In Der Briefwechsel Hasse–Scholz–Taussky. (German) [Correspondence: Hasse– Scholz–Taussky] [LR16a], chapter 4, pages 341–450. ISBN 3-86395-253-7. URL http://d-nb.

info/1090794738/04.

Lemmermeyer:2016:KEK

[LR16c] Franz Lemmermeyer and Peter Roquette. Die Korrespondenz Elisabeth Korrodi–Olga Taussky. (German) [Correspondence: Elisabeth Korrodi–Olga Taussky]. In Der Briefwechsel Hasse–Scholz–Taussky. (German) [Correspondence: Hasse– Scholz–Taussky] [LR16a], chapter 5, pages 451–462. ISBN 3-86395-253-7. URL http://d-nb.

info/1090794738/04.

Lemmermeyer:2016:OTL

[LR16d] Franz Lemmermeyer and Peter Roquette. Olga Taussky: Leben und Werk. (German) [Olga Taussky: Life and work]. In Der Briefwechsel Hasse–Scholz– Taussky. (German) [Correspondence: Hasse–Scholz–Taussky] [LR16a], chapter 2, pages 47–70. ISBN 3-86395-253-7. URL http:

//d-nb.info/1090794738/04.

Lewis:1960:SRC

[LT60] Daniel C. Lewis, Jr. and Olga Taussky. Some remarks concerning the real and imaginary parts of the characteristic roots of a finite matrix. Journal of Mathematical Physics, 1(3): 234–236, March 1960. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic).  URL http://jmp.
aip.org/resource/1/jmapaq/
v1/i3/p234_s1.

Lipman:1960:APS

[LTK^{+}60] Joe Lipman, Olga Taussky, E. S.

Keeping, H. S. Shapiro, D. J.

Newman, and R. C. Reed. Advanced Problems and Solutions: Problems for Solution: 4893-4898. American Mathematical Monthly, 67(3):294–295, 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Luchins:1987:OTT

[Luc87] Edith H. Luchins. Olga Taussky-Todd (1906–). In Grinstein and Campbell [GC87], pages 225–235. ISBN 0-313-24849-4 (hardcover). LCCN QA28 .W66 1987. With a foreword by Alice Schafer.

Lunter:1994:NPG

[Lun94] Gerton Lunter. New proofs and a generalization of inequalities of Fan, Taussky, and Todd. Journal of Mathematical Analysis and Applications, 185(2): 464–476, July 15, 1994. CO-DEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Levinger:1966:POT

[LV66] B. W. Levinger and R. S. Varga. On a problem of Olga Taussky. Pacific Journal of Mathematics, 19:473–487, 1966. CO-DEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.pjm/1102993718.

MacLane:1958:APS

[Mac58] G. R. MacLane. Advanced problems and solutions: Solutions: 4727. American Mathematical Monthly, 65(1):50, January 1958.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LET^{+}57].

Madden:2012:JTD

[Mad12] Niall Madden. John Todd and the development of modern numerical analysis. Irish Mathematical Society Bulletin, 69: 11–24, Summer 2012. ISSN 0791-5578. URL https:/

/www.maths.tcd.ie/pub/ims/
bull69/Madden.pdf.

Mauldon:1973:PSA

[Mau73] J. G. Mauldon. Problems and solutions: Advanced problems: 5916. American Mathematical Monthly, 80(6):697–698, June/ July 1973. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CTT74, GG74, HH74, MB74, WW74].

Mauldon:1974:PSS

[MB74] J. G. Mauldon and Robert Breusch. Problems and solutions: Solutions of advanced problems: 5916. American Mathematical Monthly, 81(8):911–912, October 1974. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mau73].

Moser:1961:APS

[MC61] Leo Moser and Leonard Carlitz. Advanced problems and solutions: Solutions: 4914. American Mathematical Monthly, 68 (5):516–517, May 1961. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UEM^{+}60].

McLoughlin:1996:OTT

[McL96] Mary Ann Schultz McLoughlin. Olga Taussky-Todd: Grande dame of mathematics. PhD thesis, Rensselaer Polytechnic Institute, Troy, NY, USA, 1996. vii + 203 + reproductions

                                        pp.  URL https://search.
                                        proquest.com/docview/304269254.

Thesis (Ph.D.)–Rensselaer Polytechnic Institute.

Mehr:1969:APS

[MD69] C. B. Mehr and W. G. Dotson, Jr. Advanced problems and solutions: Solutions: 5588. American Mathematical Monthly, 76 (3):316–317, March 1969. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

Meinguet:1996:NAN

[Mei96] Jean Meinguet, editor. Numerical analysis: a numerical analysis conference in honour of Jean Meinguet, Bulletin of the Belgian Mathematical Society, Simon Stevin. Belgian Mathematical Society, Brussels, Belgium, 1996. LCCN QA297 .N825 1996.

Meyer:1956:SMC

[Mey56] Herbert A. Meyer, editor. Symposium on Monte Carlo Methods: held at the University of Florida, March 16 and 17, 1954. Wiley, New York, NY, USA, 1956. LCCN QA273 U577.

Mendez:1982:PSS

[MHS82] C. G. Mendez, James P. Hatzenbuhler, and Michael Smith. Problems and solutions: Solutions of advanced problems: 6213. American Mathematical Monthly, 89(4): 279, April 1982. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [MTP^{+}78].

Hoffman:2003:SPA

[Mic03] Charles A. Micchelli, editor. Selected papers of Alan Hoffman with commentary. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2003. ISBN 981-02-4198-4, 981-279-693-2 (ebook). xliv + 446 pp. LCCN QA164 .H64 2003eb. URL http:/

/ebookcentral.proquest.com/
lib/warw/detail.action?docID=
1681706;   http://public.ebookcentral.
proquest.com/choice/publicfullrecord.
aspx?p=1681706;   http://site.
ebrary.com/id/10255838; http:
//www.myilibrary.com?id=194808.

Marcus:1960:CRT

[MK60] Marvin Marcus and N. A. Khan.

On a commutator result of Taussky and Zassenhaus. Pacific Journal of Mathematics, 10: 1337–1346, 1960. CODEN PJ-MAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

Motzkin:1952:MP

[MT52a] T. S. Motzkin and Olga Taussky.

Matrices with property-L. Bulletin of the American Mathematical Society, 58(1):41, January 1952. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Motzkin:1952:RFG

[MT52b] T. S. Motzkin and Olga Taussky. On representations of finite groups. Nederlandse Akademie van Wetenschappen. Proceedings. Series A. Indagationes Mathematicae, 55:511–512, 1952. ISSN 0023-3358.

Motzkin:1952:PMP

[MT52c] T. S. Motzkin and Olga Taussky. Pairs of matrices with property L. Transactions of the American Mathematical Society, 73:108–114, 1952. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (elec-

                                        tronic). URL http://www.ams.
                                        org/journals/tran/1952-073-
                                        01/S0002-9947-1952-0049855-
                                        8/S0002-9947-1952-0049855-
                                        8.pdf.

Motzkin:1953:PMPb

[MT53a] T. S. Motzkin and Olga Taussky. Pairs of matrices with property L. Proceedings of the National Academy of Sciences of the United States of America, 39:961– 963, 1953. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Motzkin:1953:PMPa

[MT53b] T. S. Motzkin and Olga Taussky. Pairs of matrices with property-L. II. Bulletin of the American Mathematical Society, 59(2):155, March 1953. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Moser:1954:APS

[MT54] Leo Moser and Bryant Tuckerman. Advanced problems and solutions: 4507. American Mathematical Monthly, 61(2):126–127, February 1954. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Sha52].

Motzkin:1955:PMP

[MT55] T. S. Motzkin and Olga Taussky.

Pairs of matrices with property L. II. Transactions of the American Mathematical Society, 80:387–401, 1955. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (elec-

tronic). URL http://www.ams.
org/journals/tran/1955-080-
02/S0002-9947-1955-0086781-
5/S0002-9947-1955-0086781-
5.pdf.

Marcus:1975:CAD

[MT75] Marvin Marcus and Robert C.

Thompson, editors. Collection of Articles Dedicated to Olga Taussky-Todd. Gordon and Breach, New York, NY, USA / London, UK / Paris, France, 1975. 1–168 pp. Linear and Multilinear Algebra 3 (1975/76), no. 1–2.

McWorter:1968:PSS

[MTD68] W. A. McWorter, E. J. Taft, and J. E. Delany. Problems and solutions: Solutions of advanced problems: 5462. American Mathematical Monthly, 75(2):203–204, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Mullin:1978:PSA

[MTP^{+}78] Albert A. Mullin, Olga Taussky, Alvin J. Paullay, Sidney Penner, A. A. Mullin, C. G. Mendez, Leonard Carlitz, Ki Hang Kim, and Fred Roush. Problems and solutions: Advanced problems: 6183, 6210–6215. American Mathematical Monthly, 85 (5):389–390, May 1978. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KRO80, MHS82, PPdB80, TG80].

Newman:1961:APS

[New61] D. J. Newman. Advanced problems and solutions: Solutions: 4897. American Mathematical Monthly, 68(2):187, February 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LTK^{+}60].

Newman:1980:PSO

[New80] Morris Newman. On a problem suggested by Olga Taussky-Todd. Illinois Journal of Mathematics, 24(1):156–158, 1980. CO-DEN IJMTAW. ISSN 0019-2082 (print), 1945-6581 (electronic).

URL http://projecteuclid.
org/euclid.ijm/1256047802.

Neurath:1945:K

[NT45] Otto Neurath and Olga Taussky.

Korrespondenz. Catalog entry in ÖNB (Austrian National Library) for papers in the estate of Otto and Marie Neurath., October 9, 1945.

URL https://search.onb.ac.
at/primo_library/libweb/action/
dlDisplay.do?docId=ONB_alma21303349700003338.

Newman:1956:GNB

[NT56] M. Newman and Olga Taussky.

On a generalization of the normal basis in abelian algebraic number fields. Communications on Pure and Applied Mathematics (New York), 9:85–91, 1956. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Newman:1957:CPD

[NT57] Morris Newman and Olga Taussky.

Classes of positive definite unimodular circulants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):71–73, ???? 1957. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

OLeary:2001:IMS

[O’L01]  Dianne P. O’Leary.  Itera-

tion method for the solution of the eigenvalue problem of linear differential and integral operators. In Lide [Lid01], pages 77–80. URL https:/

/nvlpubs.nist.gov/nistpubs/
sp958-lide/077-080.pdf. NIST

Special Publication.

Oliver:1995:OTT

[Oli95] Myrna Oliver. Olga Taussky-Todd; noted mathematician. Los Angeles Times, ??(??):A44, December 3, 1995. URL http://

                                        ezproxy.lib.utah.edu/docview/
                                        421082715.

Ostrowski:1983:CMP

[Ost83] Alexander M. Ostrowski, editor. Collected Mathematical Papers. Vol. 1. I Determinants II Linear Algebra III Algebraic Equations. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1983. ISBN 3-0348-9355-8. ???? pp.

                                        LCCN ???? URL http://www.
                                        springerlink.com/content/978-
                                        3-0348-9355-8.

Ostrowski:1951:VDP

[OT51] Alexander M. Ostrowski and Olga Taussky. On the variation of the determinant of a positive definite matrix. Nederlandse Akademie van Wetenschappen. Proceedings. Series A. Indagationes Mathematicae, 54: 383–385, 1951. ISSN 0023-3358.

Ostrowski:1983:VDP

[OT83] Alexander M. Ostrowski and Olga Taussky. On the variation of the determinant of a positive definite matrix. In Ostrowski [Ost83], pages 75–77. ISBN 3-0348-9355-8. LCCN ???? URL http://www.

                                        springerlink.com/content/978-
                                        3-0348-9355-8.

Ostrowski:1949:B

[OTT49] Alexander M. Ostrowski, Olga Taussky Todd, and John Todd. [Bibliography]. Report ????, U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1949. ???? pp.

Petty:1968:PSS

[PJR^{+}68] C. M. Petty, W. E. Johnson, P. V. Subba Rao, B. Ramachandra Rao, and Olga Taussky. Problems and solutions: Solutions of advanced problems: 5464. American Mathematical Monthly, 75(2): 205–206, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Paullay:1980:APS

[PPdB80] Alvin J. Paullay, Sidney Penner, and Ingrid de Buda. Advanced problems and solutions: Solutions: 6211. American Mathematical Monthly, 87(3):229, March 1980. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [MTP^{+}78].

Price:1978:DDP

[Pri78] G. Baley Price. Determinants with dominant principal diagonal: a personal historical note. Linear and Multilinear Algebra, 6 (4):247–250, 1978. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See comments [TT78].

Paige:1953:SLE

[PT53] L. J. (Lowell J.) Paige and Olga Taussky, editors. Simultaneous Linear Equations and the Determination of Eigenvalues[: proceedings of a symposium held 23– 25 August 1951 at the Institute for Numerical Analysis, Los Angeles, California], volume 29 of Applied Mathematics Series. U.S. National Bureau of Standards, Gaithersburg, MD, USA, 1953. LCCN QA3 .U5 no. 29.

Pall:1957:AQR

[PT57] G. Pall and Olga Taussky. Application of quaternions to the representations of a binary quadratic form as a sum of four squares. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 58: 23–28, 1957. CODEN PRIAAK. ISSN 0035-8975.

Pall:1965:SMQ

[PT65] Gordon Pall and Olga Taussky. Scalar matrix quadratic residues. Mathematika, 12:94–96, 1965. CODEN MTKAAB. ISSN 0025-5793 (print), 2041-7942 (electronic).

Pall:1970:FCN

[PT70] Gordon Pall and Olga Taussky. Factorization of Cayley numbers. Journal of Number Theory, 2(1):74–90, February 1970. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0022314X70900077.

Quinn:1983:ENB

[QMLT83] Grace S. Quinn, Ruth S. Mc-Kee, Marguerite Lehr, and Olga Taussky. Emmy Noether in Bryn Mawr. In Srinivasan and Sally [SS83], pages 139–146. ISBN 0-387-90838-2. LCCN QA1 .E52 1983.

Redheffer:1965:RPT

[Red65] Raymond Redheffer. Remarks on a paper of Taussky. Journal of Algebra, 2:42–47, 1965. CO-DEN JALGA4. ISSN 0021-8693 (print), 1090-266x (electronic).

Reed:1961:APS

[RH61] R. C. Reed and A. P. Hillman. Advanced problems and solutions: Solutions: 4898. American Mathematical Monthly, 68(2): 187–188, February 1961. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LTK^{+}60].

Rubinoff:1948:EPS

[RM48] A. L. Rubinoff and Leo Moser.

Elementary problems and solutions: E773. American Mathematical Monthly, 55(2):99, February 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [CBR^{+}47].

Rubel:1960:APS

[RM60] L. A. Rubel and Brockway McMillan. Advanced problems and solutions: Solutions: 4849. American Mathematical Monthly, 67(4):385–386, April 1960. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KTB^{+}59].

Romine:1991:UPT

[Rom91] Charles H. Romine. On an unsolved problem of Olga Taussky. Johns Hopkins APL Technical Digest, 12(3):271–274, July/ September 1991. CODEN JHADDQ. ISSN 0270-5214 (print), 1930-0530 (electronic).

URL http://www.jhuapl.edu/
techdigest/views/pdfs/V12_
N3_1991/V12_N3_1991_Romine.
pdf.

Robinson:1971:GPM

[RS71] Norman F. Robinson and M. F. Smiley. A generalization of the property L of Motzkin and Taussky. Linear Algebra and its Applications, 4(4): 323–328, October 1971. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379571900036.

Ratz:1970:PSS

[RWA70] Jurg Ratz, Albert Wilansky, and D. R. Anderson. Problems and solutions: Solutions of advanced problems: 5694. American Mathematical Monthly, 77(8): 895, October 1970. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DKB^{+}69].

Spohn:1968:PSS

[SAG68] W. G. Spohn, Jr., A. S. Adikesavan, and H. W. Gould. Problems and solutions: Solutions of advanced problems: 5463. American Mathematical Monthly, 75(2):

204–205, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Schur:1933:AGJ

[Sch33] I. Schur. Aufgabe 128 (gestellt in Jahresbericht D. M. V. 41 (1932), 67 kursiv). Lösungen von J. Fox & W. Schulz, M. Schiffer und Olga Taussky. Jahresbericht der Deutschen Mathematiker-Vereinigung, 42:117–120, 1933. CODEN JDMVA7. ISSN 0012-0456 (print), 1869-7135 (electronic).

Schneider:1977:OTT

[Sch77] Hans Schneider. Olga Taussky-Todd’s influence on matrix theory and matrix theorists: A discursive personal tribute. Linear and Multilinear Algebra, 5(3): 197–224, 1977. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Schneider:1998:SPRa

[Sch98a] Hans Schneider. Some personal reminiscences of Olga Taussky-Todd. arxiv.org, ??(??):????, March 14, 1998. URL https:/

/arxiv.org/abs/math/9803061.

Published in [Sch98b].

Schneider:1998:SPRb

[Sch98b] Hans Schneider. Some personal reminiscences of Olga Taussky-Todd. Linear Algebra and its Applications, 280(1):15–19, August 30, 1998. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Special issue in memory of Olga Taussky Todd.

Shapiro:1963:APS

[SD63] H. S. Shapiro and Dragomir Djokovic. Advanced problems and solutions: Solutions: 5053. American Mathematical Monthly, 70(8):904–906, October 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TWS^{+}62].

Schneider:1981:HAM

[SDGT81] Hans Schneider, Chandler Davis, Walter Gautschi, and Olga Taussky Todd. Honoring Alexander M. Ostrowski. Linear Algebra and its Applications, 38(??):3, June 1981. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379581900021.

Serre:1970:QOT

[Ser70] Jean-Pierre Serre. Sur une question d’Olga Taussky. (French) [On a question of Olga Taussky]. Journal of Number Theory, 2 (2):235–236, May 1970. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0022314X70900235.

Subbarao:1968:PSS

[SG68] M. V. Subbarao and M. G. Greening. Problems and solutions: Solutions of advanced problems: 5467. American Mathematical Monthly, 75(2):207, February 1968. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Shapiro:1952:APSb

[Sha52] Harold Shapiro. Advanced problems and solutions: Problems for solution: 4503. American Mathematical Monthly, 59(8):554–555, October 1952. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd54, MT54, Sha54, SyW54, TE54, TS54].

Shapiro:1954:APS

[Sha54] H. S. Shapiro. Advanced problems and solutions: 4503. American Mathematical Monthly, 61(1): 54, January 1954. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Sha52].

Shapiro:1997:NMO

[Sha97] Helene Shapiro. Notes from Math 223: Olga Taussky Todd’s matrix theory course, 1976– 1977. The Mathematical Intelligencer, 19(1):21–27, 1997. CO-DEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Shapiro:1998:I

[Sha98a] Helene Shapiro. Introduction.

Linear Algebra and its Applications, 280(1):5–8, August 1998. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/S0024379598100022.

Shapiro:1998:SIM

[Sha98b] Helene Shapiro. Special issue in memory of Olga Taussky Todd. Linear Algebra and its Applications, 280(1):86, 1998. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Shapiro:2005:NMC

[Sha05] Helene Shapiro. Numbers, matrices, and commutativity: a review of some of Olga Taussky Todd’s work in matrix theory. In Case and Leggett [CL05], pages 292– 307. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (e-book). LCCN QA27.5 .C66 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0668/2004048843-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0668/
2004048843-d.html;   http:/
/www.loc.gov/catdir/toc/fy0612/
2004048843.html. Based on the

conference “1999 Olga Taussky Todd Celebration of Careers in Mathematics for Women”.

Sinkhorn:1963:APS

[SK63] Richard Sinkhorn and A. G. Konheim. Advanced problems and solutions: Solutions: 5054. American Mathematical Monthly, 70 (8):906–907, October 1963. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TWS^{+}62].

Shapiro:1961:APS

[SN61] H. S. Shapiro and D. J. Newman. Advanced problems and solutions: Solutions: 4896. American Mathematical Monthly, 68(2): 186–187, February 1961. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LTK^{+}60].

Sai:2007:RTT

[SN07] Rintaro Sai and Hiroshi Nakazato.

Remark on a theorem of Taussky-Todd. Bulletin of the Graduate School of Science and Technology. Hirosaki University, 1:1–4, 2007. ISSN 1882-2509 (electronic).

Schumm:1969:PSS

[SR69] G. F. Schumm and A. L. Rubin.

Problems and solutions: Solutions of advanced problems: 5587. American Mathematical Monthly, 76(3):316, March 1969. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

Shirey:1969:PSS

[SRvR69] J. E. Shirey, Charles Riley, and Michael von Renteln. Problems and solutions: Solutions of advanced problems: 5581. American Mathematical Monthly, 76 (3):313–314, March 1969. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

Srinivasan:1983:ENB

[SS83] Bhama Srinivasan and Judith D.

Sally, editors. Emmy Noether in Bryn Mawr: Proceedings of a Symposium. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-90838-2. LCCN QA1 .E52 1983.

Schumm:1968:PSA

[SST^{+}68] G. F. Schumm, J. E. Shirey, Olga Taussky, S. W. Williams, Erwin Just, H. D. Keesing, C. B. Mehr, Richard Bumby, and Erik Ellentuck. Problems and solutions: Advanced problems: 5580–5589. American Mathematical Monthly, 75(4):414–415, April 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BEM76, JS69, KP69, KH69, MD69, SR69, SRvR69, TC69, WF69].

Scholz:1934:HKK

[ST34] Arnold Scholz and Olga Taussky. Die Hauptideale der kubischen Klassenkörper imaginärquadratischer Zahlkörper: ihre rechnerische Bestimmung und ihr Einfluß auf den Klassenkörperturm. (German) [The main ideals of cubic class-objects of imaginary-square number-fields: their mathematical purpose and their influence on the class-body-tower]. Journal für die reine und angewandte Mathematik, 171:19–41, 1934. CODEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic).

Scholz:1935:AGI

[ST35] Arnold Scholz and Olga Taussky. Aufgabe 208 (gestellt im Jber. Deutsche Math-Verein. 45 (1935), 110 kursiv). Lösungen von A. Scholz und Olga Taussky. (German) [Exercise 208 (provided in the German Mon German Math Association 45 (1935), 110 italics). Solutions by A. Scholz and Olga Taussky]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 47:47, 1935. CO-DEN JDMVA7. ISSN 0012-0456 (print), 1869-7135 (electronic).

Straus:1956:RPP

[ST56] E. G. Straus and Olga Taussky.

Remark on the preceding paper. Algebraic equations satisfied by roots of natural numbers. Pacific Journal of Mathematics, 6: 97–98, 1956. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:

//projecteuclid.org/euclid.
pjm/1103044244.

Shapiro:1980:APT

[ST80] H. Shapiro and Olga Taussky. Alternative proofs of a theorem of Moyls and Marcus on the numerical range of a square matrix. Linear and Multilinear Algebra, 8(4): 337–340, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). field of values.

Shapiro:1979:APT

[ST80] H. Shapiro and Olga Taussky. Alternative proofs of a theorem of Moyls and Marcus on the numerical range of a square matrix. Linear and Multilinear Algebra, 8(4): 337–340, 1979/80. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Stadler:1988:VVE

[Sta88] Friedrich Stadler. Vertriebene Vernunft: Emigration und Exil österreichischer Wissenschaft, Internationale Symposion, 19. bis

23. Oktober 1987 in Wien, volume 2 of Vertriebene Vernunft. Jugend und Volk, Wien, Austria, 1988. ISBN 3-224-16525-1. 1117 pp. LCCN ????

Stetsenko:1984:DCO

[Ste84] V. Ya. Stetsenko. On a development of a criterion of Olga Taussky for nonsingularity of an indecomposable matrix. Izv. Akad. Nauk Tadzh. SSR, Otd. Fiz.-Mat. Khim. Geol. Nauk, 1984(1(91)):3–8, 1984. ISSN 0002-3485.

Stoer:2011:GST

[Sto11] Josef Stoer. The Gatlinburg Symposia and their influence on the development of numerical linear algebra. IMAGE, ??(??):13– 25, Spring 2011. URL http://

                                        www.maths.manchester.ac.uk/
                                        ~higham/conferences/householder/
                                        image46_stoer.pdf. From the

start of §2: “The Gatlinburg Symposia, officially started by Alston S. Householder in 1961, were influenced in part by the existence of two earlier and similar symposia, the first organized by Olga Taussky in 1951 in Los Angeles and the second by Wallace Givens in 1957 in Detroit. ”.

Shapiro:1952:APSa

[STS^{+}52] Harold Shapiro, Olga Taussky, M. R. Spiegel, Paul Erdős, and Leo Moser. Advanced problems and solutions: Problems for solution: 4503–4507. American Mathematical Monthly, 59(8): 554–555, 1952. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Spiegel:1954:APS

[SyW54] M. R. Spiegel and Chih yi Wang.

Advanced problems and solutions: 4505. American Mathematical Monthly, 61(1):55–56, January 1954. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Sha52].

Taussky:1925:BPT

[Tau25] Olga Taussky. Vom binomial zum polynomial Theorem. (German) [From the binomial to the polynomial theorem]. Matura essay, Körnerschule Linz, Linz, Austria, 1925. ???? pp.

Taussky:1930:VHG

[Tau30a] Olga Taussky. Über eine Verschärfung des Hauptidealsatzes. (German) [On a tightening of the Principal Ideal Theorem]. Doctoral, Universität Wien, Vienna, Austria, 1930. URL https:

//www.genealogy.math.ndsu.
nodak.edu/id.php?id=11912.

Taussky:1930:MGG

[Tau30b] Olga Taussky. Zur Metrik der Gruppen. (German) [On the metric of groups]. Anzeiger der Akademie der Wissenschaften in Wien. Philosophisch-Historische Klasse, 67:140–142, 1930.

Taussky:1931:VHG

[Tau31a] Olga Taussky. Eine Verschärfung des Hauptidealsatzes. (German) [A sharpening of the Principal Ideal Theorem]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 40:29–30, 1931. CO-DEN JDMVA7. ISSN 0012-0456 (print), 1869-7135 (electronic).

Taussky:1931:SG

[Tau31b] Olga Taussky. On similarity of groups. Annals of Mathematics, 32(??):739–760, ???? 1931. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). In K. Menger, Some applications of point-set methods.

Taussky:1931:AAG

[Tau31c] Olga Taussky. Über ähnliche Abbildungen von Gruppen. (German) [On similar pictures of groups]. Ergebnisse math. Kolloquium Wien, 3:13–14, 1931.

Taussky:1931:PQG

[Tau31d] Olga Taussky. Über Pseudo-G-Quadrupel. (German) [On pseudo-G-quadruples]. Mathematische Zeitschrift, 33(??):396– 418, 1931. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic). Section 6 of K. Menger, Beiträge zur Gruppentheorie, I. Über eine Gruppenmetrick. (German) [Contributions to Group Theory, I. On a Group Metric].

Taussky:1931:MGG

[Tau31e] Olga Taussky. Zur Metrik der Gruppen. (German) [On the metric of groups]. Ergebnisse math. Kolloquium Wien, 2:38, 1931.

Taussky:1932:EDH

[Tau32a] Olga Taussky. Editor of David Hilbert Gesammelte Abhandlungen, Erster Band, ZahlenTheorie, J. Springer, Berlin, 1932.

Taussky:1932:EAL

[Tau32b] Olga Taussky. Editor to E.

Artin’s lectures on class field theory, 1932.

Taussky:1932:LGM

[Tau32c] Olga Taussky. Literaturberichte:

Gesammelte mathematische Werke. Monatshefte für Mathematik und Physik, 39(1):A9, 1932. CO-DEN MMPHA8. ISSN 1812-8076. Richard Dedekind, Herausgegeben von R. Fricke †, E. Noether, Ö. Ore. II. Band. 442 Seiten. Friedrich Vieweg & Sohn, Braunschweig 1931. Preis geh. RM 40,50, geb. 43.

Taussky:1932:VHA

[Tau32d] Olga Taussky. Über eine Verschärfung des Hauptidealsatzes für algebraische Zahlkörper. (German) [On a tightening of the Principal Ideal Theorem for algebraic number fields]. Journal für die reine und angewandte Mathematik, 168:193–210, 1932. CO-DEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic).

Taussky:1932:TKG

[Tau32e] Olga Taussky. Zur Theorie des Klassenkörpers. (German) [On the theory of the class body]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 41: 74, 1932. CODEN JDMVA7. ISSN 0012-0456 (print), 1869-7135 (electronic).

Taussky:1933:EDH

[Tau33a] Olga Taussky. Editor of David Hilbert Gesammelte Abhandlungen, Zweiter Band, Algebra, Invariantentheorie, Geometrie, J. Springer, Berlin, 1933.

Taussky:1933:LMA

[Tau33b] Olga Taussky. Literaturberichte: Moderne Algebra, B. L. van der Waerden. Bd. II. (Grundlehren der mathematischen Wissenschaft in Einzeldarstellungen. Bd. 34.) J. Springer, Berlin 1931. Preis geb. RM 15,-. Monatshefte für Mathematik und Physik, 40(1):A3–A4, 1933. CODEN MMPHA8. ISSN 1812-8076.

Taussky:1933:LEZ

[Tau33c] Olga Taussky. Literaturberichte: Einführung in die Zahlentheorie. Monatshefte für Mathematik und Physik, 40(1):A2, 1933. CODEN MMPHA8. ISSN 1812-8076. L. E. Dickson — E. Bodewig. B. G. Teubner, Leipzig 1931. Preis geb. RM 9,60.

Taussky:1933:LRD

[Tau33d] Olga Taussky. Literaturberichte: R. Dedekind, Gesammelte mathematische Werke. Monatshefte für Mathematik und Physik, 40(1): A20, 1933. CODEN MMPHA8. ISSN 1812-8076. Bd. 3. F. Vieweg u. S., Braunschweig 1932. Preis geh. RM 41,40.

Taussky:1933:LAS

[Tau33e] Olga Taussky. Lösung der Aufgabe #128 (I. Schur, 41 (2. Abt.): 67, 2. Abt.). (German) [Solution of exercise #128]. Jahresbericht der Deutschen Mathematiker-Vereinigung, 42 (2. Abt.)(??): 119–120, ???? 1933. CODEN JD-MVA7. ISSN 0012-0456 (print), 1869-7135 (electronic).

Taussky:1933:IAG

[Tau33f] Olga Taussky. Über isomorphe Abbildungen von Gruppen. (German) [On isomorphic mappings of groups]. Mathematische Annalen, 108(1):615–620, 1933. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Taussky:1933:AGG

[Tau33g] Olga Taussky. Zur Axiomatik der Gruppen. (German) [On the axioms of groups]. Ergebnisse eines math. Kolloquiums Wien, 4:2–3, 1933.

Taussky:1934:IAG

[Tau34a] Olga Taussky. Über isomorphe Abbildungen von Gruppen. (German) [On isomorphic mappings of groups]. Mathematische Annalen, 108:615–620, 1934. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Taussky:1934:AGG

[Tau34b] Olga Taussky. Zur Axiomatik der Gruppen. (German) [On the axioms of groups]. Ergebnisse math. Kolloquium Wien, 4:2–3, 1934.

Taussky:1935:AKM

[Tau35a] Olga Taussky. Abstrakte Körper und Metrik. I: Endliche Mengen und Körperpotenzen. (German) [Abstract bodies and metric. I: Finite quantities and body potencies]. Ergebnisse math. Kolloquium Wien, 6:20–23, 1935.

Taussky:1935:FCF

[Tau35b] Olga Taussky. Fields as class fields of different reference fields. Bulletin of the American Mathematical Society, 41(??):32, ???? 1935. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Taussky:1936:AMH

[Tau36a] Olga Taussky. Analytical methods in hypercomplex systems. Compositio Mathematica, 3:399– 407, 1936. CODEN CMPMAF. ISSN 0010-437x (print), 1570-5846 (electronic). URL http:/

                                        /www.numdam.org/item?id=CM_
                                        1936__3__399_0.

Taussky:1936:MPA

[Tau36b] Olga Taussky. Modern problems in algebraic number theory. British Association for the Advancement of Science, Report, page 332, 1936. CODEN BAASAX. ISSN 0365-8694. URL

                                        https://biodiversitylibrary.
                                        org/page/30739101.

Taussky:1936:RNCa

[Tau36c] Olga Taussky. Rings with noncommutative addition. Bulletin of the Calcutta Mathematical Society, 28:245–246, 1936. CODEN BCMSA5. ISSN 0008-0659.

Taussky:1936:RNCb

[Tau36d] Olga Taussky. Rings with noncommutative addition. Bulletin of the Calcutta Mathematical Society, 28:245–246, 1936. CODEN BCMSA5. ISSN 0008-0659.

Taussky:1936:TAG

[Tau36e] Olga Taussky. Zur topologischen Algebra. (German) [Towards topological algebra]. Ergebnisse math. Kolloquium Wien, 7: 60–61, 1936.

Taussky:1937:RCF

[Tau37a] Olga Taussky. A remark on the class field tower. Journal of the London Mathematical Society, 12 (2):82–85, 1937. CODEN JLM-SAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Taussky:1937:RUC

[Tau37b] Olga Taussky. A remark on unramified class fields. Journal of the London Mathematical Society, 12(2):86–88, 1937. CODEN JLM-SAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Taussky:1937:SPT

[Tau37c] Olga Taussky. Some problems of topological algebra. Comptes Rendus du Congrès International des Mathématiciens, Oslo, 2:31– 32, 1937.

Taussky:1939:APL

[Tau39] Olga Taussky. An algebraic property of Laplace’s differential equation. Quarterly Journal of Mathematics. Oxford series., 10:99– 103, 1939. CODEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Taussky:1943:OPD

[Tau43] Olga Taussky. Obituary: Prof. David Hilbert, For.Mem.R.S. Nature, 152(3850):182–183, August 14, 1943. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https:

//www.nature.com/articles/
152182a0.pdf.

Taussky:1944:NSS

[Tau44] Olga Taussky. A note on skewsymmetric matrices. Technical Report of the Aeronautical Research Committee 1944, 2:923– 940, 1944. Appendix is pp. 938– 940.

Taussky:1945:SBV

[Tau45] Olga Taussky. On some boundary value problems in the theory of the non-uniform supersonic motion of an aerofoil. Technical Report of the Aeronautical Research Council 1945, 1:452–464, 1945.

Taussky:1946:DML

[Tau46] Olga Taussky. Book review: The Development of Mathematical Logic and of Logical Positivism in Poland between the Two Wars, by Z. Jordan. Mathematical Gazette, 30(288):53– 54, February 1946. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL http://www.jstor.org/
stable/3609178.

Taussky:1947:BDG

[Tau47a] Olga Taussky. Book review: Sur les Bases du Groupe Symmétrique et les Couples de Substitutions qui Engendrent un Groupe Régulier, by Sophie Piccard. Mathematical Gazette, 31(295):187, July 1947. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Taussky:1947:MOB

[Tau47b] Olga Taussky. A method for obtaining bounds for characteristic roots of matrices with applications to flutter calculations. Report 10.508, Aeronautical Research Council of Great Britain, 1947. ???? pp.

Taussky:1948:BVPa

[Tau48a] Olga Taussky. A boundary value problem for a hyperbolic differential equation arising in the theory of the non-uniform supersonic motion of an aerofoil. Bulletin of the American Mathematical Society, 54(3):282, March 1948. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1948:BVPb

[Tau48b] Olga Taussky. A boundary value problem for a hyperbolic differential equation arising in the theory of the non-uniform supersonic motion of an aerofoil. In Studies and Essays Presented to R. Courant on his 60th Birthday, January 8, 1948 [ABB^{+}48], pages 421–435.

Taussky:1948:BCRa

[Tau48c] Olga Taussky. Bounds for characteristic roots of matrices. Bulletin of the American Mathematical Society, 54(9):831, September 1948. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1948:BCRb

[Tau48d] Olga Taussky. Bounds for characteristic roots of matrices. Duke Mathematical Journal, 15:1043– 1044, 1948. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        dmj/1077475205.

Taussky:1948:TLM

[Tau48e] Olga Taussky. On a theorem of Latimer and MacDuffee. Bulletin of the American Mathematical Society, 54(11):1057, November 1948. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1949:TLM

[Tau49a] Olga Taussky. On a theorem of Latimer and MacDuffee. Canadian Journal of Mathematics = Journal canadien de mathématiques, 1(??):300–302, ???? 1949. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Taussky:1949:RTD

[Tau49b] Olga Taussky. A recurring theorem on determinants. American Mathematical Monthly, 56(10): 672–676, December 1949. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1949:RCC

[Tau49c] Olga Taussky. A remark concerning the characteristic roots of the finite segments of the Hilbert matrix. Quarterly Journal of Mathematics, 20:80–83, 1949. CO-DEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Taussky:1950:CMQ

[Tau50a] Olga Taussky. Classes of matrices and quadratic fields. In Proceedings of the International Congress of Mathematicians (Cambridge, Mass, 1950), page 329. American Mathematical Society, Providence, RI, USA, 1950. Abstract only.

Taussky:1950:NNA

[Tau50b] Olga Taussky. Notes on numerical analysis. II. Note on the condition of matrices. Mathematical Tables and Other Aids to Computation, 4(30):111–112, May 1950. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (elec-

tronic). URL http://www.ams.
org/journals/mcom/1950-04-
030/S0025-5718-1950-0038137-
8/S0025-5718-1950-0038137-
8.pdf.

Taussky:1950:SSL

[Tau50c] Olga Taussky. Symmetrization of systems of linear equations. Bulletin of the American Mathematical Society, 56(4):342, July 1950. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1951:BBC

[Tau51a] Olga Taussky. Bibliography on bounds for characteristic roots of finite matrices. Report 1162, U.S. National Bureau of Standards, Gaithersburg, MD, USA, September 1951. 10 pp. URL https:/

                                        /nvlpubs.nist.gov/nistpubs/
                                        Legacy/RPT/nbsreport1162.pdf.

Taussky:1951:BCR

[Tau51b] Olga Taussky. Bounds for characteristic roots of matrices. II. Journal of Research of the National Bureau of Standards (1934), 46: 124–125, 1951. ISSN 0160-1741.

Taussky:1951:CMQ

[Tau51c] Olga Taussky. Classes of matrices and quadratic fields. Pacific Journal of Mathematics, 1:127– 132, 1951. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:

                                        //projecteuclid.org/euclid.
                                        pjm/1102613159.

Taussky:1952:APS

[Tau52a] Olga Taussky. Advanced problems and solutions: Problems for solution: 4504. American Mathematical Monthly, 59(8):554–555, October 1952. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Erd54, MT54, Sha54, SyW54, TE54, TS54].

Taussky:1952:ASG

[Tau52b] Olga Taussky. Arnold Scholz zum Gedächtnis. (German) [Arnold Scholz in memoriam]. Mathematische Nachrichten, 7(6):379–386, 1952. CODEN MTMNAQ. ISSN 0025-584X. URL 0049129.

Taussky:1952:CMQ

[Tau52c] Olga Taussky. Classes of matrices and quadratic fields. II. Journal of the London Mathematical Society, 27:237–239, 1952. CO-DEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Taussky:1953:GCM

[Tau53a] Olga Taussky. Generalized commutators of matrices. Bulletin of the American Mathematical Society, 59(1):57, January 1953. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1953:RQP

[Tau53b] Olga Taussky. Review of Sur quelques propriétés des valeurs caractéristiques des matrices carrés, by M. Parodi, (Mémorial des Sciences Mathématiques, no. 118, Gauthier Villars, Paris, 1952). Bulletin of the American Mathematical Society, 59 (5):464, September 1953. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

tronic). URL http://www.ams.

                                        org/journals/bull/1953-59-
                                        05/S0002-9904-1953-09722-1/
                                        S0002-9904-1953-09722-1.pdf.

Taussky:1954:CRQ

[Tau54a] Olga Taussky. Characteristic roots of quaternion matrices. Archiv der Mathematik, 5(1–3): 99–101, April 1954. CODEN ACVMAL. ISSN 0003-889X (print), 1420-8938 (electronic).

Taussky:1954:CSS

[Tau54b] Olga Taussky, editor. Contributions to the Solution of Systems of Linear Equations and the Determination of Eigenvalues, volume 39 of Applied Mathematics Series. U.S. National Bureau of Standards, Gaithersburg, MD, USA, September 30, 1954.

Taussky:1954:GCM

[Tau54c] Olga Taussky. Generalized commutators of matrices and permutations of factors in a product of three matrices. In Friends, Colleagues, and Pupils, editors, Studies in mathematics and mechanics presented to Richard von Mises, pages 67–68. Academic Press, New York, USA, 1954. ISBN 1-4832-3272-7. LCCN QA3 S93.

Taussky:1954:NMS

[Tau54d] Olga Taussky. Normal matrices in some problems in algebraic number theory. In Gerretsen and de Groot [GdG54], pages 67–68. LCCN ????

Taussky:1954:RPM

[Tau54e] Olga Taussky. Research problem 11 (multiply monotonic sequences). Bulletin of the American Mathematical Society, 60 (3):289–290, May 1954. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

tronic). URL http://www.ams.
org/journals/bull/1954-60-
03/S0002-9904-1954-09820-8/
S0002-9904-1954-09820-8.pdf.

Taussky:1954:RPH

[Tau54f] Olga Taussky. Research problem 12 (the Hilbert matrix). Bulletin of the American Mathematical Society, 60(3):290, May 1954. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

tronic). URL http://www.ams.
org/journals/bull/1954-60-
03/S0002-9904-1954-09820-8/
S0002-9904-1954-09820-8.pdf.

Solutions have been found by T. Kato, Proc. AMS 58:78–81 (1957) and Comm. PAM 11:573– 586 (1958) and by M. Rosenblum, Proc. AMS 9:137–140 and 581– 585 (1958).

Taussky:1955:AMM

[Tau55a] Olga Taussky. Abstract of Matrix Methods in Algebraic Number Theory. Pasadena, CA, USA., 1955.

Taussky:1955:BRE

[Tau55b] Olga Taussky. Book review: Economic Activity Analysis. Bulletin of the American Mathematical Society, 61(3):241–244, May 1, 1955. CODEN BAMOAD.

ISSN 0002-9904 (print), 1936-881X (electronic).

Taussky:1955:NGM

[Tau55c] Olga Taussky. A note on group matrices. Proceedings of the American Mathematical Society, 6(6):984–986, June 1955. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (elec-

                                        tronic). URL http://www.ams.
                                        org/journals/proc/1955-006-
                                        06/S0002-9939-1955-0075202-
                                        X/S0002-9939-1955-0075202-
                                        X.pdf.

Taussky:1955:REA

[Tau55d] Olga Taussky. Review of Economic activity analysis, (O. Morgenstern, ed.), Wiley, New York, 1954. Bulletin of the American Mathematical Society, 61(3):241– 243, 1955. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Taussky:1955:UIC

[Tau55e] Olga Taussky. Unimodular integral circulants. Mathematische Zeitschrift, 63:286–289, 1955. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic).

Taussky:1956:AMT

[Tau56a] Olga Taussky. Alan Mathison Turing 1912–1954. Mathematical Tables and Other Aids to Computation, 10(55):180–181, 1956. CODEN MTTCAS. ISSN 0891-6837 (print), 2326-4853 (electronic).

Taussky:1956:SCP

[Tau56b] Olga Taussky. Some computational problems in algebraic number theory. In Curtiss [Cur56], pages 187–193. LCCN QA297

.S9 1953. URL http://www.ams.
org/books/psapm/006/0079337.

Taussky:1957:APS

[Tau57a] Olga Taussky. Advanced problems and solutions: Problems for solution: 4725. American Mathematical Monthly, 64(2):116–117, February 1957. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Lam58, Mac58, TSE58, TL58].

Taussky:1957:CFM

[Tau57b] Olga Taussky. Commutativity in finite matrices. American Mathematical Monthly, 64(4):229–235, April 1957. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1957:DIH

[Tau57c] Olga Taussky. A determinantal inequality of H. P. Robertson. I. Journal of the Washington Academy of Sciences, 47 (8):263–264, August 1957. CO-DEN JWASA3. ISSN 0043-

0439. URL http://www.jstor.
org/stable/24533742.

Taussky:1957:MCC

[Tau57d] Olga Taussky. On matrix classes corresponding to an ideal and its inverse. Illinois Journal of Mathematics, 1:108–113, 1957. CO-DEN IJMTAW. ISSN 0019-2082 (print), 1945-6581 (electronic).

                                        URL http://projecteuclid.
                                        org/euclid.ijm/1255378509.

Taussky:1958:A

[Tau58a] Olga Taussky. Algebra. In Condon and Odishaw [CO58], pages 10–21. LCCN QC21 .C74 1958.

Taussky:1958:IM

[Tau58b] Olga Taussky. Integral matrices. Bulletin of the American Mathematical Society, 64(??):??, ???? 1958. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). Keynote lecture at the AMS Meeting, the first by a woman since Emmy Noether in 1934.

Taussky:1958:NGC

[Tau58c] Olga Taussky. A note on the group commutator of A and A^{∗}. Journal of the Washington Academy of Sciences, 48 (10):305, October 1958. CO-DEN JWASA3. ISSN 0043-

                                        0439. URL http://www.jstor.
                                        org/stable/24533665.

Taussky:1958:MTC

[Tau58d] Olga Taussky. On a matrix theorem of A. T. Craig and H. Hotelling. Nederlandse Akademie van Wetenschappen. Proceedings. Series A. Indagationes Mathematicae, 61:139–141, 1958. ISSN 0023-3358.

Taussky:1958:O

[Tau58e] Olga Taussky. Operators. In Condon and Odishaw [CO58], pages 95–96. LCCN QC21 .C74 1958.

Taussky:1958:ODE

[Tau58f] Olga Taussky. Ordinary differential equations. In Condon and Odishaw [CO58], pages 59–65. LCCN QC21 .C74 1958.

Taussky:1958:RP

[Tau58g] Olga Taussky. Research problems 9, 10, 11. Bulletin of the American Mathematical Society, 64(3 (part 1)):124, May 1958. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

URL https://projecteuclid.
org/euclid.bams/1183522331.

Taussky:1959:WPP

[Tau59] Olga Taussky. A weak property L for pairs of matrices. Mathematische Zeitschrift, 71:463–465, 1959. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic).

Taussky:1960:APSb

[Tau60a] Olga Taussky. Advanced problems and solutions: Problems for solution: 4894. American Mathematical Monthly, 67 (3):294–295, March 1960. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Kee61, LC61, New61, RH61, SN61, TM61].

Taussky:1960:MNQ

[Tau60b] Olga Taussky. Mathematical notes: The quadratic subfield of the field generated by the p-th root of unity. American Mathematical Monthly, 67(8):769, October 1960. CODEN AMMYAE.

ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1960:MRI

[Tau60c] Olga Taussky. Matrices of rational integers. Bulletin of the American Mathematical Society, 66:327–345, 1960. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

tronic). URL http://www.ams.
org/journals/bull/1960-66-
05/S0002-9904-1960-10439-9/
S0002-9904-1960-10439-9.pdf.

Taussky:1960:QSF

[Tau60d] Olga Taussky. The quadratic subfield of the field generated by the p-th root of unity. American Mathematical Monthly, 67: 769, 1960. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1960:RPD

[Tau60e] Olga Taussky. Research problem 2 (discuss the location of the eigenvalues of the Jordan product of two Hermitian matrices). Bulletin of the American Mathematical Society, 66(3):275, May 1960. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Taussky:1960:RCM

[Tau60f] Olga Taussky. Review of Computational Methods of Linear Algebra by V. N. Faddeeva, Dover. Mathematical Gazette, 44(348): 139–140, May 1960. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

URL http://www.jstor.org/
stable/3612579.

Taussky:1961:APSb

[Tau61a] Olga Taussky. Advanced problems and solutions: Solutions: 4916. American Mathematical Monthly, 68(6):579– 580, June/July 1961. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://links.jstor.org/
sici?sici=0002-9890(196106/
07)68:6<579:4>2.0.CO;2-9&origin=

MSN. See also [UEM^{+}60].

Taussky:1961:CUM

[Tau61b] Olga Taussky. Commutators of unitary matrices which commute with one factor. Journal of Mathematics and Mechanics, 10:175– 178, 1961. CODEN JOMMAN. ISSN 0095-9057 (print), 1943-5274 (electronic).

Taussky:1961:GTL

[Tau61c] Olga Taussky. A generalization of a theorem of Lyapunov. Journal of the Society for Industrial and Applied Mathematics, 9(4): 640–643, December 1961. CO-DEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Taussky:1961:RTL

[Tau61d] Olga Taussky. A remark on a theorem of Lyapunov. Journal of Mathematical Analysis and Applications, 2:105–107, 1961. CO-DEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Taussky:1961:RMM

[Tau61e] Olga Taussky. Review of Mathematical Methods for Digital Computers, edited by A. Ralston and H. S. Wilf. Econometrica, 29 (3):487, ???? 1961. CODEN ECMTA7. ISSN 0012-9682.

Taussky:1961:SCP

[Tau61f] Olga Taussky. Some computational problems involving integral matrices. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 65: 15–17, 1961. CODEN JNBBAU. ISSN 0098-8979; 0022-4340.

Taussky:1962:CP

[Tau62a] Olga Taussky. A characterization of property L. Journal de Mathématiques Pures et Appliquées, 41:291–293, 1962. CO-DEN JMPAAM. ISSN 0021-7824 (print), 1776-3371 (electronic).

Taussky:1962:EFM

[Tau62b] Olga Taussky. Eigenvalues of finite matrices: Some topics concerning bounds for eigenvalues of finite matrices. In Todd [Tod62], chapter 8a, pages 279–297. LCCN QA297 .T63. URL https:

                                        //babel.hathitrust.org/cgi/
                                        pt?id=mdp.39015000983901.

Taussky:1962:IM

[Tau62c] Olga Taussky. Ideal matrices. I. Archiv der Mathematik, 13: 275–282, 1962. CODEN ACV-MAL. ISSN 0003-889X (print), 1420-8938 (electronic).

Taussky:1962:MNM

[Tau62d] Olga Taussky. Mathematical notes: Matrices with trace zero. American Mathematical Monthly, 69(1):40–42, January 1962. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1962:MTZ

[Tau62e] Olga Taussky. Matrices with trace zero. American Mathematical Monthly, 69:40–42, 1962. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1962:NTS

[Tau62f] Olga Taussky. Number theory: Some computational problems in algebraic number theory. In Todd [Tod62], chapter 16b, pages 549–557. LCCN QA297 .T63. URL https:

//babel.hathitrust.org/cgi/
pt?id=mdp.39015000983901.

Taussky:1963:IMI

[Tau63]  Olga Taussky.  Ideal matri-

ces. II. Mathematische Annalen, 150:218–225, 1963. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Taussky:1964:BRW

[Tau64a] Olga Taussky. Book review:

Wolfgang Krull, Elementare und klassische Algebra, vom modernen Standpunkt. (German) [Wolfgang Krull, Elementary and classical algebra from a modern viewpoint]. Mathematics of Computation, 18(87):510, July 1964. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Taussky:1964:M

[Tau64b] Olga Taussky. Matrices C with C^{n} → 0. Journal of Algebra, 1: 5–10, 1964. CODEN JALGA4. ISSN 0021-8693 (print), 1090-266x (electronic).

Taussky:1964:VCR

[Tau64c] Olga Taussky. On the variation of the characteristic roots of a finite matrix under various changes of its elements. In Hans Schneider, editor, Recent Advances in Matrix Theory (Proc. Advanced Seminar, Math. Res. Center, U.S. Army, Univ. Wisconsin, Madison, Wis., 1963), pages 125–138. University of Wisconsin Press, Madison, WI, USA, 1964.

Taussky:1965:MT

[Tau65a] Olga Taussky. Matrix theory. Bulletin of the American Mathematical Society, 71 (5):711, September 1965. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1965:RPT

[Tau65b] Olga Taussky. Research problem 17, three problems in matrix theory (Jordan products). Bulletin of the American Mathematical Society, 71(4):711, July 1965. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

                                        tronic). URL http://www.ams.
                                        org/journals/bull/1965-71-
                                        05/S0002-9904-1965-11349-0/
                                        S0002-9904-1965-11349-0.pdf.

Taussky:1966:DIQ

[Tau66a] Olga Taussky. A determinantal identity for quaternions and a new eight square identity. Journal of Mathematical Analysis and Applications, 15:162–164, 1966. CODEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Taussky:1966:IM

[Tau66b] Olga Taussky. Integral matrices. Lectures at National Science Foundation Advanced Science Seminar on Algebraic Number Theory, Bowdoin College, 1966. Notes by Philip Laufer.

Taussky:1966:STB

[Tau66c] Olga Taussky. On the similarity transformation between an integral matrix with irreducible characteristic polynomial and its transpose. Mathematische Annalen, 166:60–63, 1966. CO-DEN MAANA3. ISSN 0025-5831 (print), 1432-1807 (electronic).

Taussky:1966:RMT

[Tau66d] Olga Taussky. Remarks on a matrix theorem arising in statistics. Monatshefte für Mathematik, 70: 461–464, 1966. CODEN MN-MTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Taussky:1967:A

[Tau67a] Olga Taussky. Algebra. In Condon and Odishaw [CO67], pages 15–26. ISBN 0-07-012403-5. LCCN QC21 .C7 1967.

Taussky:1967:TM

[Tau67b] Olga Taussky. On two matrices. SIAM Review, 9(3):590, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). Problem 67-13. Solution by D. Z. Djokovic in SIAM Review 10:379 (1968).

Taussky:1967:ODE

[Tau67c] Olga Taussky. Ordinary differential equations. In Condon and Odishaw [CO67], pages 61–66 (or 95–97??). ISBN 0-07-012403-5. LCCN QC21 .C7 1967.

Taussky:1967:PDM

[Tau67d] Olga Taussky. Positive-definite matrices. In Oved Shisha, editor, Inequalities (Proceedings of the Symposium at Wright-Patterson Air Force Base, Ohio, 1965), pages 309–319. Academic Press, New York, USA, 1967.

Taussky:1968:AGA

[Tau68a] Olga Taussky. Automorphs and generalized automorphs of quadratic forms treated as characteristic value relations. Linear Algebra and its Applications, 1(3):349–356, July 1968. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379568900141.

Taussky:1968:DMA

[Tau68b] Olga Taussky. The discriminant matrices of an algebraic number field. Journal of the London Mathematical Society, 43:

152–154, 1968. CODEN JLM-SAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Taussky:1968:FAF

[Tau68c] Olga Taussky. The factorization of the adjugate of a finite matrix. Linear Algebra and its Applications, 1(1):39–41, January 1968. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379568900463.

Taussky:1968:SM

[Tau68d] Olga Taussky. On stable matrices. In ????, editor, Programmation en mathématiques numériques, (Actes Colloq. Internat. C.N.R.S. No. 165, Besançon, 1966), pages 75–88. Éditions du Centre National de Recherche Scientifique, Paris, France, 1968.

Taussky:1968:PDM

[Tau68e] Olga Taussky. Positive-definite matrices and their role in the study of the characteristic roots of general matrices. Advances in Mathematics, 2(2):175–186, 1968. CODEN ADMTA4. ISSN 0001-8708 (print), 1090-2082 (electronic). Also published in [Tau70a].

Taussky:1968:SPC

[Tau68f] Olga Taussky. Special problems concerning the GL(n, F) and SL(n, F). Lecture at NSF Advanced Science Seminar in Algebraic Groups, Bowdoin College., 1968. Notes by John H. Smith.

Taussky:1969:GMC

[Tau69a] Olga Taussky. A generalization of matrix commutativity. Linear Algebra and its Applications, 2(3):349–353, July 1969. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379569900366.

Taussky:1969:HTI

[Tau69b] Olga Taussky. Hilbert’s theorem 94. II. Notices of the American Mathematical Society, 16(2):407, February 1969. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1969:RCH

[Tau69c] Olga Taussky. A remark concerning Hilbert’s theorem 94. Journal für die reine und angewandte Mathematik, 239–240: 435–438, 1969. CODEN JR-MAA8. ISSN 0075-4102 (print), 1435-5345 (electronic). URL

https://www.degruyter.com/
view/j/crll.1969.issue-239-
240/crll.1969.239-240.435/
crll.1969.239-240.435.xml.

Taussky:1969:ZKU

[Tau69d] Olga Taussky. Zbirovi kvadrata, uvodenje mladih u nauçni rad VI. (Serbian) [Zbirovi Squares, introduction of the young in scientific work VI]. Matematička Biblioteka (Belgrade, Yugoslavia), 41 (??):19–27, ???? 1969.

Taussky:1970:PDM

[Tau70a] Olga Taussky. Positive-definite matrices and their role in the study of the characteristic roots of general matrices. In Oved Shisha, editor, Inequalities 2, Proceedings of the Second Symposium on Inequalities held at the U.S. Air Force Academy, Colorado, August 1967, pages 389– 400. Academic Press, New York, USA, 1970.

Taussky:1970:RCS

[Tau70b] Olga Taussky. A remark concerning the similarity of a finite matrix A and A∗. Mathematische Zeitschrift, 117(1–4):189– 190, 1970. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic).

Taussky:1970:RPP

[Tau70c] Olga Taussky. Research problem 18, problems on matrices and operators. Bulletin of the American Mathematical Society, 76(5): 977–979, September 1970. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

tronic). URL http://www.ams.
org/journals/bull/1970-76-
05/S0002-9904-1970-12518-6/
S0002-9904-1970-12518-6.pdf.

Taussky:1970:RAC

[Tau70d] Olga Taussky. Reviews: The Art of Computer Programming. Volume 2: Seminumerical Algorithms, by Donald E. Knuth. American Mathematical Monthly, 77(8):900–901, October 1970.

CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1970:SCP

[Tau70e] Olga Taussky. Some computational problems in number theory. Bulletin of the Institute of Mathematics and its Applications, 6 (??):29–30, ???? 1970. CODEN IMTABW. ISSN 0950-5628.

Taussky:1970:SS

[Tau70f] Olga Taussky. Sums of squares. American Mathematical Monthly, 77(8):805–830, October 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). This article was awarded the Lester R. Ford Prize of the Mathematical Association of America for expository excellence.

Taussky:1971:SSH

[Tau71a] Olga Taussky. (1, 2, 4, 8)-sums of squares and Hadamard matrices. In Combinatorics, Proceedings of the Symposiums in Pure Mathematics, University of California, Los Angeles, California, 1968, volume 19, pages 229–233. American Mathematical Society, Providence, RI, USA, 1971.

Taussky:1971:HT

[Tau71b] Olga Taussky. Hilbert’s theorem 94. In Atkin and Birch [AB71], pages 65–71. ISBN 0-12-065750-3. LCCN QA241 .S35 1969.

Taussky:1971:RCC

[Tau71c] Olga Taussky. Result concerning classes of matrices. Notices of the American Mathematical Society, 18(6):943–??, 1971. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1971:SRM

[Tau71d] Olga Taussky. Some remarks on the matrix operator ad A. In Studies in Pure Mathematics, Papers presented to Richard Rado on the Occasion of his sixty-fifth Birthday, pages 235–237 (1971). Academic Press, London, UK, 1971.

Taussky:1972:AQF

[Tau72a] Olga Taussky. Automorphs of quadratic forms as positive operators. In Oved Shisha, editor, Inequalities, III (Proceedings of the Third Symposium, University of California, Los Angeles, California, 1969; dedicated to the memory of Theodore S. Motzkin), pages 341–345. Academic Press, New York, USA, 1972.

Taussky:1972:FIM

[Tau72b] Olga Taussky. The factorization of an integral matrix into two integral symmetric matrices. II. In ????, editor, Proceedings of the Number Theory Conference (Univ. Colorado, Boulder, Colo. 1972), pages 238–240. University of Colorado, Boulder, CO, USA, 1972.

Taussky:1972:PMH

[Tau72c] Olga Taussky. Pairs of matrices having property L. In Richard S. Varga, editor, Gatlinburg V Symposium on Numerical Algebra, June 1972, page ?? ????, ????,

                                        1972. URL http://www.math.
                                        vt.edu/HHXX/history.html.

Taussky:1972:RSM

[Tau72d] Olga Taussky. The role of symmetric matrices in the study of general matrices. Linear Algebra and its Applications, 5 (2):147–154, April 1972. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0024379572900249.

Taussky:1972:SRC

[Tau72e] Olga Taussky. Some results concerning the transition from the L- to the P -property for pairs of finite matrices. Journal of Algebra, 20:271–283, 1972. CO-DEN JALGA4. ISSN 0021-8693 (print), 1090-266x (electronic).

Taussky:1972:TPP

[Tau72f] Olga Taussky. Three problems, P91, P106, P107. Aequationes Mathematicae, 8(??):1972, 187, ???? 1972. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic).

Taussky:1973:FIMc

[Tau73a] Olga Taussky. The factorization of an integral matrix into a product of two integral symmetric matrices. Bulletin of the American Mathematical Society, 79:956–958, 1973. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

                                        tronic). URL http://www.ams.

org/journals/bull/1973-79-
05/S0002-9904-1973-13276-8/
S0002-9904-1973-13276-8.pdf.

Taussky:1973:FIMa

[Tau73b] Olga Taussky. The factorization of an integral matrix into a product of two integral symmetric matrices. I. Journal of Number Theory, 5:547–548, 1973. CO-DEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Taussky:1973:FIMb

[Tau73c] Olga Taussky. The factorization of an integral matrix into a product of two integral symmetric matrices. I. Acta Arithmetica, 24:151–156, 1973. CO-DEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic). Collection of articles dedicated to Carl Ludwig Siegel on the occasion of his seventy-fifth birthday, II.

Taussky:1973:FIMd

[Tau73d] Olga Taussky. The factorization of an integral matrix into a product of two integral symmetric matrices. II. The general case n = 2. Communications on Pure and Applied Mathematics (New York), 26:847–854, 1973. CODEN CPAMAT, CP-MAMV. ISSN 0010-3640 (print), 1097-0312 (electronic). Collection of articles dedicated to Wilhelm Magnus.

Taussky:1973:HTM

[Tau73e] Olga Taussky. Hilbert’s theorem 90 in matrix rings. Linear and Multilinear Algebra, 1 (1):5–8, 1973. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1973:RPa

[Tau73f] Olga Taussky. Research problem. Linear and Multilinear Algebra, 1(3):273, 1973. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1973:RPb

[Tau73g] Olga Taussky. Research problems. Linear and Multilinear Algebra, 1(??):273–??, ???? 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1973:SRC

[Tau73h] Olga Taussky. Some remarks concerning off diagonal part of a matrix. Notices of the American Mathematical Society, 20(6): A567–??, 1973. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1973:TRP

[Tau73i] Olga Taussky. Two research problems. Linear and Multilinear Algebra, 1(??):173–??, ???? 1973. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1974:ACB

[Tau74a] Olga Taussky. Additive commutators between 2 × 2 integral matrix representations of orders in identical or different quadratic number fields. Bulletin of the American Mathematical Society, 80:885–887, 1974. CO-DEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic).

Taussky:1974:MAC

[Tau74b] Olga Taussky. A matrix approach to a computational problem of Gauss in number theory, abstract. In ????, editor, Gatlinburg VI Symposium on Numerical Algebra, (December 1974), page ?? ????, ????, 1974.

Taussky:1974:RQS

[Tau74c] Olga Taussky. Response to query 26 (Small, Notices AMS 20:307, October 1973). Notices of the American Mathematical Society, 21(??):159, ???? 1974. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1974:RCC

[Tau74d] Olga Taussky. A result concerning classes of matrices. Journal of Number Theory, 6:64–71, 1974. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic).

Taussky:1974:SMW

[Tau74e] Olga Taussky. Some modern work on determinants. Invited address at San Francisco, annual meeting of the Mathematical Association of America, January 1974., 1974.

Taussky:1974:SRC

[Tau74f] Olga Taussky. Some results concerning the transition from the L- to the P -property for pairs of finite matrices II. Linear and Multilinear Algebra, 2(3):

195–202, 1974. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1975:ACR

[Tau75a] Olga Taussky. Additive commutators of rational 2 × 2 matrices. Linear Algebra and its Applications, 12(1):1–6, ???? 1975. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379575901214.

Taussky:1975:NQF

[Tau75b] Olga Taussky. Norms in quadratic fields and their relation to noncommuting 2 × 2 rational matrices. Notices of the American Mathematical Society, 22(2): A304–??, 1975. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1976:NQF

[Tau76a] Olga Taussky. Norms in quadratic fields and their relations to non commuting 2×2 matrices. I. Monatshefte für Mathematik, 82(3):253–255, 1976. CO-DEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic).

Taussky:1976:TFC

[Tau76b] Olga Taussky. Two facts concerning rational 2 × 2 matrices leading to integral ternary forms representing zero. In Selected topics on ternary forms and norms (Sem. Number Theory, California Inst. Tech., Pasadena, Calif., 1974/75), Paper No. 1, page 15.

California Institute of Technology, Pasadena, CA, USA, 1976. Paper number 1.

Taussky:1976:U

[Tau76c] Olga Taussky. [unknown]. In Neeta no shogai, pages 85–86. Tokyo-Tosho-Kabushiki-Kaishia, Tokyo, Japan, 1976. Japanese translation of poem on Emmy Noether in Japanese translation of the biography of Emmy Noether by Auguste Dick (see [Dic70b]).

Taussky:1977:MVP

[Tau77a] Olga Taussky. A matrix version of the principal genus in quadratic fields. In Grace Orzech, editor, Conference on quadratic forms 1976: [proceedings of the conference on quadratic forms held at Queen’s University, Kingston, Ontario, August 1–21, 1976], volume 46 of Queen’s papers in pure and applied mathematics, pages 632–633. Queen’s University, Kingston, ON, Canada, 1977.

Taussky:1977:NQFb

[Tau77b] Olga Taussky. Norms from quadratic fields and their relation to noncommuting 2 × 2 matrices. III. A link between the 4rank of the ideal class groups in Q(\sqrt{m}) and in Q(\sqrt{−m}). Mathematische Zeitschrift, 154(2):91– 95, 1977. CODEN MAZEAX. ISSN 0025-5874 (print), 1432-1823 (electronic).

Taussky:1977:NQFa

[Tau77c] Olga Taussky. Norms from quadratic fields and their relations to non commuting 2 × 2 matrices. II. The principal genus. Houston Journal of Mathematics, 3(4):543–547, 1977. ISSN 0362-1588.

Taussky:1977:RA

[Tau77d] Olga Taussky. Republication of articles 47, 85, 143. In Raymond W. Brink, editor, Selected Papers on Algebra. Mathematical Association of America, Washington, D.C., 1977.

Taussky:1978:AQD

[Tau78a] Olga Taussky. Answer to query 153 (L. D. Baumert, Notices AMS, 25:252 (1978)). Notices of the American Mathematical Society, 25(??):424, ???? 1978. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Taussky:1978:CAQ

[Tau78b] Olga Taussky. From cyclic algebras of quadratic fields to central polynomials. Journal of the Australian Mathematical Society. Series A, 25(4):503–506, 1978. CO-DEN JAMADS. ISSN 0263-6115 (print), 2396-8192 (electronic).

Taussky:1978:ICB

[Tau78c] Olga Taussky. Introduction into connections between algebraic number theory and integral matrices. In A Classical Invitation of Algebraic Numbers and Class Fields [Coh78], pages 305–326.

ISBN 1-4612-9950-0 (e-book), 0-387-90345-3 (print), 3-540-90345-3 (print). LCCN QA247 .C63.

URL http://catdir.loc.gov/
catdir/enhancements/fy0902/
78013785-d.html;   http://
catdir.loc.gov/catdir/enhancements/
fy0902/78013785-t.html;   http:
//www.gbv.de/dms/hbz/toc/ht000071512.

pdf. With two appendices by Olga Taussky: ”Artin’s 1932 Göttingen lectures on class field theory” and ”Connections between algebraic number theory and integral matrices.”.

Taussky:1978:LCF

[Tau78d] Olga Taussky. Lectures on class field theory by E. Artin (Göttingen 1932). In A Classical Invitation of Algebraic Numbers and Class Fields [Coh78], pages 277–304. ISBN 1-4612-9950-0 (e-book), 0-387-90345-3 (print), 3-540-90345-3 (print). LCCN QA247 .C63.

URL http://catdir.loc.gov/
catdir/enhancements/fy0902/
78013785-d.html;   http://
catdir.loc.gov/catdir/enhancements/
fy0902/78013785-t.html;   http:
//www.gbv.de/dms/hbz/toc/ht000071512.

pdf. With two appendices by Olga Taussky: ”Artin’s 1932 Göttingen lectures on class field theory” and ”Connections between algebraic number theory and integral matrices.”.

Taussky:1978:PCI

[Tau78e] Olga Taussky. Problem #19 (common invariant subspaces of operators and commutativity). In P. L. Butzer and B. Szökefalvi-Nagy, editors, Linear Spaces and Approximation: Proceedings of the Conference held at the Oberwolfach Mathematical Research Institute, Black Forest, August 20–27[, 1978], volume 40 of International Series of Numerical Mathematics, pages 669– 670. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1978.

Taussky:1979:DPA

[Tau79a] Olga Taussky. A Diophantine problem arising out of similarity classes of integral matrices. Journal of Number Theory, 11(3):472– 475, 1979. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). S. Chowla Anniversary Issue.

Taussky:1979:MSN

[Tau79b] Olga Taussky. My search for number theory. Lecture of MAA Conference, University of Southern California, March 10, 1979., 1979.

Taussky:1979:PNT

[Tau79c] Olga Taussky. Poem “Number theory”. In Ernest M. Robson and Jet Wimp, editors, Against infinity: an anthology of contemporary mathematical poetry, page 69. Primary Press, Parker Ford, PA, USA, 1979. ISBN 0-934982-00-7, 0-934982-01-5 (paperback). LCCN PN6110.M29 A34.

Taussky:1979:SRC

[Tau79d] Olga Taussky. Some remarks concerning matrices of the form A − A′, A^{−1}A′. Zeitschrift für Angewandte Mathematik und Physik = Journal of Applied Mathematics and Physics, 30(2):370–373, 1979. CODEN ZAMPDB. ISSN 0044-2275 (print), 1420-9039 (electronic).

Taussky:1980:A

[Tau80a] Olga Taussky. Autobiography, 1979–1980. In Olga Taussky-Todd (1906–1995): Autobiography, 1979–1980 [Ter80], pages 1–47.

URL http://oralhistories.
library.caltech.edu/43/1/OH_
Todd.pdf.

Taussky:1980:MNA

[Tau80b] Olga Taussky. More on norms from algebraic number fields, commutators and matrices which transform a rational matrix into its transpose. Linear Algebra and its Applications, 29(??): 459–464, February 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379580902566.

Taussky:1980:PST

[Tau80c] Olga Taussky. Pairs of sums of three squares of integers whose product has the same property. In General inequalities, 2 (Proceedings of the Second International Conference, Oberwolfach, 1978), volume 47, pages 29– 36. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1980. General inequalities 2, Proc. 2nd int. Conf., Oberwolfach 1978, ISNM.

Taussky:1980:SCM

[Tau80d] Olga Taussky. Sets of complex matrices which can be transformed to triangular forms. In Numerical methods, Keszthely 1977, Colloq. Math. Soc. Janos Bolyai, volume 22 of Colloq. Math. Soc. János Bolyai, pages 579–590. North-Holland Publishing Co., Amsterdam, The Netherlands, 1980.

Taussky:1980:SFC

[Tau80e] Olga Taussky. Some facts concerning integral representations of the ideals in an algebraic number field. Linear Algebra and its Applications, 31 (??):245–248, June 1980. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0024379580902232.

Taussky:1981:CBI

[Tau81a] Olga Taussky. Composition of binary integral quadratic forms via integral 2 × 2 matrices and composition of matrix classes. Linear and Multilinear Algebra, 10 (4):309–318, 1981. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1981:HSS

[Tau81b] Olga Taussky. History of sums of squares in algebra. In American mathematical heritage: algebra and applied mathematics (El Paso, Texas, 1975/Arlington, Texas, 1976), volume 13 of Math. Ser., pages 73–90. Texas Tech University, Lubbock, TX, USA, 1981.

Taussky:1981:MPR

[Tau81c] Olga Taussky. My personal recollection of Emmy Noether. In Brewer and Smith [BS81], pages 79–92. ISBN 0-8247-1550-0. LCCN QA29.N6 E47.

Taussky:1981:RPa

[Tau81d] Olga Taussky. Research problems 1, 2, 3. Linear and Multilinear Algebra, 10(??):165–166, ???? 1981. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1981:SFC

[Tau81e] Olga Taussky. Some facts concerning integral representations of ideals in an algebraic number field. Lecture Notes in Mathematics, 882:145–158, 1981. CODEN LNMAA2. ISBN 3-540-10880-7 (print), 3-540-38789-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/
10.1007/BFb0092490/. Integral

representations and applications, Proceedings of the Conference, Oberwolfach 1980.

Taussky:1982:MAP

[Tau82a] Olga Taussky. The many aspects of the Pythagorean triangles. Linear Algebra and its Applications, 43(??):285–295, March 1982. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0024379582902609.

Emmy Noether lecture at the AWM meeting in San Francisco, January 1981.

Taussky:1982:SUM

[Tau82b] Olga Taussky. Some uses of matrix theory in algebraic number theory. In ????, editor, OSU-Denison conference in honor of H. Zassenhaus, pages 44–46. ????, ????, June 1982.

Taussky:1982:TQF

[Tau82c] Olga Taussky, editor. Ternary Quadratic Forms and Norms, volume 79 of Lecture Notes in Pure and Applied Mathematics. Marcel Dekker, Inc., New York, NY, USA, 1982. ISBN 0-8247-1651-5 (paperback). LCCN QA243 .T43 1982. URL https:

//www.crcpress.com/Ternary-
Quadratic-Forms-and-Norms/
Taussky/p/book/9780824716516.

Lectures given in a numbertheory seminar at the California Institute of Technology in 1974/ 75.

Taussky:1982:TFC

[Tau82d] Olga Taussky. Two facts concerning rational 2 × 2 matrices leading to integral ternary forms representing zero. In Ternary Quadratic Forms and Norms [Tau82c], pages 39–48. ISBN 0-8247-1651-5 (paperback). LCCN QA243 .T43 1982. URL https:

//www.crcpress.com/Ternary-

Quadratic-Forms-and-Norms/
Taussky/p/book/9780824716516.

Lectures given in a numbertheory seminar at the California Institute of Technology in 1974/ 75.

Taussky:1982:TPI

[Tau82e] Olga Taussky. Two poems (Infinity and Song of the river). Totem: Caltech’s Literary and Visual Arts Magazine, ??(??):??, ???? 1982. URL http://

caltechtotem.github.io/.

Taussky:1983:CTP

[Tau83a] Olga Taussky. On the congruence transformation of a pencil of real symmetric matrices to a pencil with identical characteristic polynomial. Linear Algebra and its Applications, 52–53(??):687– 691, July 1983. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Taussky:1983:SNC

[Tau83b] Olga Taussky. Some noncommutative methods in algebraic number theory. In Srinivasan and Sally [SS83], pages 47– 57. ISBN 0-387-90838-2. LCCN QA1 .E52 1983.

Taussky:1983:SNM

[Tau83c] Olga Taussky. Some noncommutative methods in algebraic number theory. In Emmy Noether in Bryn Mawr (Bryn Mawr, PA, 1982), pages 47–57. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983.

Taussky:1984:SNL

[Tau84] Olga Taussky. The semigroup of non-left zero divisors in an algebra, infinite powers of matrices and related matters. Rocky Mountain Journal of Mathematics, 14(4):923–924, 1984. CO-DEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic). Ordered fields and real algebraic geometry (Boulder, Colo., 1983).

Taussky:1985:IMI

[Tau85a] Olga Taussky. Ideal matrices. III. Pacific Journal of Mathematics, 118(2):599–601, 1985. CO-DEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

URL http://projecteuclid.
org/euclid.pjm/1102706464.

Taussky:1985:RCP

[Tau85b] Olga Taussky. Remarks concerning possible connections between Fermat’s last theorem and integral p×p circulants. Linear Algebra and its Applications, 71(??): 295–297, November 1985. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0024379585902563.

Dedicated to H. Wielandt.

Taussky:1985:RPE

[Tau85c] Olga Taussky. Research problem “Eigenvalues, circles, ovals, and periods of complex functions”. Linear and Multilinear Algebra, 18(??):87–88, ???? 1985. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1985:TRP

[Tau85d] Olga Taussky. Two research problems on normal matrices. Linear and Multilinear Algebra, 17(??): 85–??, ???? 1985. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1986:CPC

[Tau86a] Olga Taussky. The characteristic polynomial and the characteristic curve of a pencil of matrices with complex entries. Österreichische Akademie der Wissenschaften Mathematisch-Naturwissenschaftliche Klasse. Sitzungsberichte. Abteilung II. Mathematische, Physikalische und Technische Wissenschaften, 195(1–3):175–178, 1986. ISSN 0723-9319. Dedicated to E. Hlawka.

Taussky:1986:FIM

[Tau86b] Olga Taussky. A factorization of an integral 2 × 2 matrix via a rational method. Monatshefte für Mathematik, 102(1):79–83, 1986. CODEN MNMTA2. ISSN 0026-9255 (print), 1436-5081 (electronic). Dedicated to E. Hlawka.

Taussky:1987:IMI

[Tau87a] Olga Taussky. Ideal matrices.

IV. In F. Uhlig and R. Grone, editors, Current trends in matrix theory, Proc. 3rd Conf., Auburn, Alabama 1986, pages 361–367. North-Holland Publishing Co., Amsterdam, The Netherlands, 1987.

Taussky:1987:IMA

[Tau87b] Olga Taussky. Integral matrices in algebraic number theory. In Number theory, Proceedings of the Conference, Montréal, Canada. 1985, volume 7 of CMS Conference Proceedings, pages 465–466. American Mathematical Society, Providence, RI, USA, 1987.

Taussky:1987:NCC

[Tau87c] Olga Taussky. Nonsingular cubic curves as determinantal loci. Historical remarks by John Todd. Journal of Mathematical and Physical Sciences, 21(6):665–678, 1987. ISSN 0047-2557. Ramanujan memorial volume.

Taussky:1987:RBT

[Tau87d] Olga Taussky. Radons Beitrag zur Theorie der Matrizen. (German) [Radon’s contribution to matrix theory]. In Peter Manfred Gruber, Edmund Hlawka, Wilfried Neubauer, and Leopold Schmetterer, editors, Johann Radon, Gesammelte Abhandlungen v. L, pages 299–306. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1987.

Taussky:1987:RKG

[Tau87e] Olga Taussky. Remembrances of Kurt Gödel. In Weingartner and Schmetterer [WS87], pages 29– 41. ISBN 88-7088-141-5. LCCN QA9.A1 G63 1987. L60000.

Taussky:1988:AGL

[Tau88a] Olga Taussky. Artin’s 1932 Göttingen lectures on class field theory. In A Classical Invitation to Algebraic Numbers and Class Fields [Coh88], pages 305– 326. ISBN 0-387-90345-3. LCCN QA247 .C63. With two appendices by Olga Taussky: ‘Artin’s 1932 Göttingen lectures on class field theory’ and ‘Connections between algebraic number theory and integral matrices’.

Taussky:1988:CRW

[Tau88b] Olga Taussky. Centennial reflections on women in American mathematics. AWM Newsletter, 18(6):10–11, ???? 1988.

Taussky:1988:CBA

[Tau88c] Olga Taussky. Connections between algebraic number theory and integral matrices. In A Classical Invitation to Algebraic Numbers and Class Fields [Coh88], pages 305–326. ISBN 0-387-90345-3. LCCN QA247 .C63. With two appendices by Olga Taussky: ‘Artin’s 1932 Göttingen lectures on class field theory’ and ‘Connections between algebraic number theory and integral matrices’.

Taussky:1988:PTS

[Tau88d] Olga Taussky. From Pythagoras’ theorem via sums of squares to celestial mechanics. The Mathematical Intelligencer, 10 (1):52–55, December 1988. CO-DEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Taussky:1988:HBT

[Tau88e] Olga Taussky. How I became a torchbearer for matrix theory.

American Mathematical Monthly, 95(9):801–812, November 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1988:MPR

[Tau88f] Olga Taussky. My personal recollections of Hans Heilbronn. In Kani and Smith [KS88], pages 27–37. ISBN 0-471-81655-8. LCCN QA241 .H372513 1988.

Taussky:1988:RKG

[Tau88g] Olga Taussky. Remembrances of Kurt Gödel. Engineering and Science [Caltech, Pasadena, CA], 51 (??):24–28, ???? 1988. Adapted by Jane Dietrich from [Tau87e].

Taussky:1988:ZGE

[Tau88h] Olga Taussky. Zeitzeugin. (German) [Eyewitness]. In Vertriebene Vernunft: Emigration und Exil österreichischer Wissenschaft, Internationale Symposion, 19. bis 23. Oktober 1987 in Wien. (German) [Displaced Reason: Emigration and Exile of Austrian Science, International Symposium, 19 to 23 October 1987 in Vienna] [Sta88], pages 132–134. ISBN 3-224-16525-1. LCCN ????

Taussky:1989:SSP

[Tau89a] Olga Taussky. Simultaneous similarities of pairs of 2 × 2 integral symmetric matrices. Rocky Mountain Journal of Mathematics, 19(3):957–965, 1989. CO-DEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic).

Quadratic forms and real algebraic geometry (Corvallis, OR, 1986).

Taussky:1989:SNM

[Tau89b] Olga Taussky. Some noncommutativity methods in algebraic number theory. In P. Duren, R. A. Askey, and U. Merzbach, editors, A century of mathematics in America, Part II, volume 2 of Hist. Math., pages 493–511. American Mathematical Society, Providence, RI, USA, 1989.

Taussky:1991:F

[Tau91a] Olga Taussky. Foreword. In An Introduction to multicomplex spaces and functions, pages iii–?? Marcel Dekker, Inc., New York, NY, USA, 1991.

Taussky:1991:LEB

[Tau91b] Olga Taussky. Letter to editor — Brouwer and Hilbert. The Mathematical Intelligencer, 13(3):4, Summer 1991. CO-DEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

URL https://link.springer.
com/article/10.1007/BF03023828.

Taussky:1996:WWD

[Tau96] Olga Taussky. Wielandt in Washington, DC, and Pasadena, CA, and commentaries on Wielandt’s papers. In Bertram Huppert and Hans Schneider, editors, Helmut Wielandt, Mathematische Werke, Volume 2: Linear Algebra and Analysis, pages 258–264. Walter de Gruyter, Berlin, Germany,

Taussky:1997:RHH

[Tau97] Olga Taussky. Recollections of Hans Hahn. In L. Schmetterer, K. Sigmund, and Karl Popper, editors, Hans Hahn, Gesammelte Abhandlungen, III, pages 570– 572. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997.

Taussky:19xx:VKV

[Tauxx]  Olga Taussky.  Verschiedene

kleinere Veröffentlichungen aus dem Gebiete der Mathematik. (German) [Various smaller publications from the field of mathematics]. ETH Zürich library catalog entry., 19xx.

Taussky:1979:RCC

[Tau80] Olga Taussky. Results concerning composition of sums of three squares. Linear and Multilinear Algebra, 8(3):231–233, 1979/80. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1969:PSS

[TC69] Olga Taussky and David Carlson. Problems and solutions: Solutions of advanced problems: 5582. American Mathematical Monthly, 76(3):314, March 1969. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /links.jstor.org/sici?sici=
                                        0002-9890(196903)76:3<314:
                                        5>2.0.CO;2-G&origin=MSN;
                                        http://www.jstor.org/stable/

2316402. See also [SST^{+}68].

Taussky:1954:APSa

[TE54] Olga Taussky and D. O. Ellis. Advanced problems and solutions: Solutions: 4504. American Mathematical Monthly, 61 (1):54–55, January 1954. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

URL http://links.jstor.org/
sici?sici=0002-9890(195401)
61:1<54:4>2.0.CO;2-A;   http:
//www.jstor.org/stable/2306907.

See also [Sha52, TS54].

Tent:2008:ENM

[Ten08] M. B. W. (Margaret B. W.) Tent.

Emmy Noether: the mother of modern algebra. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-430-5. xvii + 177 pp. LCCN QA29.N6 T46 2008.

Terrall:1980:OTT

[Ter80] Mary Terrall. Olga Taussky-Todd (1906–1995): Autobiography, 1979–1980. Technical report, Archives of the California Institute of Technology Oral History Project, Pasadena, CA, USA, 1980. 51 pp.

URL http://oralhistories.
library.caltech.edu/43/1/OH_
Todd.pdf.

Taussky:1985:RPF

[TF85] Olga Taussky and L. Foster. Research problem of L. Foster and Olga Taussky. Linear and Multilinear Algebra, 18(??):355–356, ???? 1985. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1980:PSS

[TG80] Olga Taussky and Robert M. Guralnick. Problems and solutions: Solutions of advanced problems: 6210. American Mathematical Monthly, 87(3):228–229, March 1980. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

                                        /links.jstor.org/sici?sici=
                                        0002-9890(198003)87:3<228:
                                        6>2.0.CO;2-2;   http://www.
                                        jstor.org/stable/2321630. See

also [MTP^{+}78]. Solution by R. M. Guralnick in American Mathematical Monthly, 87:228–229 (1980).

Taussky:1981:RPb

[TG81] Olga Taussky and R. M. Guralnick. Research problem. Linear and Multilinear Algebra, 10(??): 351, ???? 1981. CODEN LN-MLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky:1931:RBV

[TH31] Olga Taussky and H. Hahn. Review of B. L. van der Waerden’s Moderne Algebra, Bd. I. Monatshefte für Mathematik und Physik, 39(??):11–12, ???? 1931. CODEN MMPHA8.

Taussky:1932:RBV

[TH32] Olga Taussky and H. Hahn. Review of B. L. van der Waerden’s Moderne Algebra, Bd. II. Monatshefte für Mathematik und Physik, 40(??):3–4, ???? 1932. CODEN MMPHA8.

Taussky:1980:RCC

[TH80] Olga Taussky and J. Hsia. Results concerning composition of sums of three squares. Linear and Multilinear Algebra, 8(3): 231–233, 1980. CODEN LNM-LAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Thompson:1983:COT

[Tho83] Robert C. Thompson. A converse to the Ostrowski–Taussky determinantal inequality. Linear Algebra and its Applications, 52–53:693–699, July 1983. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Thebault:1958:APS

[TL58] Victor Thébault and Joe Lipman. Advanced problems and solutions: 4726. American Mathematical Monthly, 65(1):49–50, January 1958. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LET^{+}57].

Taussky:1961:APSa

[TM61] Olga Taussky and Marvin Marcus. Advanced problems and solutions: Solutions: 4894. American Mathematical Monthly, 68(2): 185–186, February 1961. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [LTK^{+}60].

Taussky:1963:APS

[TM63] Olga Taussky and Imanuel Marx. Advanced problems and solutions: Solutions: 5051.

American Mathematical Monthly, 70(8):903–904, October 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

links.jstor.org/sici?sici=
0002-9890(196310)70:8<903:
5>2.0.CO;2-B;   http://www.
jstor.org/stable/2312699. See

also [TWS^{+}62].

Todd:1944:NSS

[Tod44] Olga Todd. A note on skewsymmetric matrices. Reports and Memoranda 2006 (8095 & 8447), Ministry of Aircraft Production, Aeronautical Research Committee, London, UK, October 1944. 16–18 pp. Appendix to Duncan, W. J., Properties of characteristic numbers and modes deduced by matrix methods.

Todd:1945:SBV

[Tod45] Olga Todd. On some boundary value problems in the theory of the non-uniform supersonic motion of an aerofoil. Reports and memoranda 2141(8961), Ministry of Supply, Aeronautical Research Council, London, UK, 1945. 13 pp.

Todd:1962:SNA

[Tod62]  John Todd, editor.  Survey

of Numerical Analysis. Mc-Graw-Hill, New York, NY, USA, 1962. xvi + 589 pp. LCCN QA297 .T63. URL https:

//babel.hathitrust.org/cgi/
pt?id=mdp.39015000983901.

Todd:1978:BNM

[Tod78] John Todd. Basic Numerical Mathematics, volume 14, 22 of International series of numerical mathematics. Academic Press, New York, USA, 1978. ISBN 0-12-692401-5 (vol. 1), 0-12-692402-3 (vol. 2). 253 (vol. 1, est.), 250 (vol. 2, est.) pp. LCCN QA297 .T58 1978.

Todd:1983:O

[Tod83] John Todd. Oberwolfach — 1945. In Beckenbach and Walter [BW83], pages 19–22. ISBN 3-7643-1539-3. LCCN QA295 .I57 1981. URL https://

link.springer.com/chapter/
10.1007/978-3-0348-6290-5_2.

Todd:1997:VLG

[Tod97] Olga Taussky Todd. Der vogel auf dem leim. (German) [The bird on the glue]. The Mathematical Intelligencer, 19(1):17, December 1997. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Taussky:1960:APSa

[TP60] Olga Taussky and W. V. Parker.

Advanced problems and solutions: Solutions: 4846. American Mathematical Monthly, 67(2): 192–193, February 1960. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [KTB^{+}59].

Trafford:1992:NID

[Tra92] Paul Joseph Trafford. Norms of ideals in direct sums of number fields and applications to the circulants problems of Olga Taussky-Todd. M.Sc. thesis, University of Glasgow, Glasgow, UK, 1992. ????

pp.  URL https://search.
proquest.com/docview/301521435.

Tropp:1973:TIJ

[Tro73] H. S. Tropp. Transcript of interview of J.T. and O.T. Smithsonian Museum, July 2, 1973.

Taussky:1934:HKK

[TS34] Olga Taussky and A. Scholz. Die Hauptideale der kubischen Klassenkörper imaginärquadratischer Zahlkörper: ihre rechnerische Bestimmung und ihr Einfluß auf den Klassenkörperturm. (German) [The main ideals of the cubic class body imaginary quadratic number field: their mathematical determination and its influence on the class field tower]. Journal für die reine und angewandte Mathematik, 171:19– 41, 1934. CODEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic).

Taussky:1954:APSb

[TS54] Olga Taussky and D. W. Sasser. Advanced problems and solutions: Solutions: 4504. American Mathematical Monthly, 61(5):351, May 1954. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/links.jstor.org/sici?sici=
0002-9890(195405)61:5<351:
4>2.0.CO;2-3;   http://www.
jstor.org/stable/2307487. See

also [Sha52, TE54].

Taussky:1970:PSS

[TS70] Olga Taussky and J. B. Skinner. Problems and solutions: Solutions of advanced problems: 5692. American Mathematical Monthly, 77(8):894, October 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/links.jstor.org/sici?sici=
0002-9890(197010)77:8<894:
5>2.0.CO;2-I;   http://www.
jstor.org/stable/2317048. See

also [DKB^{+}69].

Taussky:1958:APS

[TSE58] Olga Taussky, P. P. Saworotnow, and Howard Eves. Advanced problems and solutions: Solutions: 4725. American Mathematical Monthly, 65(1):49, January 1958. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:/

/links.jstor.org/sici?sici=
0002-9890(195801)65:1<49:4>
2.0.CO;2-H;   http://www.jstor.
org/stable/2310318. See also

[LET^{+}57].

Taussky:1940:CAN

[TT40a] Olga Taussky and John Todd.

A characterisation of algebraic numbers. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 46:1–8, 1940. CODEN PRIAAK. ISSN 0035-8975.

Taussky:1940:DQ

[TT40b] Olga Taussky and John Todd.

Determinants of quaternions.

Bulletin of the American Mathematical Society, 46(5):431–432, May 1940. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (electronic). URL http:/

                                        /www.ams.org/journals/bull/
                                        1940-46-05/S0002-9904-1940-
                                        07220-9/S0002-9904-1940-07220-
                                        9.pdf.

Taussky:1940:MFP

[TT40c] Olga Taussky and John Todd. Matrices with finite period. Proceedings of the Edinburgh Mathematical Society (2), 6:128–134, 1940. CODEN PEMSA3. ISSN 0013-0915 (print), 1464-3839 (electronic).

Taussky:1941:IG

[TT41a] Olga Taussky and John Todd. Inversion in groups. Quarterly Journal of Mathematics. Oxford series., 12:65–67, 1941. CO-DEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Taussky:1941:MFP

[TT41b] Olga Taussky and John Todd. Matrices of finite period. Proceedings of the Royal Irish Academy, Section A: Mathematical and Physical Sciences, 46:113–121, 1941. CODEN PRIAAK. ISSN 0035-8975.

Taussky:1942:IPM

[TT42] Olga Taussky and John Todd. Infinite powers of matrices. Journal of the London Mathematical Society, 17:146–151, 1942. CO-DEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Todd-Taussky:1944:ANS

[TT44] Olga Todd-Taussky. Appendix.

A note on skew-symmetric matrices. Report and Memoranda 2006 (8095 & 8447), Ministry of Aircraft Production, Aeronautical Research Committee, London, UK, 1944. 16–18 pp.

Taussky:1947:SAM

[TT47] Olga Taussky and John Todd.

Some aspects of modern algebra. Science Progress, 138:253–268, 1947. ISSN 0036-8075 (print), 1095-9203 (electronic).

Taussky:1948:CTGa

[TT48a] Olga Taussky and John Todd.

Covering theorems for groups. Bulletin of the American Mathematical Society, 54(3):276, March 1948. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1948:CTGb

[TT48b] Olga Taussky and John Todd.

Covering theorems for groups. Annales de la Societé Polonaise de Mathématique, 21:303– 305 (1949), 1948.

Taussky:1949:CTG

[TT49] Olga Taussky and John Todd.

Covering theorems for groups. Ann. Soc. Polon. Math., 21:303– 305, 1949.

Taussky:1951:CRP

[TT51] Olga Taussky and John Todd.

The characteristic roots of products of matrices with rational integral elements. Bulletin of the American Mathematical Society, 57(2):123, March 1951. CO-DEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Taussky:1952:SEMa

[TT52a] Olga Taussky and John Todd. Systems of equations, matrices and determinants. I. Mathematics Magazine, 26(2):9–20, September/October 1952. CODEN MAMGA8. ISSN 0025-570X (print), 1930-0980 (electronic).

Taussky:1952:SEMb

[TT52b] Olga Taussky and John Todd. Systems of equations, matrices and determinants. II. Mathematics Magazine, 26(2):71–88, September/October 1952. CO-DEN MAMGA8. ISSN 0025-570X (print), 1930-0980 (electronic).

Taussky:1954:GTP

[TT54] Olga Taussky and John Todd. Generation and testing of pseudorandom numbers. Report 3370, U.S. National Bureau of Standards, Gaithersburg, MD, USA, June 22, 1954. ii + 16 pp. URL https:/

/nvlpubs.nist.gov/nistpubs/
Legacy/RPT/nbsreport3370.pdf.

Taussky:1956:CBT

[TT56a] Olga Taussky and John Todd. Commuting bilinear transformations and matrices. Journal of the Washington Academy of Sciences, 46(12):373–375, December 1956. CODEN JWASA3. ISSN 0043-0439. URL http://www.

jstor.org/stable/24533908.

Taussky:1956:GTP

[TT56b] Olga Taussky and John Todd.

Generation and testing of pseudorandom numbers. In Meyer [Mey56], pages 15–28. LCCN QA273 U577.

Taussky:1957:ASE

[TT57a] Olga Taussky and John Todd.

Applications of systems of equations, matrices and determinants. In James and Beckenbach [JB57], chapter 24, pages 310–337. LCCN

QA37 .J26. URL https://www.
questia.com/read/86015503/
the-tree-of-mathematics.

Taussky:1957:SEM

[TT57b] Olga Taussky and John Todd.

Systems of equations, matrices and determinants. In James and Beckenbach [JB57], chapter 23, pages 305–318. LCCN

QA37 .J26. URL https://www.
questia.com/read/86015489/
the-tree-of-mathematics.

Taussky:1960:SDV

[TT60] Olga Taussky and John Todd.

Some discrete variable computations. Proceedings of Symposia in Applied Mathematics, 10:201– 209, 1960. CODEN PSYMA7. ISSN 0160-7634. URL http:

//www.ams.org/books/psapm/
010/0115261.

Taussky-Todd:1977:OTT

[TT77]  Olga Taussky-Todd.  Olga

Taussky-Todd. In Zassenhaus [Zas77], pages xxxiv–xlvi + 1. ISBN 0-12-776350-3. LCCN QA241 .N866.

Taussky-Todd:1978:SCC

[TT78] Olga Taussky-Todd. Some comments concerning G. B. Price’s note on: “Determinants with dominant principal diagonal: a personal historical note” (Linear and Multilinear Algebra 6 (1978/79), no. 4, 247–250). Linear and Multilinear Algebra, 6(4): 251, 1978. CODEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic). See [Pri78].

Taussky-Todd:1985:AE

[TT85a] Olga Taussky-Todd. Autobiographical essay. In Albers and Alexanderson [AA85], pages 321–

352. ISBN 0-8176-3191-7, 3-7643-3191-7. LCCN QA28 .M37 1985.

Taussky-Todd:1985:RP

[TT85b] Olga Taussky-Todd. Research problems. Linear and Multilinear Algebra, 17(1):85, 1985. CO-DEN LNMLAZ. ISSN 0308-1087 (print), 1563-5139 (electronic).

Taussky-Todd:1988:HOW

[TT88] Olga Taussky-Todd. [In her own words]. AWM Newsletter, 18(6):10–11, November/ December 1988. URL https:

                                        //www.drivehq.com/file/df.
                                        aspx/isGallarytrue/shareID8755087/
                                        fileID748748559. Reprinted in

a similar form in [TT05].

Taussky:1991:ALM

[TT91] Olga Taussky and John Todd.

Another look at a matrix of Mark Kac. Linear Algebra and its Applications, 150:341–360, 1991. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Proceedings of the First Conference of the International Linear Algebra Society (Provo, UT, 1989).

Taussky-Todd:2005:HOW

[TT05] Olga Taussky-Todd. In her own words. In Case and Leggett [CL05], pages 4–5. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (e-book). LCCN QA27.5 .C66 2005. URL

http://www.loc.gov/catdir/
enhancements/fy0668/2004048843-
b.html;   http://www.loc.gov/
catdir/enhancements/fy0668/
2004048843-d.html;   http:/
/www.loc.gov/catdir/toc/fy0612/
2004048843.html. Based on a

talk at the panel “Centennial Reflections on Women in American Mathematics”, organized by Judy Green and Jeanne LaDuke, AMS Centennial Meeting, Providence, 1988 as reported in the AWM Newsletter 18(6) 1988, 10– 11. Taussky Todd (1906–1995) was then a professor emerita of mathematics, California Institute of Technology.

Taussky:2006:CTT

[TT06] Olga Taussky and John Todd.

Cholesky, Toeplitz and the triangular factorization of symmetric matrices. Numerical Algorithms, 41(2):197–202, February 2006. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

                                        tronic).  URL http://www.
                                        springerlink.com/openurl.asp?
                                        genre=article&issn=1017-1398&
                                        volume=41&issue=2&spage=197.

Taussky:2012:ASE

[TT12a] Olga Taussky and John Todd. Applications of systems of equations, matrices and determinants. In James and Beckenbach [JB12], chapter 24, pages 310–337. ISBN 1-258-28562-2. LCCN QA37 .J26.

Taussky:2012:SEM

[TT12b] Olga Taussky and John Todd. Systems of equations, matrices and determinants. In James and Beckenbach [JB12], chapter 23, pages 305–318. ISBN 1-258-28562-2. LCCN QA37 .J26.

Taussky-Todd:1957:ASE

[TTT57] Olga Taussky-Todd and John Todd. Applications of systems of equations, matrices, and determinants. Mathematics Magazine, ?? (??), 1957. CODEN MAMGA8. ISSN 0025-570X. Special issue: “The Tree of Mathematics”.

Taussky:1948:MNN

[TW48a] Olga Taussky and L. A. Wigglesworth. Mathematical notes: Note on a theorem in ndimensional geometry. American Mathematical Monthly, 55 (8):492–494, October 1948. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1948:NTD

[TW48b] Olga Taussky and L. A. Wigglesworth. Note on a theorem in n-dimensional geometry. American Mathematical Monthly, 55: 492–494, 1948. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Taussky:1962:LRB

[TW62a] Olga Taussky and Helmut Wielandt. Linear relations between higher additive commutators. Proceedings of the American Mathematical Society, 13:732–735, 1962. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (elec-

tronic). URL http://www.ams.
org/journals/proc/1962-013-
05/S0002-9939-1962-0148684-
6/S0002-9939-1962-0148684-
6.pdf.

Taussky:1962:MF

[TW62b] Olga Taussky and Helmut Wielandt. On the matrix function AX + X′A′. Archive for rational mechanics and analysis, 9:93–96, 1962. CODEN AVR-MAW. ISSN 0003-9527 (print), 1432-0673 (electronic).

Taussky:1963:RDS

[TW63] Olga Taussky and Helmut Wielandt. On the role of the determinant in semigroups of matrices. Quarterly Journal of Mathematics, 14(1):123–130, 1963. CO-DEN QJMAAT. ISSN 0033-5606 (print), 1464-3847 (electronic).

Tarwater:1981:AMH

[TWHM81] Dalton Tarwater, John T. White, Carl Hall, and Marion E. Moore, editors. American Mathematical Heritage: Algebra and Applied Mathematics[: papers of the third conference, held at the University of Texas at El Paso, November 1975, and the fourth conference, held at the University of Texas at Arlington, October 1976], volume 13 of Mathematics Series. Texas Technological University Press, Lubbock, TX, USA, 1981.

Taussky:1962:APS

[TWS^{+}62] Olga Taussky, Seth Warner, H. S. Shapiro, Richard Sinkhorn, Peter Yff, C. W. Kohls, M. E. Mahowald, and F. M. Hornyak. Advanced problems and solutions: Problems for solution: 5051– 5057. American Mathematical Monthly, 69(9):925–926, November 1962. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [HC63, KMCS63, KMS64, SD63, SK63, TM63, WB63, War64, YCJ63, Yff64].

Taussky:1959:STB

[TZ59] Olga Taussky and Hans Zassenhaus. On the similarity transformation between a matrix and its transpose. Pacific Journal of Mathematics, 9:893– 896, 1959. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic). URL http:

//projecteuclid.org/euclid.
pjm/1103039127.

Taussky:1969:SCC

[TZ69] Olga Taussky and H. Zassenhaus. Short communications: On the l-cohomology of the general and special linear group I. Aequationes Mathematicae, 3(1–2): 207–209, February 1969. CO-DEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic).

Taussky:1970:CGS

[TZ70] Olga Taussky and Hans Zassenhaus. On the 1-cohomology of the general and special linear groups. Aequationes Mathematicae, 5:129–201, 1970. CO-DEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic).

Ungar:1960:APS

[UEM^{+}60] Peter Ungar, Paul Erdős, Leo Moser, S. W. Golomb, and Olga Taussky. Advanced Problems and Solutions: Problems for Solution: 4911-4916. American Mathematical Monthly, 67(6):596–597, June/July 1960. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EY61, ES61, GLF61, MC61, Tau61a, UvL61].

Uhlig:2007:OTT

[Uhl07] Frank Uhlig. Olga Taussky, a torchbearer for mathematics and teacher of mathematicians. Proceedings in Applied Mathematics and Mechanics, 7(1):1170501– 1170503, December 2007. ISSN 1617-7061.

Ungar:1961:APS

[UvL61] Peter Ungar and J. H. van Lint. Advanced problems and solutions: Solutions: 4911. American Mathematical Monthly, 68(5): 512, May 1961. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UEM^{+}60].

vanderWaerden:1976:QFT

[van76] B. L. van der Waerden. On the 1st and the 2nd quadratic form. Two letters of B. L. van der Waerden to Olga Taussky-Todd. In Selected Topics on Ternary Forms and Norms, Seminar in Number Theory, California Institute of Technology 1974/75, Exp. 2, 12 p. (1976). California Institute of Technology, Pasadena, CA, USA, 1976. Translated from the German originals by H. Zassenhaus.

Varga:1962:MIA

[Var62] Richard S. Varga. Matrix Iterative Analysis. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1962. 322 pp. LCCN QA263 .V3 1962.

Varga:2000:MIA

[Var00] Richard S. Varga. Matrix Iterative Analysis, volume 27 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 3-642-05154-5 (paperback), 3-540-66321-5 (hardcover), 3-642-05156-1 (e-book). x + 358 pp. LCCN QA297.8 .V37 2000.

URL http://lcweb.loc.gov/
catdir/toc/99050196.html;
http://www.loc.gov/catdir/
enhancements/fy0815/99050196-
d.html;   http://www.springerlink.
com/openurl.asp?genre=book&
isbn=978-3-642-05161-6.

Varga:2004:GHC

[Var04] Richard S. Varga. Geršgorin and His Circles, volume 36 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21100-4 (hardcover), 3-642-17798-0 (e-book). ISSN 0179-3632. x + 226 pp. LCCN QA193 .V37 2004. URL

http://www.loc.gov/catdir/
enhancements/fy0815/2004104814-
d.html.

Varga:2009:MIA

[Var09] Richard S. Varga. Matrix Iterative Analysis, volume 27 of Springer series in computational mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second revised and expanded edition, 2009. ISBN 3-642-05154-5 (softcover), 3-642-05156-1 (ebook). ISSN 0179-3632 (print), 2198-3712 (electronic). x + 358 pp. LCCN QA263 .V3 2009.

vanderWaerden:1931:MAB

[vdW31] B. L. van der Waerden. Moderne Algebra. Band II. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1931. ???? pp.

Volk:1968:PSS

[VM68] Benjamin Volk and M. E. Muldoon. Problems and solutions: Solutions of advanced problems: 5466. American Mathematical Monthly, 75(2):206, February 1968. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Warner:1964:APS

[War64] Seth Warner. Advanced problems and solutions: Solutions: 5052. American Mathematical Monthly, 71(3):328, March 1964. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TWS^{+}62, WB63].

Warner:1963:APS

[WB63] Seth Warner and I. N. Baker. Advanced problems and solutions: Solutions: 5052. American Mathematical Monthly, 70(8): 904, October 1963. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TWS^{+}62, War64].

Weiss:1996:OTT

[Wei96] E. Weiss. Olga Taussky-Todd — obituary. IEEE Annals of the History of Computing, 18(3):60– 61, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:

                                        //ieeexplore.ieee.org/iel4/
                                        85/11673/00539918.pdf.

Williams:1968:APS

[WF68] S. W. Williams and S. P. Franklin. Advanced problems and solutions: Solutions: 5468. American Mathematical Monthly, 75(2):208, February 1968. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BKM^{+}67].

Williams:1969:PSS

[WF69] S. W. Williams and R. V. Fuller.

Problems and solutions: Solutions of advanced problems: 5583. American Mathematical Monthly, 76(3):314–315, March 1969. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SST^{+}68].

White:1986:ACM

[Whi86] Marge White. AAAS Council meeting, 1986. Science, 233(4763):575–577, August 1986. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). The article reports that Olga Taussky-Todd is made a Member of the American Association for the Advancement of Science on 29 May 1986.

Godel:1987:GRS

[WS87] Paul Weingartner and Leopold Schmetterer, editors. Gödel remembered: Salzburg 10–12 July 1983, volume 4 of History of logic. Bibliopolis, Napoli [Naples], Italy, 1987. ISBN 88-7088-141-5. LCCN QA9.A1 G63 1987. L60000.

Washington:1974:PSS

[WW74] L. C. Washington and Seth Warner. Problems and solutions: Solutions of advanced problems: 5920. American Mathematical Monthly, 81(8):912–913, October 1974. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Mau73].

Yff:1963:APS

[YCJ63] Peter Yff, David Carlson, and H. G. Jacob. Advanced problems and solutions: Solutions: 5055. American Mathematical Monthly, 70(9):1016–1017, November 1963. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TWS^{+}62, Yff64].

Yff:1964:APS

[Yff64] Peter Yff. Advanced problems and solutions: Solutions: 5055. American Mathematical Monthly, 71(3):328, March 1964. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [TWS^{+}62, YCJ63].

Yin:1994:CIF

[Yin94] Xiangrong Yin. A converse inequality of Fan, Taussky, and Todd. Journal of Mathematical Analysis and Applications, 182 (3):654–657, March 15, 1994. CO-DEN JMANAK. ISSN 0022-247x (print), 1096-0813 (electronic).

Zassenhaus:1961:RPO

[Zas61] Hans Zassenhaus. A remark on a paper of Olga Taussky. Journal of Mathematics and Mechanics, 10: 179–180, 1961. CODEN JOM-MAN. ISSN 0095-9057 (print), 1943-5274 (electronic).

Zassenhaus:1977:NTA

[Zas77] Hans Zassenhaus, editor. Number Theory and Algebra: Collected Papers Dedicated to Henry B. Mann, Arnold E. Ross, and Olga Taussky-Todd. Academic Press, New York, USA, 1977. ISBN 0-12-776350-3. xviii + 393 pp. LCCN QA241 .N866.

Zund:2000:TTO

[Zun00] Joseph D. Zund. Taussky-Todd, Olga (1906–1995), mathematician. In Carnes [Car00], page [various]. ISBN 0-19-520635-5 (24-volume set), 0-19-860669-9 (online). LCCN CT213. URL

http://www.anb.org.

Zhang:2005:SNP

[ZX05] Zhi-Hua Zhang and Zhen-Gang Xiao. A simple new proof of Fan–Taussky–Todd inequalities. The Australian Journal of Mathematical Analysis and Applications, 2(1):5, 2005. ISSN 1449-5910 (electronic). URL http:

//ajmaa.org/cgi-bin/paper.
pl?string=v2n1/V2I1P5.tex;
 http://ajmaa.org/searchroot/
files/pdf/v2n1/v2i1p5.pdf.